### v2

# On von Neumann regularity of ample groupoid algebras

Abstract.
We study the question of when the algebra of an ample groupoid over a field is von Neumann regular and, more generally, when the algebra of a graded ample groupoid is graded von Neumann regular. For fields of characteristic 00 we are able to provide a complete characterization. Over fields of positive characteristic, we merely have some sufficient and some necessary conditions for regularity. As applications we recover known results on regularity and graded regularity of Leavitt path algebras and inverse semigroups. We also prove a number of new results, in particular concerning graded regularity of algebras of Deaconu-Renault groupoids and Nekrashevych-Exel-Pardo algebras of self-similar groups.
Key words and phrases:
von Neumann regular rings, Leavitt path algebras, ample groupoids, groupoid algebras
2020 Mathematics Subject Classification:
20M25, 16S88, 22A22, 20M18
The first author was supported by a Simons Foundation Collaboration Grant, award number 849561, and the Australian Research Council Grant DP230103184.
1. Introduction
The notion of a (von Neumann) regular ring was invented by von Neumann in the 1930s in his study of what are now called von Neumann algebras. A ring (or more generally a semigroup) is regular if for every $a$, there is a $b$ with $aba=a$. It was later observed that regular rings are precisely the rings for which every module is flat, and so from the point of view of homological algebra they are a natural generalization of semisimple rings.
A famous result of Connell [12] says that a group ring $KG$, over a field $K$, is regular if and only if $G$ is locally finite and the characteristic of $K$ does not divide the order of any element of $G$. In particular, this says that regularity is a rare feature for groups rings. Leavitt path algebras [1] are also rarely regular: they are regular if and only if the (directed) graph defining the Leavitt path algebra is acyclic [3]. However, every Leavitt path algebra is graded von Neumann regular by a result of Hazrat [20]. A graded ring $R$ is graded regular if the semigroup of homogeneous elements is regular. This has an important impact on the graded $K$-theory of Leavitt path algebras.
Okniński tried to generalize Connell’s results to inverse semigroup algebras (see his book [28]). He was able to prove that if $K$ is a field of characteristic 00 and $S$ is an inverse semigroup, then $KS$ is regular if and only if $S$ is locally finite. In positive characteristic he was able to give necessary conditions: $S$ is periodic, each maximal subgroup of $S$ is locally finite and the characteristic of $K$ cannot divide the order of any element of a maximal subgroup. It had been proven earlier that if $S$ is locally finite and the characteristic of $K$ divides the order of no element of any maximal subgroup, then $KS$ is regular. But being periodic with locally finite maximal subgroups is a strictly weaker condition than being locally finite. The reason that Okniński was able to obtain a complete characterization in characteristic 00 is that he could embed $\mathbb{C}S$ into the Banach algebra $\ell_{1}(S)$ and use some analysis.
Group algebras, inverse semigroup algebras and Leavitt path algebras are all special cases of algebras of ample groupoids [31]. It seems natural then to investigate regularity of ample groupoid algebras. A first result in this direction was obtained in [4], where it was show that if $K$ is a field and $\mathscr{G}$ is an ample groupoid, then a necessary condition for the groupoid algebra $K\mathscr{G}$ to be regular is that each isotropy group of $\mathscr{G}$ is locally finite and the order of no element of any isotropy group is divisible by the characteristic of $K$. The proof is a reduction to Connell’s result.
Our main result is a complete characterization of regularity of $K\mathscr{G}$ when $K$ is a field of characteristic 00. Following the Bourbaki tradition, we say a space is quasi-compact if every open covering has a finite subcover and reserve “compact” to mean quasi-compact and Hausdorff. Then $K\mathscr{G}$ is regular if and only if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids. For principal, second countable Hausdorff groupoids this is equivalent to approximate finiteness in the sense of [18], hence we call such groupoids approximately quasi-compact. Like Okniński, we use analytic methods, but this time we embed $\mathbb{C}\mathscr{G}$ into the reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$.
In positive characteristic, we are able to show that if $\mathscr{G}$ is approximately quasi-compact and the characteristic of $K$ does not divide the order of any isotropy element of $\mathscr{G}$, then $K\mathscr{G}$ is regular. But we are unable to prove in general the necessity of being approximately quasi-compact in positive characteristic.
We also consider graded regularity of groupoid algebras for ample groupoids equipped with a locally constant cocycle to a group. We prove the algebra is graded regular if and only if the homogeneous component of the identity is regular. Since the homogeneous component of the identity is the algebra of the groupoid of elements mapping to the identity under the cocycle, our results for regularity of groupoid algebras apply. From this we easily recover Hazrat’s result on graded regularity of Leavitt path algebras [20]. We extend these results to algebras of Deaconu-Renault groupoids under mild hypotheses. These include algebras of row finite higher rank graphs. We consider further graded regularity of partial group skew products, Leavitt path algebras of labeled graphs and Nekrashevych-Exel-Pardo algebras of self-similar graphs.
The paper is organized as follows. Section 2 is devoted to preliminaries concerning inverse semigroups, ample groupoids and regular rings. Then Section 3 proves our main results on regularity of ample groupoid algebras. Section 4 addresses the question of graded regularity of the algebra of an ample groupoid equipped with a locally constant cocycle. The final section, applies the main results to study regularity and graded regularity for the algebras of a number of well-studied families of ample groupoids.
2. Preliminaries
We collect here some preliminaries on inverse semigroups, ample groupoids and von Neumann regular rings.
2.1. Inverse semigroups and groupoids
The reader is referred to [22] for details on the theory of inverse semigroups. A semigroup is a set together with an associative binary operation. An inverse semigroup is a semigroup $S$ such that for every $s\in S$ there is a unique $s^{*}\in S$ such that $s^{*}ss^{*}=s^{*}$ and $ss^{*}s=s$. Notice that $ss^{*},s^{*}s$ are idempotents for all $s\in S$. The set $E(S)$ of idempotents of $S$ is a commutative subsemigroup and hence a meet semilattice, with the partial order given by $e\leq f$ if and only if $ef=e$, and the meet by $e\wedge f=ef$. The order extends to $S$ by defining $s\leq t$ if $s=te$ for some $e\in E(S)$. If $e\in E(S)$, then $G_{e}=\{s\in S\mid s^{*}s=e=ss^{*}\}$ is a group with identity $e$ known as the maximal subgroup of $S$ at $e$.
A topological groupoid is a groupoid $\mathscr{G}$ endowed with a locally compact topology such that composition and inversion are continuous. We do not assume that $\mathscr{G}$ is Hausdorff in general, but we do assume that the unit space $\mathscr{G}^{(0)}=\{\gamma\gamma^{-1}\mid\gamma\in\mathscr{G}\}$ is Hausdorff with the relative topology inherited from $\mathscr{G}$. Since we do not assume that $\mathscr{G}$ is Hausdorff, locally compact is defined as every point in $\mathscr{G}$ having a compact neighborhood, where compact means Hausdorff and open covers have finite subcovers. Following the Bourbaki tradition, we say that a space $X$ is quasi-compact if every open covering of $X$ has a finite subcovering (but with no assumption of being Hausdorff). Note that a finite union of quasi-compact sets is quasi-compact and a closed subset of a quasi-compact set is quasi-compact.
An étale groupoid is a topological groupoid $\mathscr{G}$ such that the range map $\mathop{\boldsymbol{r}}\colon\mathscr{G}\to\mathscr{G}^{(0)}$ (or equivalently, the source map $\mathop{\boldsymbol{s}}$) is a local homeomorphism. A bisection of an étale groupoid $\mathscr{G}$ is an open subset $U\subseteq\mathscr{G}$ such that the restrictions $\mathop{\boldsymbol{r}}|_{U}$ and $\mathop{\boldsymbol{s}}|_{U}$ are homeomorphisms onto an open set in $\mathscr{G}^{(0)}$. In particular, bisections are always Hausdorff subspaces. An étale groupoid $\mathscr{G}$ is an ample groupoid if $\mathscr{G}^{(0)}$ is a locally compact Hausdorff space with a basis of compact open sets.
If $x\in\mathscr{G}^{(0)}$, then the isotropy group at $x$ is
$\mathscr{G}^{x}_{x}=\{\gamma\in\mathscr{G}\mid\mathop{\boldsymbol{s}}(\gamma)=x=\mathop{\boldsymbol{r}}(\gamma)\}.$
Then $\mathop{\mathrm{Iso}}(\mathscr{G})=\bigcup_{x\in\mathscr{G}^{(0)}}\mathscr{G}^{x}_{x}$ is a closed subgroupoid called the isotropy subgroupoid (or bundle). A groupoid consisting of only isotropy is called a group bundle. A groupoid with trivial isotropy groups is called principal.
If $X\subseteq\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is the subgroupoid of all arrows $\gamma$ with $\mathop{\boldsymbol{s}}(\gamma),\mathop{\boldsymbol{r}}(\gamma)\in X$, i.e., $\mathscr{G}|_{X}=\mathop{\boldsymbol{s}}^{-1}(X)\cap\mathop{\boldsymbol{r}}^{-1}(X)$. In particular, if $X$ is open (respectively, closed) in $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is open (respectively, closed) in $\mathscr{G}$.
The orbit of $x\in\mathscr{G}^{(0)}$ is the set $\mathcal{O}_{x}$ consisting of all $y\in\mathscr{G}^{(0)}$ such that there is an arrow $\gamma\colon x\to y$. Note that $\mathcal{O}_{x}=\mathop{\boldsymbol{r}}(\mathop{\boldsymbol{s}}^{-1}(x))=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(x))$. The set of all orbits of $\mathscr{G}$ is denoted $\mathscr{G}^{(0)}/\mathscr{G}$. A subset $X$ of $\mathscr{G}^{(0)}$ is invariant if it is a union of orbits.
Suppose that $\mathscr{G}$ is an ample groupoid. Then the set $\Gamma_{c}(\mathscr{G})$ of compact open bisections form a basis for its topology. Moreover, the set $\Gamma_{c}(\mathscr{G})$ is an inverse semigroup with multiplication and inversion defined by
$UV=\{\gamma\eta\in\mathscr{G}\mid\gamma\in U,\eta\in V\},\text{ and }U^{-1}=\{\gamma^{-1}\mid\gamma\in U\},$
respectively. The semilattice of idempotents of $\Gamma_{c}(\mathscr{G})$ is given by the set of compact open subsets of $\mathscr{G}^{(0)}$.
2.2. Algebras of ample groupoids
Fix an ample groupoid $\mathscr{G}$ and a field $K$. The Steinberg algebra of $\mathscr{G}$ over $K$ is
$K\mathscr{G}=\mathrm{span}_{K}\{1_{U}\mid U\in\Gamma_{c}(\mathscr{G})\},$
with multiplication given by convolution:
$f\ast g(\gamma)=\sum_{\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{s}}\alpha)}f(\gamma\alpha^{-1})g(\alpha)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
Notice that $1_{U}*1_{V}=1_{UV}$ for all $U,V\in\Gamma_{c}(\mathscr{G})$. Thus $K\mathscr{G}$ is a quotient of the semigroup algebra $K\Gamma_{c}(\mathscr{G})$. If $K$ is a field with involution $\sigma$ (possibly trivial), then $K\mathscr{G}$ is a $\ast$-algebra with involution given by $f^{\ast}(\gamma)=\sigma(f(\gamma^{-1}))$. We shall only use this in the case $K=\mathbb{C}$ and $\sigma$ is complex conjugation. Note that $K\mathscr{G}$ is unital if and only if $\mathscr{G}^{(0)}$ is compact. See [31] for details.
If $U$ is an open invariant subset of $\mathscr{G}^{(0)}$ and $X=\mathscr{G}^{(0)}\setminus U$ is the complementary closed invariant subspace, then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$ via restriction of functions by [10, Pages 1596–1597], a fact we shall use in the sequel without comment.
2.3. Groupoid of germs and the universal groupoid of an inverse semigroup
We briefly review the construction of the universal groupoid of an inverse semigroup; see [31] or [29] for a detailed discussion.
Fix an inverse semigroup $S$ and let $E$ denote the semilattice of idempotents of $S$ (since $S$ will be fixed throughout this subsection). Let $X$ be a locally compact totally disconnected space. We denote by $I_{X}$ the inverse semigroup of all partial homeomorphisms of $X$ with compact open domain and range. An action of $S$ on $X$ is a homomorphism $\varphi\colon S\to I_{X}$ such that if $X_{e}$ denotes the domain of an idempotent $e$, then $\bigcup_{e\in E}X_{e}=X$; this last condition says that the action is non-degenerate. Notice that $\psi_{s}\colon X_{s^{*}s}\to X_{ss^{*}}$ for each $s\in S$. We put $sx=\psi_{s}(x)$ for the action when no confusion can arise.
If $S$ acts on a locally compact totally disconnected space $X$, the groupoid of germs $\mathscr{G}$ of the action is described as follows. The groupoid $\mathscr{G}$ consists of all equivalence classes of pairs $(s,x)$ with $x\in X_{s^{*}s}$ where $(s,x)\sim(t,y)$ if $x=y$ and there exists $u\in S$ with $x\in X_{u^{*}u}$ and $u\leq s,t$ (equivalently, there exists $e\in E$ with $x\in X_{e}$ and $se=te$) . The class of $(s,x)$ is denoted by $[s,x]$. A basis of compact open bisections for the topology on $\mathscr{G}$ is given by the sets
$(s,U)=\{[s,x]\mid x\in U\}$
where $U\subseteq X_{s^{*}s}$ is compact open. The set of composable pairs is $\mathscr{G}_{S}^{(2)}=\{([s,x],[t,y])\mid x=ty,\}$ with multiplication given by
$[s,ty][t,y]=[st,y],$
and inversion given by
$[s,x]^{-1}=[s^{*},sx].$
The unit space $\mathscr{G}_{S}^{(0)}=\{[e,x]\mid x\in X_{e}\}$ is identified with $X$ via the homeomorphism $[e,x]\mapsto x$. From this point of view, the domain and range maps are given by $\mathop{\boldsymbol{s}}([s,x])=x$ and $\mathop{\boldsymbol{r}}([s,x])=sx$.
There is a homomorphism $\psi\colon S\to\Gamma_{c}(\mathscr{G})$ given by $\psi(s)=(s,X_{s^{*}s})$ and $\mathscr{G}=\bigcup\psi(S)$. The reader should consult [15, 29] for more details on groupoids of germs.
Of special importance is the action of $S$ on the spectrum $\widehat{E}$ of $E$. Let $\widehat{E}$ denote the space of nonzero (semi)characters $x\colon E\to\{0,1\}$, endowed with the topology of pointwise convergence. Then $\widehat{E}$ is a locally compact, totally disconnected Hausdorff space. If $e,e_{1},\ldots e_{n}\in E$ and $x\in\widehat{E}$, then
[…]
is a compact open subset of $\widehat{E}$, and such sets form a basis $\mathscr{B}$ of compact open sets for $\widehat{E}$. Basis elements of the form
$D_{e}=\{x\in\widehat{E}\mid x(e)=1\},$
(for which $n=0$) play a particularly important role in the construction of the universal groupoid. They allow us to define an action of $S$ on $\widehat{E}$ as follows. For each $s\in S$, the map $\alpha_{s}\colon D_{s^{*}s}\to D_{ss^{*}}$ defined by
$\alpha_{s}(x)(e)=x(s^{*}es)$
is a homeomorphism, and the map $s\mapsto\alpha_{s}$ defines an action $\alpha\colon S\to I_{\widehat{E}}$ of $S$ on $\widehat{E}$ (see [29, Proposition 4.3.2] [^1] Note that Paterson works with right actions in [29].).
The universal groupoid $\mathscr{G}_{S}$ associated with $S$ is the groupoid of germs $\mathscr{G}_{S}$ of the action of $S$ on $\widehat{E}$. For $s\in S$, define $\psi(s)=(s,D_{s^{*}s})$. Then $\psi$ is a homomorphism of $S$ into $\Gamma_{c}(\mathscr{G}_{S})$[29, Theorem 3.3.2] [^2] In fact, $\psi$ is a restriction of Paterson’s $\psi_{X}$ in [29, Theorem 3.3.2]. In fact, $\psi$ is injective, and extends to an isomorphism $\widehat{\psi}\colon KS\to K\mathscr{G}_{S}$; see [31, Theorem 6.3].
If $e\in E$ and $x_{e}$ is the principal character given by
$x_{e}(f)=\begin{cases}1,&\text{if}\ f\geq e\\ 0,&\text{else.}\end{cases}$
then the isotropy group of $\mathscr{G}_{S}$ at $e$ is isomorphic to the maximal subgroup $G_{e}$. The principal characters form a dense subset of $\widehat{E}$. See [29, 31] for more details.
2.4. Regular rings
A ring $R$ is von Neumann regular, or regular for short, if for all $a\in R$, there exists $b\in R$ with $aba=a$. We do note require rings to be unital. The following facts are quite standard [19].
Proposition 2.1.
The following statements all hold.
  1. (1)
    The direct product of regular rings is regular.
  2. (2)
    If $R$ is a ring and $I$ is an ideal of $R$, then $R$ is regular if and only if $I$ and $R/I$ are regular.
  3. (3)
    A direct limit of regular rings is regular.
  4. (4)
    A corner in a regular ring is regular.
  5. (5)
    A matrix algebra over a regular ring is regular.
  6. (6)
    The center of a regular ring is regular.
  7. (7)
    Every semisimple Artinian ring is regular.
  8. (8)
    A ring $R$ is regular if and only if it has local units and each unitary $R$-module is flat.
Recall that a semigroup is locally finite if all its finitely generated subsemigroups are finite, or, equivalently, it is a direct limit of finite semigroups. Note that any subsemigroup of a locally finite semigroup is locally finite. The following is a special case of a classical theorem of Birkhoff in universal algebra, cf. [8, Theorem 10.16]. We include a proof for the sake of completeness.
Theorem 2.2 (Birkhoff).
Let $S$ be a subsemigroup of a direct product $\prod_{\alpha\in A}S_{\alpha}$ of finite semigroups of uniformly bounded order. Then $S$ is locally finite.
Proof.
Without loss of generality, we may assume that $S$ is generated by a finite set $X$. Suppose that $|S_{\alpha}|\leq d$ for all $\alpha\in A$. Let $T_{d+1}$ be the (finite) monoid of all self-maps on a $(d+1)$-element set. Let $S_{\alpha}^{1}$ be the monoid obtained by adding an identity to $S_{\alpha}$. Then the action of $S_{\alpha}$ on $S_{\alpha}^{1}$ by left multiplication is faithful and so $S_{\alpha}$ embeds in $T_{d+1}$ as $|S_{\alpha}^{1}|\leq d+1$. Hence combining the projections $\pi_{\alpha}\colon S\to S_{\alpha}$ with embeddings into $T_{d+1}$, we see that the homomorphisms from $S$ to $T_{d+1}$ separate points of $S$. But since $X$ is finite and a homomorphism from $S$ to $T_{d+1}$ is determined by its restriction to $X$, we see that $S$ has at most $r=|T_{d+1}|^{|X|}$ many distinct homomorphisms to $T_{d+1}$. Thus $S$ embeds in $T_{d+1}^{r}$ (via the product of these homomorphisms) and hence is finite. ∎
The following result is due to Connell [12].
Theorem 2.3 (Connell).
Let $G$ be a group and $R$ a ring. Then the group ring $RG$ is regular if and only if $R$ is regular, $G$ is locally finite and the order of each element of $G$ is a unit in $R$.
Attempts have been made to extend this to inverse semigroups. The best result to date is in Okniński [27, Chapter 15].
Theorem 2.4.
Let $K$ be a field and $S$ an inverse semigroup.
  1. (1)
    If $K$ has characteristic 00, then $KS$ is regular if and only if $S$ is locally finite.
  2. (2)
    If $K$ has characteristic $p>0$, $S$ is locally finite and $p$ does not divide the order of any element of any maximal subgroup of $S$, then $KS$ is regular.
  3. (3)
    If $KS$ is regular and $K$ has characteristic $p>0$, then $S$ is periodic and each maximal subgroup of $S$ is locally finite with no element of order divisible by $p$.
It is not known in positive characteristic if $KS$ regular implies that $S$ is locally finite for inverse semigroups in general, although this is known for Clifford semigroups, cf. [28, Chapter 15, Proposition 11].
Since we wish to prove the difficult implications of Theorem 2.4 as consequences of our main results, we prove the straightforward implications of Theorem 2.4, which we shall need in our arguments.
Proposition 2.5.
Let $K$ be a field and $S$ a locally finite inverse semigroup such that the characteristic of $K$ does not divide the order of any element of any maximal subgroup of $S$. Then $KS$ is regular.
Proof.
Since the class of regular rings is closed under direct limits by Proposition 2.1, we may assume that $S$ is a finite inverse semigroup and $K$ is a field whose characteristic does not divide the order of any maximal subgroup of $S$. But then $KS$ is semisimple (in fact, isomorphic to a direct product of matrix algebras over group algebras of maximal subgroups) and hence regular, cf. [32, Corollary 9.4]. ∎
3. von Neumann regularity of Steinberg algebras
The following result is proved in [4]. We provide a proof for the reader’s convenience.
Theorem 3.1.
Let $K$ be a field and $\mathscr{G}$ an ample groupoid. If $K\mathscr{G}$ is regular, then each isotropy group of $\mathscr{G}$ is locally finite and the characteristic of $K$ divides the order of no isotropy group element.
Proof.
By Connell’s theorem, it is enough to show that $K\mathscr{G}_{x}^{x}$ is regular for each $x\in\mathscr{G}^{(0)}$. Let $a\in K\mathscr{G}_{x}^{x}$, which we view as a finitely supported function $\mathscr{G}_{x}^{x}\to K$. Choose a compact open bisection $U_{\gamma}$ containing $\gamma$ for each $\gamma\in\mathscr{G}_{x}^{x}$ with $a(\gamma)\neq 0$. Consider $f=\sum_{\gamma\in\mathscr{G}^{x}_{x}}a(\gamma)U_{\gamma}$. By construction, $f|_{\mathscr{G}_{x}^{x}}=a$. Since $K\mathscr{G}$ is regular, we can find $g\in K\mathscr{G}$ with $fgf=f$. Notice that if $\alpha\in\mathscr{G}_{x}^{x}$, then
$f(\alpha)=\sum_{\beta\gamma\nu=\alpha}f(\beta)g(\gamma)f(\nu).$ (3.1)
But if $\alpha=\beta\gamma\nu$, then by construction of $f$ and the definition of a bisection if $f(\beta)\neq 0\neq f(\nu)$, we must have $\beta,\nu\in\mathscr{G}^{x}_{x}$ and hence $\gamma\in\mathscr{G}^{x}_{x}$. Therefore, (3.1) becomes
$a(\alpha)=f(\alpha)=\sum_{\beta,\gamma,\nu\in\mathscr{G}_{x}^{x},\beta\gamma\nu=\alpha}a(\beta)g(\gamma)a(\nu).$
Thus if we put $b=g|_{\mathscr{G}_{x}^{x}}$, then $b$ has finite support and $aba=a$. ∎
We aim to improve on this, particularly in characteristic zero, where we can use some analysis. We shall also see that the necessary condition in Theorem 3.1 is by no means sufficient.
Let us say that an ample groupoid $\mathscr{G}$ is uniformly bounded if there is an integer $M\geq 1$ with $|\mathop{\boldsymbol{r}}^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$. Note that any open or closed subgroupoid of a uniformly bounded ample groupoid is uniformly bounded. Let us say that $\mathscr{G}$ is approximately uniformly bounded if it is a directed union of uniformly bounded open subgroupoids. An open or closed subgroupoid of an approximately uniformly bounded ample groupoid is again approximately uniformly bounded. Let us say that $\mathscr{G}$ is approximately quasi-compact if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids. It is not difficult to verify that a principal second countable Hausdorff ample groupoid is approximately quasi-compact if and only if it is an approximately finite groupoid in the sense of [26, Section 11.5] (i.e., an approximately finite étale equivalence relation in the sense of [18]). Note that if $\mathscr{G}$ is a quasi-compact ample groupoid, then $\mathscr{G}^{(0)}$ is compact being the image of $\mathscr{G}$ under $\mathop{\boldsymbol{r}}$.
The following proposition will be used without comment.
Proposition 3.2.
Let $\mathscr{G}$ be an ample groupoid and suppose that it is a directed union of open approximately quasi-compact subgroupoids. Then $\mathscr{G}$ is approximately quasi-compact.
Proof.
Suppose that $\mathscr{G}=\bigcup_{\alpha\in D}\mathscr{G}_{\alpha}$ with the $\mathscr{G}_{\alpha}$ a directed family of open approximately quasi-compact subgroupoids. Write $\mathscr{G}_{\alpha}=\bigcup_{\beta\in I_{\alpha}}\mathscr{G}_{\alpha,\beta}$ with the $\mathscr{G}_{\alpha,\beta}$ a directed family of open quasi-compact subgroupoids of $\mathscr{G}_{\alpha}$. Then each $\mathscr{G}_{\alpha,\beta}$ is open and quasi-compact in $\mathscr{G}$ and $\mathscr{G}=\bigcup_{\alpha\in D}\bigcup_{\beta\in I_{\alpha}}\mathscr{G}_{\alpha,\beta}$. It remains to show that this family of open quasi-compact subgroupoids is directed. Given $\mathscr{G}_{\alpha,\beta}$ and $\mathscr{G}_{\alpha^{\prime},\beta^{\prime}}$, there exists $\alpha_{0}$ with $\mathscr{G}_{\alpha},\mathscr{G}_{\alpha^{\prime}}\subseteq\mathscr{G}_{\alpha_{0}}$. Since $\mathscr{G}_{\alpha,\beta}\cup\mathscr{G}_{\alpha^{\prime},\beta^{\prime}}$ is quasi-compact and $\mathscr{G}_{\alpha_{0}}=\bigcup_{\beta_{0}\in I_{\alpha_{0}}}\mathscr{G}_{\alpha_{0},\beta_{0}}$, which is a directed union of open subgroupoids, we deduce that $\mathscr{G}_{\alpha,\beta},\mathscr{G}_{\alpha^{\prime},\beta^{\prime}}\subseteq\mathscr{G}_{\alpha_{0},\beta_{0}}$ for some $\beta_{0}\in I_{\alpha_{0}}$. This establishes directedness. ∎
Our next observation is trivial.
Lemma 3.3.
Let $\mathscr{G}$ be an ample groupoid and $\mathscr{H}$ an open subgroupoid. If $\mathscr{H}$ is contained in a quasi-compact subspace of $\mathscr{G}$, then $\mathscr{H}$ is uniformly bounded. In particular, quasi-compact ample groupoids are uniformly bounded.
Proof.
Let $C$ be a quasi-compact subspace of $\mathscr{G}$ containing $\mathscr{H}$. We may cover $C$ by finitely many compact open bisections $U_{1}\cup\cdots\cup U_{n}$ of $\mathscr{G}$. Then we have $|\mathop{\boldsymbol{r}}^{-1}(x)\cap\mathscr{H}|\leq n$ for all $x\in\mathscr{H}^{(0)}$ as $|\mathop{\boldsymbol{r}}^{-1}(x)\cap U_{i}|\leq 1$ for $i=1,\ldots,n$ and $\mathscr{H}\subseteq U_{1}\cup\cdots\cup U_{n}$. ∎
The following result shows, amongst other things, that the classes of approximately quasi-compact and approximately uniformly bounded groupoids coincide.
Proposition 3.4.
Let $\mathscr{G}$ be an ample groupoid. Then the following are equivalent.
  1. (1)
    $\mathscr{G}$ is approximately quasi-compact.
  2. (2)
    $\mathscr{G}$ is approximately uniformly bounded.
  3. (3)
    The inverse semigroup $\Gamma_{c}(\mathscr{G})$ is locally finite.
  4. (4)
    There is a locally finite inverse semigroup $S\subseteq\Gamma_{c}(\mathscr{G})$ with $\bigcup S=\mathscr{G}$.
Proof.
Lemma 3.3 shows that (1) implies (2). Trivially, (3) implies (4). If $T$ is any inverse semigroup of compact open bisections of $\mathscr{G}$, then $\bigcup T$ is an open subgroupoid. Suppose that $S$ is a locally finite inverse semigroup of compact open bisections with $\bigcup S=\mathscr{G}$. Then $S$ is the directed union of its finitely generated inverse subsemigroups, each of which is finite, and hence $\mathscr{G}$ is the directed union of the groupoids of the form $\mathscr{H}=\bigcup T$ with $T$ a finite inverse subsemigroup. But since $T$ is finite and each element of $T$ is compact open, it follows that $\mathscr{H}$ is quasi-compact and open, and so (4) implies (1). It remains to show that (2) implies (3).
Suppose that (2) holds. If $\mathscr{G}$ is the directed union of uniformly bounded open subgroupoids $\mathscr{G}_{\alpha}$ with $\alpha\in D$, then $\Gamma_{c}(\mathscr{G})$ is the directed union of the $\Gamma_{c}(\mathscr{G}_{\alpha})$ since any compact set belongs to some $\mathscr{G}_{\alpha}$. Hence any finitely generated subsemigroup of $\Gamma_{c}(\mathscr{G})$ is contained in some $\Gamma_{c}(\mathscr{G}_{\alpha})$. Thus we may assume without loss of generality that $\mathscr{G}$ is uniformly bounded. Suppose that $|\mathop{\boldsymbol{r}}(x)|\leq n$ for all $x\in\mathscr{G}^{(0)}$. Then, for each $x\in\mathscr{G}^{(0)}$, we have that $|\mathscr{G}^{x}_{x}|\cdot|\mathcal{O}_{x}|=|\mathop{\boldsymbol{r}}^{-1}(x)|\leq n$. It follows that $|\mathscr{G}|_{\mathcal{O}_{x}}|=|\mathcal{O}_{x}|^{2}\cdot|\mathscr{G}_{x}^{x}|\leq n^{2}$, and hence $\mathscr{G}|_{\mathcal{O}_{x}}$ is a finite closed subgroupoid and $\Gamma_{c}(\mathscr{G}|_{\mathcal{O}_{x}})$ is a finite inverse semigroup of cardinality at most $2^{n^{2}}$.
We have an injective homomorphism $\psi\colon\Gamma_{c}(\mathscr{G})\to\prod_{\mathcal{O}\in\mathscr{G}^{(0)}/\mathscr{G}}\Gamma_{c}(\mathscr{G}|_{\mathcal{O}})$ given by $\psi(U)_{\mathcal{O}}=U\cap\mathscr{G}|_{\mathcal{O}}$. We deduce from Theorem 2.2 that $\Gamma_{c}(\mathscr{G})$ is locally finite, yielding (3). ∎
Corollary 3.5.
Let $S$ be an inverse semigroup acting on a locally compact totally disconnected Hausdorff space $X$ by partial homeomorphisms between compact open subsets. Let $\mathscr{G}$ be the groupoid of germs of the action. If $S$ is locally finite, then $\mathscr{G}$ is approximately quasi-compact. The converse holds if the mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G})$ given by $\psi(s)=(s,X_{s^{*}s})$ is injective.
Proof.
We have a homomorphism $\psi\colon S\to\Gamma_{c}(\mathscr{G})$ given by $\psi(s)=(s,X_{s^{*}s})$. Moreover, $\mathscr{G}=\bigcup\psi(S)$. If $S$ is locally finite, then so is $\psi(S)$ and hence $\mathscr{G}$ is approximately quasi-compact by Proposition 3.4. Conversely, if $\mathscr{G}$ is approximately quasi-compact and $\psi$ is injective, then $\Gamma_{c}(\mathscr{G})$ and all its subsemigroups are locally finite by Proposition 3.4, and hence $S\cong\psi(S)$ is locally finite. ∎
To prove our main result, let us first recall the construction of the reduced $C^{*}$-algebra of an ample groupoid as described in [11]. Let $\mathscr{G}$ be an ample groupoid. Then $\mathbb{C}\mathscr{G}$ is a $\ast$-algebra with involution $f^{\ast}(\gamma)=\overline{f(\gamma^{-1})}$. For each $x\in\mathscr{G}^{(0)}$, we consider a $\ast$-representation $\pi_{x}\colon\mathbb{C}\mathscr{G}\to\mathcal{B}(\ell_{2}(\mathscr{G}_{x}))$ given on the orthonormal basis $\delta_{\gamma}$, $\gamma\in\mathscr{G}_{x}$, by
$\pi_{x}(f)\delta_{\gamma}=\sum_{\alpha\in\mathop{\boldsymbol{s}}^{-1}(\mathop{\boldsymbol{r}}(\gamma))}f(\alpha)\delta_{\alpha\gamma}.$
Notice that if $U$ is a compact open bisection, then $\|\pi_{x}(1_{U})\|\leq 1$. Indeed, $U$ induces a partial permutation $\psi_{U}$ of $\mathscr{G}_{x}$ by observing that if $\gamma\in\mathscr{G}_{x}$, then either $U\{\gamma\}$ is empty, in which case $\psi_{U}(\gamma)$ is undefined, or $U\{\gamma\}=\{\alpha\gamma\}$ where $\alpha$ is the unique element of $U$ with $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\gamma)$, in which case $\psi_{U}(\gamma)=\alpha\gamma$. Therefore if $a=\sum_{\gamma\in\mathscr{G}_{x}}c_{\gamma}\delta_{\gamma}\in\ell_{2}(\mathscr{G}_{x})$, then writing $\mathrm{dom}(\psi_{U})$ for the domain of $\psi_{U}$, we have that
$\pi_{x}(1_{U})a=\sum_{\gamma\in\mathscr{G}_{x}}\sum_{\alpha\in\mathop{\boldsymbol{r}}^{-1}(\gamma)}1_{U}(\alpha)c_{\gamma}\delta_{\alpha\gamma}=\sum_{\gamma\in\mathrm{dom}(\psi_{U})}c_{\gamma}\delta_{\psi_{U}(\gamma)}$
and so $\|\pi_{x}(1_{U})a\|\leq\|a\|$. Thus $\|\pi_{x}(1_{U})\|\leq 1$. It follows that if $f=\sum_{i=1}^{n}c_{i}1_{U_{i}}\in\mathbb{C}\mathscr{G}$ with the $U_{i}$ compact open bisections, then $\|\pi_{x}(f)\|\leq\sum_{i=1}^{n}|c_{i}|$, and so $\|\pi_{x}(f)\|\leq|f|$ where ||$\cdot$|| ⋅ | is the norm defined in [33] by
$|f|=\inf\left\{\sum_{i=1}^{n}|d_{i}|:f=\sum_{i=1}^{n}d_{i}1_{U_{i}},U_{i}\in\Gamma_{c}(\mathscr{G})\right\}.$
(Note that this latter norm also bounds the $I$-norm.) The reduced $C^{*}$-norm on $\mathbb{C}\mathscr{G}$ is defined by $\|f\|_{r}=\sup_{x\in\mathscr{G}^{(0)}}\|\pi_{x}(f)\|$. The reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$ of $\mathscr{G}$ is then the completion of $\mathbb{C}\mathscr{G}$ with respect to $\|\cdot\|_{r}$; see [11] for details. The fact that the reduced norm is truly a norm, and hence $\mathbb{C}\mathscr{G}$ embeds in $C^{*}_{r}(\mathscr{G})$, follows from the following estimate that we shall need in any event.
Proposition 3.6.
If $f\in\mathbb{C}\mathscr{G}$, then $\|f\|_{\infty}\leq\|f\|_{r}\leq|f|$.
Proof.
We already observed that $\|f\|_{r}\leq|f|$. If $f\in\mathbb{C}\mathscr{G}$ and $\gamma\in\mathscr{G}$, then for $x=\mathop{\boldsymbol{s}}(\gamma)$, we have $\langle\pi_{x}(f)\delta_{x},\delta_{\gamma}\rangle=f(\gamma)$, and so by the Cauchy-Schwarz inequality $|f(\gamma)|\leq\|\pi_{x}(f)\|\leq\|f\|_{r}$. ∎
We now aim to characterize ample groupoids with regular algebras over fields of characteristic zero, generalizing Theorem 2.4. Our proof is inspired by Okniński [28, Chapter 13, Lemma 2], who traces his approach to an idea of Formanek. We shall need the well-known fact that any countable field $L$ of characteristic 00 can be embedded into $\mathbb{C}$. Indeed, choose a transcendence basis $S$ for $L/\mathbb{Q}$. Then $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{Q}(S)/\mathbb{Q}$ is purely transcendental. Since $L$ is countable, $S$ is also countable. On the other hand, $\mathbb{C}$ is uncountable and hence has an uncountable transcendence basis $T$ over $\mathbb{Q}$. Thus we can find an injective map $S\to T$ which extends to a field embedding $\mathbb{Q}(S)\hookrightarrow\mathbb{C}$. Since $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{C}$ is algebraically closed, we can then extend this embedding to $L$.
Theorem 3.7.
Let $K$ be a field of characteristic zero and $\mathscr{G}$ an ample groupoid. Then the following are equivalent.
  1. (1)
    $K\mathscr{G}$ is regular.
  2. (2)
    $\mathscr{G}$ is approximately uniformly bounded.
  3. (3)
    $\mathscr{G}$ is approximately quasi-compact.
  4. (4)
    $\Gamma_{c}(\mathscr{G})$ is locally finite.
  5. (5)
    There is a locally finite inverse semigroup $S\subseteq\Gamma_{c}(\mathscr{G})$ with $\bigcup S=\mathscr{G}$.
Proof.
We already know that (2)–(5) are equivalent by Proposition 3.4. Let $S=\Gamma_{c}(\mathscr{G})$ be the inverse semigroup of compact open bisections of $\mathscr{G}$. Note that (4) implies that $KS$ is regular by Proposition 2.5, and hence $K\mathscr{G}$ is regular by Proposition 2.1, as it is quotient of $KS$. Thus (4) implies (1). It remains to show that (1) implies (2).
Note that if $U$ is any compact open subset of $\mathscr{G}^{(0)}$, then $U\mathscr{G}U=\mathscr{G}|_{U}$ is an open subgroupoid with compact unit space $U$ and $K\mathscr{G}|_{U}=1_{U}K\mathscr{G}1_{U}$, and hence is regular by Proposition 2.1. Since $\mathscr{G}$ is the directed union of the $\mathscr{G}|_{U}$ as $U$ runs over all compact open subsets of $\mathscr{G}^{(0)}$, we may assume without loss of generality that $\mathscr{G}^{(0)}$ is compact, and hence $K\mathscr{G}$ is unital, by Propositions 3.2 and 3.4.
Note that if $T$ is any inverse subsemigroup of $S$, then $\bigcup T$ is an open subgroupoid of $\mathscr{G}$. Also, we have that $\mathscr{G}=\bigcup S$. Since $S$ is a directed union of its finitely generated inverse subsemigroups, it suffices by Proposition 3.4 to prove that $\bigcup T$ is uniformly bounded for every finitely generated inverse subsemigroup $T$ of $S$. So let $X\subseteq S$ be a finite subset closed under inversion. We will show that $\mathscr{H}=\bigcup\langle X\rangle$ is uniformly bounded.
Say $X=\{U_{1},\ldots,U_{n}\}$ and let $a=\frac{1}{2n}(1_{U_{1}}+\cdots+1_{U_{n}})\in K\mathscr{G}$. Then since $K\mathscr{G}$ is regular, we can find $b\in K\mathscr{G}$ such that $(1-a)b(1-a)=(1-a)$. We can write $b=c_{1}1_{V_{1}}+\cdots+c_{k}1_{V_{k}}$ with $V_{1},\ldots,V_{k}$ compact open bisections. Then $L=\mathbb{Q}(c_{1},\ldots,c_{k})$ is a finitely generated extension of $\mathbb{Q}$ such that $a,b\in L\mathscr{G}$. Since $L$ is countable, we can embed it over $\mathbb{Q}$ into $\mathbb{C}$ and hence $a,b\in\mathbb{C}\mathscr{G}$. Note that by Proposition 3.6, we have that $\|a\|_{r}\leq|a|\leq\frac{1}{2}$, and so $\sum_{m=0}^{\infty}a^{m}$ converges to $(1-a)^{-1}$ in $C^{*}_{r}(\mathscr{G})$. But in $\mathbb{C}\mathscr{G}\subseteq C^{*}_{r}(\mathscr{G})$, we have $(1-a)b(1-a)=(1-a)$, and so we deduce that $b=(1-a)^{-1}$ by cancelling $1-a$ on both sides in $C^{*}_{r}(\mathscr{G})$. Thus $b=\sum_{m=0}^{\infty}a^{m}$ in $C^{*}_{r}(\mathscr{G})$, that is, $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{r}=0$. But then since $\|f\|_{\infty}\leq\|f\|_{r}$, for $f\in\mathbb{C}\mathscr{G}$ by Proposition 3.6, we have $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{\infty}=0$. From the definition of $a$, we see that
$a^{m}=\frac{1}{(2n)^{m}}\sum_{U_{i_{1}},\ldots,U_{i_{m}}\in X}1_{U_{i_{1}}\cdots U_{i_{m}}}.$
Now if $\gamma\in\mathscr{H}=\bigcup\langle X\rangle$, then $\gamma\in U_{i_{1}}\cdots U_{i_{r}}$ for some $U_{i_{1}},\ldots,U_{i_{r}}\in X$ (not necessarily distinct). But then $\sum_{m=0}^{N}a^{m}(\gamma)\geq a^{r}(\gamma)\geq\frac{1}{(2n)^{r}}$ for any $N\geq r$. It follows that $b(\gamma)\neq 0$ for all $\gamma\in\mathscr{H}$. Thus $\mathscr{H}$ is contained in the quasi-compact subspace $V_{1}\cup\cdots\cup V_{k}$ of $\mathscr{G}$. It follows from Lemma 3.3 that $\mathscr{H}$ is uniformly bounded, completing the proof. ∎
Remark 3.8.
Note that Theorem 3.7 gives us some additional information about the structure of $K\mathscr{G}$ when it is regular in characteristic 00. Namely, $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple subalgebras. In particular, if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial (i.e., a direct limit of finite direct products of matrix algebras over $K$). Indeed, if $\Gamma_{c}(\mathscr{G})$ is locally finite, then since $KT$ is a finite dimensional semisimple $K$-algebra for any finite inverse subsemigroup of $T$ [32, Chapter 9], we see that $K\Gamma_{c}(\mathscr{G})$ is of the desired form. Since a quotient algebra of such a directed union is again such a directed union, it follows that $K\mathscr{G}$ is of this form.
In an attempt to understand the positive characteristic case, we will give sufficient conditions for the algebra of an approximately quasi-compact ample groupoid to be regular.
Proposition 3.9.
Let $\mathscr{G}$ be a quasi-compact ample group bundle and $K$ a field whose characteristic does not divide the order of any isotropy group element. Then $K\mathscr{G}$ is regular.
Proof.
We know that $S=\Gamma_{c}(\mathscr{G})$ is locally finite by Proposition 3.4. We show that no maximal subgroup of $S$ contains an element of order divisible by the characteristic of $K$. It will then follow that $KS$ is regular by Proposition 2.5, and hence its quotient $K\mathscr{G}$ is regular by Proposition 2.1. Let $U\subseteq\mathscr{G}^{(0)}$ be the identity of a maximal subgroup $H$, and note that $H$ consists of those $V\in S$ with $\mathop{\boldsymbol{s}}(V)=U=\mathop{\boldsymbol{r}}(V)$. If $V\in H$ and $x\in U$, then denote by $v_{x}$ the unique element of $V$ with $\mathop{\boldsymbol{r}}(v_{x})=x$. Since $\mathscr{G}$ consists of isotropy, $\mathop{\boldsymbol{s}}(v_{x})=x$ as well. Then we have an injective group homomorphism $\psi\colon H\to\prod_{x\in U}\mathscr{G}_{x}^{x}$ given by $\psi(V)_{x}=v_{x}$. Suppose that $K$ has characteristic $p>0$. Note that if $V\in H$ satisfies $V^{p}=U$, then $\psi(V)^{p}=\psi(U)$ and hence $v_{x}^{p}=x$ for all $x\in U$. Since no isotropy group has an element of order $p$, we deduce that $v_{x}=x$ for all $x\in U$, and hence $V=U$. Thus no element of $H$ has order divisible by $p$. This concludes the proof. ∎
We now consider a class of quasi-compact ample groupoids whose members are building blocks of more complicated ones. The remaining proofs are very much inspired by [18, Lemma 3.4].
Lemma 3.10.
Let $\mathscr{G}$ be a quasi-compact ample groupoid and $K$ a field whose characteristic does not divide the order of any isotropy group element. Suppose that all orbits of $\mathscr{G}$ have the same cardinality. Then $K\mathscr{G}$ is regular.
Proof.
It follows from Lemma 3.3 that the size of each orbit is finite, say that this common size is $n$. Choose for each $x\in\mathscr{G}^{(0)}$, elements $\gamma_{x,1},\ldots,\gamma_{x,n}$ with $\mathop{\boldsymbol{s}}(\gamma_{x,i})=x$ and $\mathcal{O}_{x}=\{\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})\}$. We may assume without loss of generality that $\gamma_{x,1}=x$ (the identity at $x$). Choose for each $i$ a compact open bisection $U^{x}_{i,1}$ containing $\gamma_{x,i}$. We may assume that $U^{x}_{1,1}\subseteq\mathscr{G}^{(0)}$. Since $\mathscr{G}^{(0)}$ is Hausdorff with a basis of compact open sets and $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x_{n}})$ are pairwise distinct, we can find pairwise disjoint compact open neighborhoods $V_{1},\ldots,V_{n}$ of $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})$, respectively, in $\mathscr{G}^{(0)}$. Replacing $U^{x}_{i,1}$ by $V_{i}U^{x}_{i,1}$, we may assume that the sets $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,1})$ are pairwise disjoint. Also, putting $U=\bigcap_{i=1}^{n}\mathop{\boldsymbol{s}}(U^{x}_{i,1})$ and replacing $U^{x}_{i,1}$ by $U^{x}_{i,1}U$, we may assume that $\mathop{\boldsymbol{s}}(U^{x}_{i,1})=\mathop{\boldsymbol{s}}(U^{x}_{1,1})=U^{x}_{1}$ for all $i$. Put $U^{x}_{i,j}=U^{x}_{i,1}(U^{x}_{j,1})^{-1}$. Note that $\mathop{\boldsymbol{s}}(U^{x}_{i,j})=U^{x}_{j}$ and $\mathop{\boldsymbol{r}}(U^{x}_{i,j})=U^{x}_{i}$ for all $i,j$. Also note that $U^{x}_{i,i}=U^{x}_{i}$ for $i=1,\ldots,n$. Let $U^{x}=\bigcup_{i=1}^{n}U^{x}_{i}$. Then $U^{x}$ is compact open and we claim that it is invariant. For suppose that $y\in U^{x}_{j}$. Then since $U^{x}_{j}=\mathop{\boldsymbol{s}}(U^{x}_{i,j})$ and $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,j})$ for all $i=1,\ldots,n$, we can find elements of the orbit of $y$ in each of $U^{x}_{1},\ldots,U^{x}_{n}$. Since these sets are pairwise disjoint, this yields $n$ elements in the orbit of $y$ and hence the whole orbit by assumption. Thus each $U^{x}$ is invariant and each $U^{x}_{i}$ contains exactly one element from each orbit of $U^{x}$. Also note that
$U^{x}_{i,j}U^{x}_{k,\ell}=\begin{cases}U^{x}_{i,\ell},&\text{if}\ j=k\\ \emptyset,&\text{otherwise.}\end{cases}$ (3.2)
Another observation is that the isotropy subgroupoid $\mathop{\mathrm{Iso}}(\mathscr{G})$ of $\mathscr{G}$ is clopen, and hence quasi-compact. It is closed since $\mathscr{G}^{(0)}$ is Hausdorff and it is the equalizer of $\mathop{\boldsymbol{s}},\mathop{\boldsymbol{r}}$. It is open because if $\gamma\colon x\to x$ is isotropy and $V$ is a compact open bisection containing $\gamma$, then $U^{x}_{1}VU^{x}_{1}$ is a compact open bisection containing $\gamma$ and consisting of isotropy as $x\in U^{x}_{1}$, and $U^{x}_{1}$ contains exactly only one element from each orbit of the invariant compact open set $U^{x}$, whence $\mathop{\boldsymbol{s}}(\alpha),\mathop{\boldsymbol{r}}(\alpha)\in U^{x}_{1}$ implies $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\alpha)$.
Since $\mathscr{G}^{(0)}$ is compact and the $U^{x}$ cover $\mathscr{G}^{(0)}$, we can find finitely may points $x_{1},\ldots,x_{r}$ such that $\mathscr{G}^{(0)}=U^{x_{1}}\cup\cdots\cup U^{x_{r}}$. Without loss of generality, we may assume that $U^{x_{i}}\nsubseteq\bigcup_{j\neq i}U^{x_{j}}$ for all $i$. Put $V_{1}=U^{x_{1}},V_{2}=U^{x_{2}}\setminus V_{1},\ldots,V_{r}=U^{x_{r}}\setminus\bigcup_{i=1}^{r-1}V_{i}$. Then $V_{1},\ldots,V_{r}$ is partition of $\mathscr{G}^{(0)}$ into compact open invariant sets. In particular note that $V_{i}U=UV_{i}$ for any compact open bisection $U$. It follows that if $\mathscr{G}_{i}=\mathscr{G}|_{V_{i}}$, then $K\mathscr{G}\cong\prod_{k=1}^{r}K\mathscr{G}_{k}$ and $K\mathscr{G}_{k}=1_{V_{k}}K\mathscr{G}1_{V_{k}}$.
Note that since $\mathscr{G}_{k}=\mathop{\boldsymbol{r}}^{-1}(V_{k})$ is clopen in $\mathscr{G}$ it is quasi-compact. Also $\mathop{\mathrm{Iso}}(\mathscr{G}_{k})=\mathop{\mathrm{Iso}}(\mathscr{G})\cap\mathscr{G}_{k}$ is clopen and hence quasi-compact. Put $U^{k}_{i,j}=V_{k}U^{x_{k}}_{i,j}=U^{x_{k}}_{i,j}V_{k}$; this is a compact open bisection in $\mathscr{G}_{k}$. Also note that $U^{k}_{1,1}\subseteq U^{x_{k}}_{1}$ does not contain any two points in the same orbit by the previous discussion. Hence $U^{k}_{1,1}\mathscr{G}_{k}U^{k}_{1,1}=\mathscr{G}_{k}|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})\cap\mathop{\boldsymbol{r}}^{-1}(U^{k}_{1,1})$ is clopen (as $U^{k}_{1,1}$ is a clopen subspace of $\mathscr{G}^{(0)}$) and thus a quasi-compact group bundle. Note that $V_{k}=V_{k}U^{x_{k}}=\bigcup_{i=1}^{n}U^{k}_{i,i}$ and hence it follows from (3.2) that the $1_{U^{k}_{i,j}}$ form a set of matrix units for $K\mathscr{G}_{k}$. Thus $K\mathscr{G}_{k}\cong M_{n}(1_{U^{k}_{1,1}}K\mathscr{G}_{k}1_{U^{k}_{1,1}})=M_{n}(K\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}})$. We deduce that $K\mathscr{G}$ is regular from Propositions 2.1 and 3.9. ∎
We need one last lemma before proving our next main result.
Lemma 3.11.
Let $\mathscr{G}$ be an étale groupoid and $k\geq 1$ be a natural number. Then the set $O_{k}$ of elements $x\in\mathscr{G}^{(0)}$ whose orbit has at least $k$ elements is an open invariant subset.
Proof.
Clearly $O_{k}$ is invariant; we claim that it is open. Suppose that $x\in O_{k}$. Let $x=y_{1},y_{2}\ldots,y_{k}$ be $k$ distinct elements in the orbit of $x$. As in the proof of Lemma 3.10, we can find open bisections $U_{1},\ldots,U_{k}$ with $x\in\mathop{\boldsymbol{s}}(U_{i})$, $y_{i}\in\mathop{\boldsymbol{r}}(U_{i})$, for $i=1,\ldots,k$, and $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ pairwise disjoint (using that $\mathscr{G}^{(0)}$ is Hausdorff). Then $U=\bigcap_{i=1}^{k}\mathop{\boldsymbol{s}}(U_{i})$ is a neighborhood of $x$ and if $y\in U$, then each of $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ contains an element of the orbit of $y$. Since these sets are pairwise disjoint, we conclude that $y\in O_{k}$. ∎
We can now prove our second main theorem.
Theorem 3.12.
Suppose that $K$ is a field and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that the order of no isotropy element is divisible by the characteristic of $K$. Then $K\mathscr{G}$ is regular.
Proof.
Since an approximately quasi-compact ample groupoid is a directed union of quasi-compact open subgroupoids, it suffices to prove the result for quasi-compact groupoids as regular rings are closed under direct limits. So assume that $\mathscr{G}$ is quasi-compact and hence uniformly bounded by Lemma 3.3. In particular, the cardinalities of the orbits of $\mathscr{G}$ are uniformly bounded.
Let $d_{1}<d_{2}<\cdots<d_{m}$ be the distinct orbit sizes. We prove the result by induction on $m$. If $m=1$, then all orbits of $\mathscr{G}$ have the same size and the result follows from Lemma 3.10. Let $U=\{x\in\mathscr{G}^{(0)}\mid|\mathcal{O}_{x}|=d_{m}\}$. Then $U$ is open and invariant by Lemma 3.11 once we note that $U=O_{d_{m}}$ in the notation of that lemma. Let $X=\mathscr{G}^{(0)}\setminus U$. Then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$. Note that $\mathscr{G}|_{U}=\mathop{\boldsymbol{r}}^{-1}(U)$ is open and $\mathscr{G}|_{X}=\mathop{\boldsymbol{r}}^{-1}(X)$ is closed and hence quasi-compact. By construction, the orbits sizes in $\mathscr{G}|_{X}$ are $d_{1}<d_{2}<\cdots<d_{m-1}$, and so by induction $K\mathscr{G}|_{X}$ is regular. If we can show that $K\mathscr{G}|_{U}$ is regular, then Proposition 2.1 will imply that $K\mathscr{G}$ is regular.
Let $V\subseteq U$ be a compact open set. Then $V$ is clopen in $\mathscr{G}^{(0)}$ and hence $\mathop{\boldsymbol{r}}^{-1}(V)\subseteq\mathscr{G}$ is clopen, hence quasi-compact and open. Thus $W=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(V))$ is a compact open invariant subset of $U$ containing $V$. Hence the compact open invariant subsets of $U$ are cofinal among compact open subsets of $U$. Moreover, if $W$ is a compact open invariant subspace of $U$, then $W$ is clopen in $\mathscr{G}^{(0)}$ (as the latter is Hausdorff), and so $\mathscr{G}|_{W}=\mathop{\boldsymbol{r}}^{-1}(W)$ is a clopen subgroupoid of $\mathscr{G}$, hence quasi-compact, with all orbits of size $d_{m}$ (as $W\subseteq U$ is invariant). Hence $K\mathscr{G}|_{W}$ is regular by Lemma 3.10. Since $K\mathscr{G}|_{U}$ is the direct limit of the $K\mathscr{G}|_{W}$ where $W$ runs over the compact open invariant subsets of $U$, we conclude that $K\mathscr{G}|_{U}$ is regular from Proposition 2.1. This completes the proof. ∎
Remark 3.13.
If one follows the proofs carefully, it is not difficult to see that if $K$ is a field of characteristic $p>0$ and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that no isotropy element has order divisible by $p$, then $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple algebras, and hence if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial. This just uses that the class under consideration is closed under matrix amplification and if an algebra has a finite ideal filtration with factors of this form, then the algebra has this form.
4. Graded von Neumann regularity of Steinberg algebras
4.1. Graded regularity
In this section we characterize when ample groupoid algebras are graded von Neumann regular.
Let $K$ be a field, $R$ a $K$-algebra and $H$ a group with identity $e$. Then $R$ is $H$-graded if $R$ has a direct sum decomposition $R=\bigoplus_{h\in H}R_{h}$, where each $R_{h}$ is a $K$-vector subspace of $R$ such that $R_{h}R_{k}\subseteq R_{hk}$. The summand $R_{h}$ is called the homogeneous component of degree $h$. The homogeneous component $R_{e}$ is a $K$-subalgebra of $R$. An element is homogenous if it belongs to a homogeneous component.
A graded ring $R$ is graded (von Neumann) regular if for every homogeneous element $x\in R$ there exists $y\in R$ such that $x=xyx$.
Let $\mathscr{G}$ be an ample groupoid and let $H$ be a discrete group. A cocycle from $\mathscr{G}$ to $H$ is a map $c\colon\mathscr{G}\to H$ such that $c(\gamma^{-1})=c(\gamma)^{-1}$, for all $\gamma\in\mathscr{G}$, and $c(\gamma\eta)=c(\gamma)c(\eta)$ for all $(\gamma,\eta)\in\mathscr{G}^{(2)}$. A continuous cocycle $c\colon\mathscr{G}\to H$ induces an an $H$-grading $\{(K\mathscr{G})_{h}\}_{h}$ of $K\mathscr{G}$. This is done in detail in [9] for Hausdorff groupoids. The details for non-Hausdorff groupoids are similar, but we present them for completeness. Call a compact open bisection $U\in\Gamma_{c}(\mathscr{G})$ homogenous of degree $h\in H$ if $U\subseteq c^{-1}(h)$ for some $h\in H$. Then $(K\mathscr{G})_{h}$ is the $K$-span of the characteristic functions of homogeneous compact open bisections of degree $h$. Note that each $c^{-1}(h)$ with $h\in H$ is clopen, and so if $U$ is any compact open bisection, then $U=\bigcup_{h\in H}(U\cap c^{-1}(h))$ and the $U\cap c^{-1}(h)$ are pairwise disjoint homogeneuous compact open bisections. By compactness of $U$, only finitely many of these intersections, say corresponding to $h_{1},\ldots,h_{n}$, are nonempty, and so $1_{U}=\sum_{i=1}^{n}1_{(c^{-1}(h_{i})\cap U)}$, showing that $\sum_{h\in H}(K\mathscr{G})_{h}=K\mathscr{G}$. The sum is direct since the $c^{-1}(h)$ are pairwise disjoint and elements of $(K\mathscr{G})_{h}$ are supported on $c^{-1}(h)$. One can, in fact, show that $(K\mathscr{G})_{h}$ consists of those elements of $K\mathscr{G}$ supported on $c^{-1}(h)$, but we shall never need this description. This argument also shows that the homogeneous compact open bisections form a basis for the topology of $\mathscr{G}$.
Notice that $\mathscr{G}^{(0)}\subseteq c^{-1}(e)=\mathscr{G}_{e}$ and that $\mathscr{G}_{e}$ is a clopen subgroupoid and hence is ample. Moreover, $(K\mathscr{G})_{e}=K\mathscr{G}_{e}$ by definition. From now on, we write $K\mathscr{G}_{h}$ instead of the more cumbersome $(K\mathscr{G})_{h}$.
The following is [21, Definition 2.9]. Let $R=\bigoplus_{h\in H}R_{h}$ be an $H$-graded $K$-algebra. If, for every $h\in H$ and $r\in R_{h}$, there exist $\epsilon_{h}(r)\in R_{h}R_{h^{-1}}$ and $\epsilon_{h}^{\prime}(r)\in R_{h^{-1}}R_{h}$ such that $\epsilon_{g}(r)r=r=r\epsilon_{g}^{\prime}(r)$, then $R$ is called nearly epsilon-strongly $H$-graded. It is shown in [21] that being nearly epsilon-strongly $H$-graded is a necessary condition for graded regularity.
Proposition 4.1.
Let $\mathscr{G}$ be an ample groupoid with a continuous cocycle $c\colon\mathscr{G}\to H$ and let $K$ be a field. Then $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
Proof.
Let $\mathscr{G}=\bigoplus_{h\in H}K\mathscr{G}_{h}$ be the $H$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to H$ as above.
Fix any $h\in H$ and let $U\subseteq c^{-1}(h)$ be a compact open bisection. Define $\epsilon_{h}(1_{U})=1_{UU^{-1}}$. Then
$\epsilon_{h}(1_{U})=1_{UU^{-1}}=1_{U}1_{U^{-1}}\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}},$
and
$\epsilon_{h}(1_{U})1_{U}=1_{UU^{-1}}1_{U}=1_{UU^{-1}U}=1_{U}.$
Similarly, defining $\epsilon^{\prime}_{h}(1_{U})=1_{U^{-1}U}$ yields that $\epsilon^{\prime}_{h}(1_{U})\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$ and $1_{U}\epsilon^{\prime}_{h}(1_{U})=1_{U}$.
Now consider an arbitrary $f\in K\mathscr{G}_{h}$. Then there are compact open bisections $U_{1},U_{2},\ldots,U_{n}$, with each $U_{i}\subseteq c^{-1}(h)$, and $k_{1},k_{2},\cdots,k_{n}\in K$ such that $f=\sum_{i=1}^{n}k_{i}1_{U_{i}}$. In particular, $\epsilon_{h}(1_{U_{i}})\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$ for each $i$. Put $U=\bigcup_{i=1}^{n}U_{i}U_{i}^{-1}$. Then $U$ is a compact open set in $\mathscr{G}^{(0)}$. Put $\epsilon_{h}(f)=1_{U}$. Since $U\subseteq\mathscr{G}^{(0)}$ and $\mathscr{G}^{(0)}$ is Hausdorff, it follows from the inclusion-exclusion principle that $1_{U}=\epsilon_{h}(f)$ is an integer combination of the $1_{U_{i}U_{i}^{-1}}$, and hence $\epsilon_{h}(f)\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$. Similarly, putting $U^{\prime}=\bigcup_{i=1}^{n}U_{i}^{-1}U_{i}$ and defining $\epsilon^{\prime}_{h}(f)=1_{U^{\prime}}$, we have that $\epsilon^{\prime}_{h}(f)\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$. It is straightforward from the construction that $\epsilon_{h}(f)f=f=f\epsilon^{\prime}_{h}(f)$. Hence, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded. ∎
Theorem 4.2.
Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle. Then $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular.
Proof.
By Proposition 4.1, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded. Thus, by [21, Theorem 1.2], $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular. ∎
Theorem 4.2 allows us to upgrade Theorems 3.7 and 3.12 to a theorem about graded regularity.
Theorem 4.3.
Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle.
  1. (1)
    If $\mathscr{G}_{e}$ is approximately quasi-compact and the characteristic of $K$ divides the order of no isotropy element of $\mathscr{G}_{e}$, then $K\mathscr{G}$ is graded regular with respect to the $H$-grading induced by $c$.
  2. (2)
    If $K$ is of characteristic 00, then $K\mathscr{G}$ is graded regular (with respect to the $H$-grading induced by $c$) if and only if $\mathscr{G}_{e}$ is approximately quasi-compact.
5. Applications
Our first application is to recover the nontrivial part of Okniński’s result on regularity of inverse semigroups in characteristic zero from Theorem 3.7.
Theorem 5.1 (Okniński).
Let $S$ be an inverse semigroup and $K$ a field of characteristic zero. Then $KS$ is regular if and only if $S$ is locally finite.
Proof.
We already proved in Proposition 2.5 that if $S$ is locally finite, then $KS$ is regular. For the converse, assume that $KS$ is regular. Let $\mathscr{G}_{S}$ denote the universal groupoid of $S$. Since $KS\cong K\mathscr{G}_{S}$, we have that $K\mathscr{G}_{S}$ is regular. Then, by Theorem 3.7, we have that $\mathscr{G}_{S}$ is approximately quasi-compact, and hence $S$ is locally finite by Corollary 3.5 since the canonical mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G}_{S})$ given by $\psi(s)=(s,D_{s^{*}s})$ is injective. ∎
Next we give an example to show that the necessary conditions in Theorem 3.1 are not sufficient to guarantee regularity of the algebra of a groupoid by giving examples of principal groupoids with nonregular algebras. Recall that if a discrete group $G$ acts on a locally compact, Hausdorff and totally disconnected space $X$ by homeomorphisms, then the groupoid of germs of this action is just the so-called transformation groupoid $\mathscr{G}$ with underlying space $G\times X$ (with the product topology) and unit space $\{1\}\times X$ (identified with $X$) with $\mathop{\boldsymbol{s}}(g,x)=x$ and $\mathop{\boldsymbol{r}}(g,x)=gx$. Products are given by $(h,gx)(g,x)=(hg,x)$ and $(g,x)^{-1}=(g^{-1},gx)$. The isotropy group of the transformation groupoid at $x\in X$ is just stabilizer of $x$ in $G$. In particular, if the action of $G$ is free, then the transformation groupoid is principal. It is well known that $K\mathscr{G}$ is isomorphic to the skew group ring $C_{c}(X,K)\rtimes G$ where $C_{c}(X,K)$ is the ring of compactly supported locally constant functions $X\to K$ and $G$ acts on this ring via $(gf)(x)=f(g^{-1}x)$. The first part of the following result extends a result in [18].
Theorem 5.2.
Let $X$ be compact totally disconnected space and $G$ a discrete group acting on $X$. Let $\mathscr{G}$ be the corresponding transformation groupoid.
  1. (1)
    $\mathscr{G}$ is approximately quasi-compact if and only if $G$ is locally finite.
  2. (2)
    If $K$ is a field of characteristic 00, then $K\mathscr{G}$ is regular if and only if $G$ is locally finite.
Proof.
The first item follows from Corollary 3.5 since the map $\psi\colon G\to\Gamma_{c}(\mathscr{G})$ given by $\psi(g)=\{g\}\times X$ is injective. The second item follows from the first and Theorem 3.7. ∎
For example, if we consider a minimal (nonperiodic) action of $\mathbb{Z}$ on the Cantor set $X$ (e.g., a nonperiodic minimal shift space), then the associated transformation groupoid $\mathscr{G}$ is a principal groupoid (and hence trivially satisfies the necessary conditions in Theorem 3.1), but $\mathbb{C}\mathscr{G}$ is not regular.
5.1. Leavitt path algebras
We follow the conventions and notation of [1] for Leavitt path algebras. For the boundary path groupoid of a graph we use the terminology and conventions of [26].
A directed graph $E=(E^{0},E^{1},r,s)$ consists of a set of vertices $E^{0}$, a set of edges $E^{1}$ and maps $r,s\colon E^{1}\to E^{0}$. A finite path in $E$ is a finite sequence of edges $\mu=e_{1}e_{2}\cdots e_{n}$ such that $r(e_{i})=s(e_{i+1})$. The length of $\mu$ is $|\mu|=n$. We also admit an empty path of length 00 at each vertex that we often identify with the vertex. An infinite path is an infinite sequence of edges $e_{1}e_{2}\cdots$ such that $r(e_{i})=s(e_{i+1})$ for all $i\geq 1$. Let $E^{*}$ denote the set of all finite paths in $E$ and let $E^{\infty}$ denote the set of all infinite paths in $E$. A finite path $e_{1}e_{2}\cdots e_{n}$ is a closed path if $s(e_{1})=r(e_{n})$, and $E$ is acyclic if it has no nonempty closed paths. A vertex $v\in E^{0}$ is a sink if $|s^{-1}(v)|=0$, an infinite emitter if $|s^{-1}(v)|=\infty$ and a regular vertex otherwise. Let $E_{\text{reg}}$ denote all regular vertices, and call $E_{\text{sing}}=E^{0}\setminus E_{\text{reg}}$ the set of singular vertices. The mappings $r,s$ extend to paths in the natural way.
The boundary path space of $E$ is the set $\partial E=E^{\infty}\cup\{\mu\in E^{*}\mid r(\mu)\in E_{\text{sing}}\}$. If $\mu\in E^{*}$, we let $Z(\mu)=\{\mu x\mid r(\mu)=s(x),x\in\partial E\}$. Sets of the form
$Z(\mu\setminus F)=Z(\mu)\setminus\left(\cup_{e\in F}Z(\mu e)\right),$
where $\mu\in E^{*}$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, form a basis of compact open sets for a locally compact Hausdorff topology on $\partial E$.
Let $\partial E^{\geq n}=\{\mu\in\partial E\mid|\mu|\geq n\}$. Then each $\partial E^{\geq n}$ is an open set. The shift map $\sigma\colon E^{\geq 1}\to\partial E$ is given by $\sigma(e_{1}e_{2}e_{3}\cdots)=e_{2}e_{3}\cdots$ for $e_{1}e_{2}e_{3}\cdots\in\partial E^{\geq 2}$ and $\sigma(e)=r(e)$ if $e\in E^{1}$. Let $\sigma^{n}\colon\partial E^{\geq n}\to\partial E$ denote the $n$-fold composition of $\sigma$, and let $\sigma^{0}=\mathrm{id}$. Then each $\sigma^{n}$ is a local homeomorphism between open subsets of $\partial E$.
As a set, the boundary path groupoid associated with $E$ is
$\mathscr{G}_{E}=\{(x,m-n,y)\in\partial E\times\mathbb{Z}\times\partial E\mid x\in\partial E^{\geq m},y\in\partial E^{\geq n},\sigma^{m}(x)=\sigma^{n}(y)\}.$
The set of composable pairs is
$\mathscr{G}_{E}^{(2)}=\{((x,m,y),(y,n,z))\mid(x,m,y),(y,n,z)\in\mathscr{G}_{E}\}.$
Multiplication and inverses are given by $(x,m,y)(y,n,z)=(x,m+n,z)$ and $(x,m,y)^{-1}=(y,-m,x)$, respectively. For $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ such that $\sigma^{|\mu|}(Z(\mu\setminus F))=\sigma^{|\nu|}(Z(\nu\setminus F))$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, let
$Z(\mu,F,\nu)=\{(x,|\mu|-|\nu|,y)\in\mathscr{G}_{E}\mid x\in Z(\mu\setminus F),y\in Z(\nu\setminus F)\}.$
Then the collection $\{Z(\mu,F,\nu)\}$, ranging over $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, forms a basis of compact open bisections for a locally compact Hausdorff topology on $\mathscr{G}_{E}$ [26, Lemma 9.2]. In particular, $\mathscr{G}_{E}$ is an ample groupoid. The unit space $\mathscr{G}_{E}^{(0)}$ is identified with $\partial E$.
All isotropy groups in $\mathscr{G}_{E}$ are (isomorphic to) subgroups of $\mathbb{Z}$ via $(x,m,x)\mapsto m$ for $x\in\mathscr{G}_{E}^{(0)}$. The isotropy group at $x\in\partial E$ is nontrivial if and only if $x$ is eventually periodic, that is, $x=e_{1}e_{2}\cdots e_{n}\mu^{\infty}$ for some $n\in\mathbb{N}\cup\{0\}$ and closed path $\mu$ of positive length. Hence, $\mathscr{G}_{E}$ is principal if and only if $E$ is acyclic.
Let $K$ be a field and let $L_{K}(E)$ denote the Leavitt path algebra of $E$ over $K$ (see [2] for the definition of $L_{K}(E)$). Then $L_{K}(E)\cong K\mathscr{G}_{E}$ with the isomorphism given by $\mu\nu^{*}\mapsto 1_{Z(\mu,\nu)}$ [9, Proposition 4.3]. In [3, Theorem 1] it is proved that $L_{K}(E)$ is regular if and only if $E$ acyclic. Using our groupoid characterization, we can give a short proof of this equivalence, and expand on it.
Theorem 5.3.
Let $E$ be a directed graph and $K$ a field. The following are equivalent.
  1. (1)
    $L_{K}(E)$ is von Neumann regular.
  2. (2)
    $E$ is acyclic.
  3. (3)
    $\mathscr{G}_{E}$ is principal.
  4. (4)
    $\mathscr{G}_{E}$ is an approximately quasi-compact groupoid.
Proof.
We already observed that acyclic is equivalent to principal. If $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular, then $\mathscr{G}_{E}$ has locally finite isotropy by Theorem 3.1 and hence $E$ must be principal as $\mathscr{G}_{E}$ has no infinite cyclic isotropy group. Suppose now that $E$ is acyclic. We show that $\mathscr{G}_{E}$ is approximately quasi-compact.
Let $F$ be a finite subgraph of $E$. Then consider the union $\mathscr{G}_{E}(F)$ of all compact open bisections $Z(\alpha,\emptyset,\beta)$ with $\alpha,\beta$ finite paths in $F$ such that $r(\alpha)=r(\beta)$. Then $\mathscr{G}_{E}(F)$ is compact open because there are only finitely many such paths $\alpha,\beta$ as $F$ is finite and acyclic. It is a subgroupoid since if $\alpha,\beta,\gamma,\delta$ are finite paths in $F$ with $r(\alpha)=r(\beta)$ and $r(\gamma)=r(\delta)$, then
$Z(\alpha,\emptyset,\beta)Z(\gamma,\emptyset,\delta)=\begin{cases}Z(\alpha\mu,\emptyset,\delta),&\text{if}\ \gamma=\beta\mu\\ Z(\alpha,\emptyset,\delta\nu),&\text{if}\ \beta=\gamma\nu\\ \emptyset,&\text{else.}\end{cases}$ (5.1)
Since $\mathscr{G}_{E}$ is the directed union of the $\mathscr{G}_{E}(F)$ as $F$ ranges over all finite subgraphs of $E$, we conclude that $\mathscr{G}_{E}$ is approximately quasi-compact.
If $\mathscr{G}_{E}$ is approximately quasi-compact (and hence approximately uniformly bounded), then all its isotropy groups must be locally finite. Since the isotropy groups of $\mathscr{G}_{E}$ are always subgroups of $\mathbb{Z}$, they must then be trivial and so $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular by Theorem 3.12. ∎
Next we show that we can recover [20, Theorem 10], which states that the Leavitt path algebra of an arbitrary graph is graded von Neumann regular. Every Leavitt path algebra $L_{K}(E)$ over a field $K$ is $\mathbb{Z}$-graded, with the grading given by homogeneous components
$L_{K}(E)_{n}=\mathrm{span}_{K}\{\alpha\beta^{*}\mid\alpha,\beta\in E^{*},|\alpha|-|\beta|=n\}.$
The corresponding grading on $K\mathscr{G}_{E}$ is induced by the continuous cocycle $c\colon\mathscr{G}_{E}\to\mathbb{Z}$ defined by $(x,m,y)=m$.
Theorem 5.4 (Hazrat).
Let $E$ be an arbitrary graph and $K$ a field. Then the Leavitt path algebra $L_{K}(E)$ is a graded von Neumann regular ring.
Proof.
To simplify notation, let $\mathscr{G}=\mathscr{G}_{E}$ be the boundary path groupoid of $E$. Then $\mathscr{G}_{0}=\{(\alpha x,0,\beta x)\in\mathscr{G}\mid|\alpha|=|\beta|\}$. In particular, the isotropy groups of $\mathscr{G}_{0}$ are trivial as $(x,m,x)\to m$ is an isomorphism of the isotropy group of $x$ in $\mathscr{G}$ with a subgroup of $\mathbb{Z}$. For a finite subgraph $F$ of $E$ and $n\geq 0$, we let
$\mathscr{G}_{0}(F,n)=\bigcup_{\alpha,\beta\in F^{*},\mathop{\boldsymbol{r}}(\alpha)=\mathop{\boldsymbol{r}}(\beta),|\alpha|=|\beta|\leq n}Z(\alpha,\emptyset,\beta).$
Then $\mathscr{G}_{0}(F,n)$ is compact open, as there are only finitely many choices of $\alpha,\beta$ given $F$ and $n$, and is a subgroupoid by (5.1). If $F\subseteq F^{\prime}$ and $n\leq n^{\prime}$, then $\mathscr{G}_{0}(F,n)\subseteq\mathscr{G}_{0}(F^{\prime},n^{\prime})$. It follows that $\mathscr{G}_{0}$ is approximately quasi-compact. We deduce now from Theorem 4.3 that $L_{K}(E)\cong K\mathscr{G}_{E}$ is graded regular. ∎
5.2. Algebraic partial skew group rings induced by topological partial actions
We refer the reader to [16] for a detailed treatment on partial actions and partial skew group rings.
Let $X$ be a locally compact Hausdorff totally disconnected space, let $G$ be a discrete group with identity element $e$, and let $K$ be a field. Let $\phi=(\{X_{g}\}_{g\in G},\{\phi_{g}\}_{g\in G})$ be a topological partial action of $G$ on $X$. In particular, each $X_{g}$ is a clopen subset of $X$, and $\phi_{g}\colon X_{g^{-1}}\to X_{g}$ is a homeomorphism with inverse $\phi_{g}^{-1}=\phi_{g^{-1}}$. For any $g\in G$, let $C_{c}(X_{g},K)$ denote the ring of all locally constant compactly supported functions $f\colon X_{g}\to K$. Since $X_{g}$ is open in $X$, we may view $C_{c}(X_{g},K)$ as an ideal in $C_{c}(X,K)$. Let $D_{g}=C_{c}(X_{g},K)$ and define $\varphi_{g}\colon D_{g^{-1}}\to D_{g}$ by $\varphi_{g}(f)(x)=f\circ\phi_{g^{-1}}(x)$. Then $\varphi_{g}$ is an isomorphism of $K$-algebras and $\varphi=(\{D_{g}\}_{g\in G},\{\varphi_{g}\}_{g\in G})$ is an algebraic partial action of $G$ on $C_{c}(X,K)$. The partial skew group ring is the $K$-algebra with underlying $K$-vector space
$C_{c}(X,K)\rtimes_{\varphi}G=\bigoplus_{g\in G}D_{g}\delta_{g}$
and multiplication defined by $(a\delta_{g})(b\delta_{h})=\varphi_{g}((\varphi_{g^{-1}}(a)b)\delta_{gh}$. The above direct sum decomposition is a $G$-grading on $C_{c}(X,K)\rtimes_{\varphi}G$ (by definition of the product) with the identity component given by $(C_{c}(X,K)\rtimes_{\varphi}G)_{e}=D_{e}\delta_{e}\cong C_{c}(X,K)$.
The transformation groupoid associated with the partial action $\phi$ is the set
$\mathscr{G}=\{(x,g,y)\in X\times G\times X\mid\,y\in X_{g^{-1}}\text{ and }x=\phi_{g}(y)\},$
with the set of composable pairs given by
$\mathscr{G}^{(2)}=\{((x,g,y),(y,h,z))\mid(x,g,y),(y,h,z)\in\mathscr{G}\}$
and multiplication and inverses defined by
$(x,g,y)(y,h,z)=(x,gh,z)\text{ and }(x,g,y)^{-1}=(y,g^{-1},x).$
With the relative topology from the product topology on $X\times G\times X$, $\mathscr{G}$ is an ample Hausdorff groupoid. The unit space $\mathscr{G}^{(0)}$ of $\mathscr{G}$ is identified with $X$ (as topological spaces). The groupoid has a natural $G$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to G$ given by $c(x,g,y)=g$. Hence, $K\mathscr{G}$ is $G$-graded. Moreover, $K\mathscr{G}$ is graded isomorphic to $C_{c}(X,K)\rtimes_{\varphi}G$. Note that $K\mathscr{G}_{e}=K\mathscr{G}^{(0)}\cong C_{c}(X,K)$.
Proposition 5.5.
Let $\phi=(\{X_{g}\}_{g\in G},\{\phi_{g}\}_{g\in G})$ be a topological partial action of a discrete group $G$ on locally compact Hausdorff and totally disconnected space $X$. Then $C_{c}(X,K)\rtimes_{\varphi}G$ is graded regular.
Proof.
Let $\mathscr{G}$ denote the transformation groupoid associated with $\phi$, and let $e\in G$ denote the identity element. By Theorem 4.2, $C_{c}(X,K)\rtimes_{\varphi}G\cong K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}\cong C_{c}(X,K)$ is regular. But $C_{c}(X,K)$ is trivially regular since if $f\in C_{c}(X,K)$ and then $ff^{\prime}f=f$ where $f^{\prime}(x)=1/f(x)$ whenever $f(x)\neq 0$, and $f^{\prime}(x)=0$ otherwise. ∎
5.3. Leavitt labeled path algebras
Let $E=(E^{0},E^{1},r,s)$ be a directed graph and let $\mathscr{A}$ be a set, called the alphabet. Let $\mathscr{A}^{*}$ denote the set of all finite words over $\mathscr{A}$ together with the empty word $\varepsilon$, and let $\mathscr{A}^{\infty}$ denote all infinite words over $\mathscr{A}$. An edge-labeling on $E$ is a surjective map $\mathscr{L}\colon E^{1}\to\mathscr{A}$. The pair $(E,\mathscr{L})$ is called a labeled graph. If $\mu=e_{1}e_{2}\cdots e_{n}$ is a finite or an infinite path in $E$, then $\mathscr{L}$ extends to a labeled path $\mathscr{L}(\mu)=\mathscr{L}(e_{1})\mathscr{L}(e_{2})\cdots\mathscr{L}(e_{n})$. The length of a labeled path $\alpha=\mathscr{L}(\mu)$ is defined as the length of the path $\mu$ in $E$.
Let $\mathscr{L}^{n}=\mathscr{L}(E^{n})$ denote the set of labeled paths of length $n$, let $\mathscr{L}^{\infty}$ denote the set of labeled paths of infinite length, and let $|\varepsilon|=0$. Set $\mathscr{L}^{\geq 1}=\cup_{n\geq 1}\mathscr{L}^{n}$, =$\mathscr{L}^{*}=\{\varepsilon\}\cup\mathscr{L}^{\geq 1}$, and $\mathscr{L}^{\leq\infty}=\mathscr{L}^{*}\cup\mathscr{L}^{\infty}$.
A labeled path $\alpha$ is a beginning of a labeled path $\beta$ if $\beta=\alpha\beta^{\prime}$ for some labeled path $\beta^{\prime}$. If $1\leq i\leq j\leq|\alpha|$, let $\alpha_{i,j}=\alpha_{i}\alpha_{i+1}\ldots\alpha_{j}$ if $j<\infty$ and $\alpha_{i,j}=\alpha_{i}\alpha_{i+1}\ldots$ if $j=\infty$. If $j<i$ set $\alpha_{i,j}=\varepsilon$. Define $\overline{\mathscr{L}^{\infty}}$ to be the set of all infinite words such that all beginnings are finite labeled paths. We write $\overline{\mathscr{L}^{\leq\infty}}=\mathscr{L}^{*}\cup\overline{\mathscr{L}^{\infty}}$.
Let $\mathscr{P}(E^{0})$ denote the power set of $E^{0}$. The relative range of $\alpha\in\mathscr{L}^{*}$ with respect to $A\in\mathscr{P}(E^{0})$ is the set
$r(A,\alpha)=\begin{cases}\{r(\mu)\ |\ \mu\in E^{\ast},\ \mathscr{L}(\mu)=\alpha,\ s(\mu)\in A\},&\text{if }\alpha\in\mathscr{L}^{\geq 1}\\ A,&\text{if }\alpha=\varepsilon.\end{cases}$
The range $r(\alpha)$ of $\alpha$ is the set $r(\alpha)=r(E^{0},\alpha)$. We define
$\mathscr{L}(AE^{1})=\{\mathscr{L}(e)\mid e\in E^{1}\ \mbox{and}\ s(e)\in A\}=\{a\in\mathscr{A}\mid r(A,a)\neq\emptyset\}.$
A labeled space is a triple $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ where $(E,\mathscr{L})$ is a labeled graph and $\mathscr{B}$ is a subset of $\mathscr{P}(E^{0})$, which is closed under finite intersections and finite unions, contains $r(\alpha)$ for every $\alpha\in\mathscr{L}^{*}$, and is closed under relative ranges, that is, $r(A,\alpha)\in\mathscr{B}$ for all $A\in\mathscr{B}$ and all $\alpha\in\mathscr{L}^{*}$. We call $\mathscr{B}$ an accommodating family for $\mathscr{E}$. A labeled space $\mathscr{E}$ is weakly left-resolving if for all $A,B\in\mathscr{B}$ and all $\alpha\in\mathscr{L}^{\geq 1}$ we have $r(A\cap B,\alpha)=r(A,\alpha)\cap r(B,\alpha)$. A weakly left-resolving labeled space is normal if $\mathscr{B}$ is closed under relative complements. A non-empty set $A\in\mathscr{B}$ is regular if for all $\emptyset\neq B\subseteq A$, we have that $0<|\mathscr{L}(BE^{1})|<\infty$. The subset of all regular elements of $\mathscr{B}$ together with the empty set is denoted by $\mathscr{B}_{reg}$. For $\alpha\in\mathscr{L}^{*}$, define
$\mathscr{B}_{\alpha}=\mathscr{B}\cap\mathscr{P}(r(\alpha))=\{A\in\mathscr{B}\mid A\subseteq r(\alpha)\}.$
If a labeled space is normal, then $\mathscr{B}_{\alpha}$ is a Boolean algebra for each $\alpha\in\mathscr{L}^{*}$.
Definition 5.6 (Definition 3.1 of [7]).
Let $K$ be a field and let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be a normal labeled space. The Leavitt labelled path algebra $L_{K}(\mathscr{E})$ associated with $\mathscr{E}$ and with coefficients in $K$ is the universal $K$-algebra with generators $\{p_{A}\mid A\in\mathscr{B}\}\cup\{s_{a},s_{a}^{*}\mid a\in\mathscr{A}\}$ subject to the relations
  1. (i)
    $p_{A\cap B}=p_{A}p_{B}$, $p_{A\cup B}=p_{A}+p_{B}-p_{A\cap B}$ and $p_{\emptyset}=0$, for every $A,B\in\mathscr{A}$;
  2. (ii)
    $p_{A}s_{a}=s_{a}p_{r(A,a)}$ and $s_{a}^{*}p_{A}=p_{r(A,a)}s_{a}^{*}$, for every $A\in\mathscr{B}$ and $a\in\mathscr{A}$;
  3. (iii)
    $s_{a}^{*}s_{a}=p_{r(a)}$ and $s_{b}^{*}s_{a}=0$ if $b\neq a$, for every $a,b\in\mathscr{A}$;
  4. (iv)
    $s_{a}s_{a}^{*}s_{a}=s_{a}$ and $s_{a}^{*}s_{a}s_{a}^{*}=s_{a}^{*}$ for every $a\in\mathscr{A}$;
  5. (v)
    For every $A\in\mathscr{B}_{reg}$,
    $p_{A}=\sum_{a\in\mathscr{L}(AE^{1})}s_{a}p_{r(A,a)}s_{a}^{*}.$
Leavitt labeled path algebras generalize Leavitt path algebras of graphs, Leavitt path algebras of ultragraphs and algebraic Exel-Laca algebras.
Next, we describe a partial action associated with a labeled space and its transformation groupoid. The Steinberg algebra of this transformation groupoid is isomorphic to the Leavitt labeled path algebra.
5.3.1. Inverse semigroup of a labeled space
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be a normal labeled space and let
$S_{\mathscr{E}}=\{(\alpha,A,\beta)\mid\alpha,\beta\in\mathscr{L}^{*}\ \mbox{and}\ A\in\mathscr{B}_{\alpha}\cap\mathscr{B}_{\beta}\ \mbox{with}\ A\neq\emptyset\}\cup\{0\}.$
Define a binary operation on $S_{\mathscr{E}}$ by letting $s\cdot 0=0\cdot s=0$ for all $s\in S_{\mathscr{E}}$ and, if $s=(\alpha,A,\beta)$ and $t=(\gamma,B,\delta)$ are in $S_{\mathscr{E}}$, then
$s\cdot t=\left\{\begin{array}[]{ll}(\alpha\gamma^{\prime},r(A,\gamma^{\prime})\cap B,\delta),&\mbox{if}\ \ \gamma=\beta\gamma^{\prime}\ \mbox{and}\ r(A,\gamma^{\prime})\cap B\neq\emptyset,\\ (\alpha,A\cap r(B,\beta^{\prime}),\delta\beta^{\prime}),&\mbox{if}\ \ \beta=\gamma\beta^{\prime}\ \mbox{and}\ A\cap r(B,\beta^{\prime})\neq\emptyset,\\ 0,&\mbox{otherwise}.\end{array}\right.$
Also, define $(\alpha,A,\beta)^{*}=(\beta,A,\alpha)$. With these operations $S_{\mathscr{E}}$ is an inverse semigroup with zero element 00; see [5, Proposition 3.4]. The semilattice of idempotents of $S_{\mathscr{E}}$ is
$E(S_{\mathscr{E}})=\{(\alpha,A,\alpha)\mid\alpha\in\mathscr{L}^{*}\ \mbox{and}\ A\in\mathscr{B}_{\alpha}\}\cup\{0\}.$
The natural order on the semilattice $E(S_{\mathscr{E}})$ is given by $p\leq q$ if only if $pq=p$. In our case, the order can described by noticing that if $p=(\alpha,A,\alpha)$ and $q=(\beta,B,\beta)$, then $p\leq q$ if and only if $\alpha=\beta\alpha^{\prime}$ and $A\subseteq r(B,\alpha^{\prime})$ [5, Proposition 4.1].
5.3.2. Filters in $E(S_{\mathscr{E}})$
The space on which we will define a partial action is the tight spectrum $\mathsf{T}$ of $E(S_{\mathscr{E}})$ (see [15, Section 12]). One may view the tight spectrum as the closure of the set of ultrafilters (maximal proper filters) in $E(S_{\mathscr{E}})$, by using the correspondence between characters on $E(S_{\mathscr{E}})$ (with the topology of pointwise convergence) and filters in $E(S_{\mathscr{E}})$. Namely, characters are precisely the characteristic functions of filters.
Filters in $E(S_{\mathscr{E}})$ are related to labeled paths in $\mathscr{E}$ in the following way. Let $\alpha\in\overline{\mathscr{L}^{\leq\infty}}$ and $\{\mathcal{F}_{n}\}_{0\leq n\leq|\alpha|}$ (understanding that ${0\leq n\leq|\alpha|}$ means $0\leq n<\infty$ when $\alpha\in\overline{\mathscr{L}^{\infty}}$) be a family such that $\mathcal{F}_{n}$ is a filter in $\mathscr{B}{\alpha_{1,n}}$ for every $n>0$, and $\mathcal{F}_{0}$ is either a filter in $\mathscr{B}$ or $\mathcal{F}_{0}=\emptyset$. The family $\{\mathcal{F}_{n}\}_{0\leq n\leq|\alpha|}$ is a complete family for $\alpha$ if
$\mathcal{F}_{n}=\{A\in\mathscr{B}{\alpha_{1,n}}\mid r(A,\alpha_{n+1})\in\mathcal{F}_{n+1}\}$
for all $n\geq 0$.
Theorem 5.7 (Theorem 4.13 of [5]).
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be a normal labeled space. Then there is a bijective correspondence between filters in $E(S_{\mathscr{E}})$ and pairs $(\alpha,\{\mathcal{F}_{n}\}_{0\leq n\leq|\alpha|})$, where $\alpha\in\overline{\mathscr{L}^{\leq\infty}}$ and $\{\mathcal{F}_{n}\}_{0\leq n\leq|\alpha|}$ is a complete family for $\alpha$.
A filter $\xi$ in $E(S_{\mathscr{E}})$ with associated labeled path $\alpha\in\overline{\mathscr{L}^{\leq\infty}}$ is sometimes denoted by $\xi^{\alpha}$ to stress the word $\alpha$; in addition, the filters in the complete family associated with $\xi^{\alpha}$ will be denoted by $\xi^{\alpha}_{n}$ (or simply $\xi_{n}$). Specifically,
$\xi^{\alpha}_{n}=\{A\in\mathscr{B}\mid(\alpha_{1,n},A,\alpha_{1,n})\in\xi^{\alpha}\}.$ (5.2)
Remark 5.8.
It follows from [5, Propositions 4.4 and 4.8] that for a filter $\xi^{\alpha}$ in $E(S_{\mathscr{E}})$ and an element $(\beta,A,\beta)\in E(S_{\mathscr{E}})$ we have that $(\beta,A,\beta)\in\xi^{\alpha}$ if and only if $\beta$ is a beginning of $\alpha$ and $A\in\xi_{|\beta|}$.
See [5, Theorems 5.10 and 6.7] for a characterization of tight filters in terms of labeled paths.
5.3.3. Partial action on the tight filters
We give a brief overview of the partial action on $\mathsf{T}$ introduced in [14].
Fix a normal labeled space $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$. We describe the topology on $\mathsf{T}$, the tight spectrum of $\mathscr{E}$, inherited from the product topology on $\{0,1\}^{E(S_{\mathscr{E}})}$, where $\{0,1\}$ has the discrete topology. For each $e\in E(S_{\mathscr{E}})$, define
$V_{e}=\{\xi\in\mathsf{T}\mid e\in\xi\},$
and for $\{e_{1},\ldots,e_{n}\}$ a finite (possibly empty) set in $E(S)$, we define
$V_{e\colon e_{1},\ldots,e_{n}}=V_{e}\cap V_{e_{1}}^{c}\cap\cdots\cap V_{e_{n}}^{c}=\{\xi\in\mathsf{T}\mid e\in\xi,e_{1}\notin\xi,\ldots,e_{n}\notin\xi\}.$
Sets of the form $V_{e\colon e_{1},\ldots,e_{n}}$ form a basis of compact open sets for a Hausdorff topology on $\mathsf{T}$ [6, Remark 4.2].
Let $\mathbb{F}$ be the free group generated by $\mathscr{A}$ (identifying the identity of $\mathbb{F}$ with $\varepsilon$). For every $t\in\mathbb{F}$ there is a clopen set $V_{t}\subseteq\mathsf{T}$, which is compact if $t\neq\varepsilon,$ and a homeomorphism $\varphi_{t}\colon V_{t^{-1}}\to V_{t}$ such that
$\varphi=(\{V_{t}\}_{t\in\mathbb{F}},\{\varphi_{t}\}_{t\in\mathbb{F}})$
is a topological partial action of $\mathbb{F}$ on $\mathsf{T}$ [14, Proposition 3.12]. In particular, $V_{\varepsilon}=\mathsf{T}$ and if $\alpha,\beta\in\mathscr{L}^{*}$, then $V_{\alpha}=V_{(\alpha,r(\alpha),\alpha)},V_{\alpha^{-1}}=V_{(\varepsilon,r(\alpha),\varepsilon)}$, and $V_{(\alpha\beta^{-1})^{-1}}=\varphi_{\beta^{-1}}^{-1}(V_{\alpha^{-1}})$, with $V_{(\alpha\beta^{-1})^{-1}}\neq\emptyset$ if and only if $r(\alpha)\cap r(\beta)\neq\emptyset$ [14, Lemma 3.10]. If $V_{t}\neq\emptyset$, then $t=\alpha\beta^{-1}$ for some $\alpha,\beta\in\mathscr{L}^{*}$ [14, Lemma 3.11(ii)], in which case
$V_{\alpha\beta^{-1}}=V_{(\alpha,r(\alpha)\cap r(\beta),\alpha)}$
by [13, Lemma 4.3]. If $\xi\in V_{\beta\alpha^{-1}}$ with associated word $\beta\gamma$, then the associated word of $\varphi_{\alpha\beta^{-1}}(\xi)$ is $\alpha\gamma$ and, for $0\leq n\leq|\gamma|$, we have that
$\varphi_{\alpha\beta^{-1}}(\xi)_{|\alpha|+n}=\{A\cap r(\alpha\gamma_{1,n})\mid A\in\uparrow_{\mathscr{B}_{\gamma_{1,n}}}\xi_{|\beta|+n}\},$
where $\uparrow_{\mathscr{B}_{\gamma_{1,n}}}\xi_{|\beta|+n}$ is the upper set of $\mathscr{B}_{\gamma_{1,n}}$ generated by $\xi_{|\beta|+n}$
$\varphi_{\beta^{-1}}(\xi)_{n}=\uparrow_{\mathscr{B}_{\gamma_{1,n}}}\xi_{|\beta|+n}.$
The transformation groupoid associated with the partial action $\Phi=(\{V_{t}\}_{t\in\mathbb{F}},\{\phi_{t}\}_{t\in\mathbb{F}})$ is
$\mathscr{G}=\{(\xi,t,\eta)\in\mathsf{T}\times\mathbb{F}\times\mathsf{T}\mid\eta\in V_{t^{-1}},\text{ and }\xi=\phi_{t}(\eta)\}$ (5.3)
with products and inverses defined by
$(\xi,s,\eta)(\eta,t,\rho)=(\xi,st,\rho)\text{ and }(\xi,t,\eta)^{-1}=(\eta,t^{-1},\xi),$
respectively. Endow $\mathscr{G}$ with the relative topology from the product topology on $\mathsf{T}\times\mathbb{F}\times\mathsf{T}$. The unit space $\mathscr{G}^{0}$ is identified with $\mathsf{T}$. By [7, Theorem 6.1], the Leavitt labeled graph algebra of $L_{K}(\mathscr{E})$ is isomorphic to $K\mathscr{G}$. The transformation groupoid $\mathscr{G}$ is isomorphic to a Deaconu-Renault groupoid associated with $\mathscr{E}$ [14, Theorem 5.5]. Thus, all isotropy subgroups of $\mathscr{G}$ are subgroups of the integers.
The Leavitt path algebra of an acyclic graph is a direct limit of Leavitt path algebras of finite acyclic graphs [1, Proposition 1.6.15], which implies the Leavitt path algebra is locally matricial [1, Proposition 2.6.20]. This fact is used to prove regularity of Leavitt path algebras of acyclic graphs. The case of labeled spaces is more intricate. For example, a row-finite graph may produce two labeled spaces, one with a regular Leavitt labeled path algebra and the other with a non-regular Leavitt labeled path algebra, depending on the labeling and accommodating family chosen (see Example 5.12). In addition, it is not known if, or under which conditions, Leavitt labeled path algebras are locally matricial.
Using our groupoid characterization we obtain some regularity results, but not an equivalence. For graphs, being acyclic is equivalent to the boundary path groupoid being principal. In the transformation groupoid of a labeled space, a tight filter $\xi^{\alpha}$ has non-trivial isotropy in $\mathscr{G}$ if and only if there exist $\beta,\gamma\in\mathscr{L}^{\geq 1}$ such that $\alpha=\beta\gamma^{\infty}$ and for all $n\in\mathbb{N}\cup\{0\}$ and all $A\in\xi_{|\beta\gamma^{n}|}$ we have that $A\cap r(A,\gamma)\neq\emptyset$ [14, Proposition 6.10]. The negation of this characterization then characterizes when the groupoid is principal. However, assuming the existence of a tight filter with associated word $\alpha=\beta\gamma^{\infty}$ such that there exists $n\in\mathbb{N}\cup\{0\}$ and $A\in\xi_{|\beta\gamma^{n}|}$ with $A\cap r(A,\gamma)=\emptyset$ seems to be an obstacle to finding an equivalence between principality and regularity. Nonetheless, we have the following results.
Theorem 5.9.
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field. If $L_{K}(\mathscr{E})$ is regular, then $\mathscr{G}_{\mathscr{E}}$ is principal.
Proof.
Suppose $L_{K}(\mathscr{E})\cong K\mathscr{G}$ is regular. Then the isotropy subgroups of $\mathscr{G}$ are locally finite by Theorem 3.1. Since all isotropy subgroups of $\mathscr{G}$ are infinite cyclic, it follows that $\mathscr{G}$ is principal. ∎
Corollary 5.10.
If $L_{K}(\mathscr{E})$ is regular, then for all $\xi^{\alpha}\in\mathsf{T}$, either $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$, or if $\alpha=\beta\gamma^{\infty}$ then there is $n\in\mathbb{N}$ and $A\in\xi_{|\beta\gamma^{n}|}$ such that $A\cap r(A,\gamma)=\emptyset$.
Proof.
The result follows from Theorem 5.9 and the characterization of nontrivial isotropy in $\mathscr{G}_{\mathscr{E}}$ [14, Proposition 6.10]. ∎
Theorem 5.11.
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field. Suppose that if $\xi^{\alpha}\in\mathsf{T}$ and $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$. Then $L_{K}(\mathscr{E})$ is regular.
Proof.
Let $F$ be a finite subgraph of $E$. Let $\mathscr{L}_{F}\colon F^{1}\to\mathscr{A}_{F}$ denote the restriction of $\mathscr{L}$ to $F^{1}\subset E^{1}$, where $\mathscr{A}_{F}=\mathscr{L}(F^{1})\subset\mathscr{A}$. Then $(F,\mathscr{L}_{F})$ is a labeled subgraph. Let $\mathscr{B}(F)=\mathscr{B}\cap\mathscr{P}(F^{0})$. Then $\mathscr{F}=(F,\mathscr{L}_{F},\mathscr{B}(F))$ is a normal labeled space, which is set-finite in the sense that $|\mathscr{L}_{F}(AF^{1})|<\infty$ for all $A\in\mathscr{B}(F)$. Notice that since we assume $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$, it follows that $|\mathscr{L}^{*}_{F}|<\infty$. And, since $F^{0}$ is finite, it follows that $|\mathscr{B}(F)|<\infty$. Hence, $|E(S_{\mathscr{F}})|<\infty$. Let
[…]
$e,e_{i},f,f_{j}\in E(S_{\mathscr{F}}),\text{ and }\alpha,\beta\in\mathscr{L}^{*}_{F}\}.$
We claim that $\mathscr{G}(F)$ is a compact open subgroupoid of $\mathscr{G}_{\mathscr{E}}$. Notice that $E(S_{\mathscr{F}})\subseteq E(S_{\mathscr{E}})$, $\mathscr{B}_{F}\subseteq\mathscr{B}$, and $\{\alpha\beta^{-1}\mid\alpha,\beta\in\mathscr{L}^{*}_{F}\}\subseteq\mathbb{F}$ (where $\mathbb{F}$ is the free group generated by $\mathscr{A}$). Thus, since each $V_{e\colon e_{1},\ldots,e_{m}}$ is compact open in $\mathsf{T}$, it follows that $\mathscr{G}(F)$ is an open subset of $\mathscr{G}_{\mathscr{E}}$. It is clear that $\mathscr{G}(F)$ is closed under inverses.
We show that $\mathscr{G}(F)$ is closed under taking products. Let $(\xi,s,\eta),(\eta,t,\rho)\in\mathscr{G}(F)$ with the product $(\xi,s,\eta)(\eta,t,\rho)$ defined in $\mathscr{G}_{\mathscr{E}}$. Then there are $p,q,m,n\in\mathbb{N}\cup\{0\}$ and $d,d_{1},\ldots,d_{p},e,e_{1},\ldots,e_{m},f,f_{1},\ldots,f_{n},g,g_{1},\ldots,g_{q}\in E(S_{\mathscr{F}})$ such that
$(\xi,s,\eta)\in V_{d\colon d_{1},\ldots,d_{p}}\times\{s\}\times\times V_{e\colon e_{1},\ldots,e_{m}},\text{ and }$
$(\eta,t,\rho)\in V_{f\colon f_{1},\ldots,f_{n}}\times\{t\}\times\times V_{g\colon g_{1},\ldots,g_{q}}.$
Suppose $e=(\alpha,A,\alpha)$ and $f=(\beta,B,\beta)$ with $\alpha,\beta\in\mathscr{L}_{F}^{*}\subset\mathscr{L}^{*}$ and $A,B\in\mathscr{B}_{F}$. Hence, $\eta=\eta^{\alpha\gamma}=\eta^{\beta\delta}$, for some $\alpha\gamma,\beta\delta\in\overline{\mathscr{L}^{\leq\infty}}$ with $A\in\eta_{|\alpha|}$ and $B\in\eta_{|\beta|}$. The one-to-one correspondence of Theorem 5.7 between tight filters in $\mathsf{T}$ and pairs consisting of a labeled path and a complete family of filters implies that $\alpha\gamma=\beta\delta$. Hence, $\alpha=\beta\alpha^{\prime}$ or $\beta=\alpha\beta^{\prime}$.
Suppose first that $\alpha=\beta\alpha^{\prime}$. Then $\beta\delta=\beta\alpha^{\prime}\gamma$. By [14, Lemma 5.3], there are $\tau,\nu\in\mathscr{L}_{F}^{*}$ such that
$(\xi,s,\eta)=(\xi^{\tau\gamma},\tau\alpha^{-1},\eta^{\alpha\gamma})=(\xi^{\tau\gamma},\tau(\beta\alpha^{\prime})^{-1},\eta^{\beta\alpha^{\prime}\gamma})$
and
$(\eta,t,\rho)=(\eta^{\beta\delta},\beta\nu^{-1},\rho^{\nu\delta})=(\eta^{\beta\alpha^{\prime}\gamma},\beta\alpha^{\prime}(\nu\alpha^{\prime})^{-1},\rho^{\nu\alpha^{\prime}\gamma}).$
Thus,
$(\xi,s,\eta)(\eta,t,\rho)$ $=$ $(\xi^{\tau\gamma},\tau(\nu\alpha^{\prime})^{-1},\rho^{\nu\alpha^{\prime}\gamma})$ (5.4)
$\in$ $V_{d\colon d_{1},\ldots,d_{p}}\times\{\tau(\nu\alpha^{\prime})^{-1}\}\times V_{g\colon g_{1},\ldots,g_{q}}.$
On the other hand, if $\beta=\alpha\beta^{\prime}$, then $\alpha\gamma=\alpha\beta^{\prime}\delta$. Similarly to above, we see that
$(\xi,s,\eta)(\eta,t,\rho)=(\xi^{\tau\beta^{\prime}\delta},\tau\beta^{\prime}\nu^{-1},\rho^{\nu\delta})\in V_{d\colon d_{1},\ldots,d_{p}}\times\{\tau\beta^{\prime}\nu^{-1}\}\times V_{g\colon g_{1},\ldots,g_{q}}.$
Hence $\mathscr{G}(F)$ is closed under multiplication from $\mathscr{G}_{\mathscr{E}}$, and thus an open subgroupoid of $\mathscr{G}_{\mathscr{E}}$.
Since $|\mathscr{L}^{*}_{F}|<\infty$, $|E(S_{\mathscr{F}})|<\infty$ and $V_{e\colon e_{1},\ldots,e_{m}}$ is compact for all $e,e_{i}\in E(S_{\mathscr{F}})$, it follows that $\mathscr{G}(F)$ is a finite union of compact sets, and therefore itself compact. Then, $\mathscr{G}_{\mathscr{E}}$ is a directed union of $\mathscr{G}(F)$, ranging over all finite subgraphs $F$ of $E$, and we conclude that $\mathscr{G}_{\mathscr{E}}$ is approximately quasi-compact.
Notice that the assumption $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$ implies that all isotropy in $\mathscr{G}_{\mathscr{E}}$ are trivial [14, Proposition 6.10]. Thus, the order of no isotropy element is divisible by the characteristic of $K$. Hence, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is regular by Theorem 3.12. ∎
Example 5.12.
We give an example that illustrates the significant role that the labeling and the accommodating family plays in the algebraic structure of algebras associated with labeled spaces. Let $E$ be the graph with vertices $E^{0}=\{v_{i}\mid i\in\mathbb{Z}\}$ and edges $E^{1}=\{e_{i}\mid i\in\mathbb{Z}\}$ such that $r(e_{i})=s(e_{i+1})$ for all $i\in\mathbb{Z}$. $E$ is an acyclic row-finite graph. Let $\mathscr{A}=\{a\}$ and let $\mathscr{L}(e_{i})=a$ for all $i\in\mathbb{Z}$. Let $\mathscr{B}$ be the Boolean algebra generated by all finite and cofinite subsets of $E^{0}$. Then $\mathscr{B}$ is an accommodating family for the labeled graph $(E,\mathscr{L})$ and $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ is a normal labeled space. Here, =$\mathscr{L}^{*}=\{a^{n}\mid n\in\mathbb{N}\}\cup\{\epsilon\}$ and, therefore, $a^{\infty}\in\overline{\mathscr{L}^{\leq\infty}}$. This implies that $\mathscr{G}_{\mathscr{E}}$ is not principal and, therefore, $L_{K}(\mathscr{E})$ is not regular by Theorem 5.9. The underlying graph, however, is acyclic and thus its Leavitt path algebra $L_{K}(E)$ is regular by Theorem 5.3.
Now let $\mathscr{A}^{\prime}=E^{1}$, let $\mathscr{C}$ be the set of all finite subsets of $E^{0}$ and let $\mathscr{K}\colon E^{1}\to\mathscr{A}^{\prime}$ be the identity map. Then, $\mathscr{E}^{\prime}=(E,\mathscr{K},\mathscr{C})$ is a normal label space. In this instance, $L_{K}(\mathscr{E}^{\prime})\cong L_{K}(E)$ (by [7, Example 7.1]) and is, therefore, regular.
Remark 5.13.
Theorem 5.11 gives a partial converse to Corollary 5.10 and, thus, to Theorem 5.9. However, we don’t see how to show that the property “if $\alpha\neq\beta\gamma^{\infty}$, then there exists $n\in\mathbb{N}\cup\{0\}$ and $A\in\xi_{|\beta\gamma^{n}|}$ such that $A\cap r(A,\gamma)=\emptyset$” implies regularity. In fact, we do not even have an example of a labeled space that exhibits this property, nor do we know whether it influences regularity of $L_{K}(\mathscr{E})$.
Every Leavitt labeled path algebra $L_{R}(\mathscr{E})$ is $\mathbb{Z}$-graded, with homogeneous components given by
$L_{R}(\mathscr{E})_{n}=\mathrm{span}_{K}\{s_{\alpha}p_{A}s_{\beta}^{*}\mid\alpha,\beta\in\mathscr{L}^{*},\ A\in\mathscr{B}_{\alpha}\cap\mathscr{B}_{\beta}\ \mbox{and}\ |\alpha|-|\beta|=n\}$
for $n\in\mathbb{Z}$ [7, Proposition 3.8].
Theorem 5.14.
Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field. Then the Leavitt labeled path algebra $L_{K}(\mathscr{E})$ is graded regular.
Proof.
To ease notation, let $\mathscr{G}=\mathscr{G}_{\mathscr{E}}$ denote the transformation groupoid of the partial action associated with $\mathscr{E}$. The map $c\colon(\xi^{\alpha\gamma},\alpha\beta^{-1},\eta^{\beta\gamma})\mapsto|\alpha|-|\beta|$ is a continuous cocycle $c\colon\mathscr{G}\to\mathbb{Z}$ that induces the $\mathbb{Z}$-grading on $L_{R}(\mathscr{E})$ via the isomorphism $L_{K}(\mathscr{E})\cong K\mathscr{G}$. Then
$\mathscr{G}_{0}=\{(\eta^{\alpha\gamma},\alpha\beta^{-1},\eta^{\beta\gamma})\mid\alpha,\beta\in\mathscr{L}^{*},\,r(\alpha)\cap r(\beta)\neq\emptyset,|\alpha|=|\beta|\}.$
If $(\eta,t,\eta)\in\mathscr{G}_{0}$ is an isotropy group element, then $\eta=\eta^{\alpha\gamma^{\infty}}$, for some $\alpha\in\mathscr{L}^{*}$ and $\gamma\in\mathscr{L}^{\geq 1}$, and either $t=\alpha\gamma\alpha^{-1}$ or $t=\alpha\gamma^{-1}\alpha^{-1}$ [7, Remark 8.2]. Without loss of generality, suppose $t=\alpha\gamma\alpha^{-1}$. Then, $c(\eta,t,\eta)=|\gamma|=0$, which implies that $t=\varepsilon$. Hence $\mathscr{G}_{0}$ is has trivial isotropy groups.
Let $F$ be a finite subgraph of $E$, let $\mathscr{L}_{F}\colon F^{1}\to\mathscr{A}$ be the restriction of $\mathscr{L}$ to $F^{1}\subset E^{1}$, and let $\mathscr{B}(F)=\mathscr{B}\cap\mathscr{P}(F^{0})$. Then $\mathscr{F}=(F,\mathscr{L}_{F},\mathscr{B}(F))$ is a normal labeled space. For $n\geq 0$, let
$\mathscr{G}_{0}(F,n)=\bigcup_{\begin{subarray}{c}\alpha,\beta\in\mathscr{L}_{F}^{*}\\ A\in\mathscr{B}(F)\\ A\cap r(\alpha)\cap r(\beta)\neq\emptyset\\ |\alpha|=|\beta|\leq n\end{subarray}}V_{(\alpha,A,\alpha)}\times\{\alpha\beta^{-1}\}\times V_{(\beta,A,\beta)}$
Then $\mathscr{G}_{0}(F,n)$ is an open subgroupoid of $\mathscr{G}_{0}$ by (5.4). Moreover, given $n$ and that $F$ is finite, it follows that there are only finitely many choices of $A$, $\alpha$, $\beta$ and $B$. Thus, $\mathscr{G}_{0}(F,n)$ is compact, as it is a finite union of compact sets.
If $F\subseteq F^{\prime}$ and $n\leq n^{\prime}$, then $\mathscr{G}_{0}(F,n)\subseteq\mathscr{G}_{0}(F^{\prime},n^{\prime})$. Thus, $\mathscr{G}_{0}$ is approximately quasi-compact. By Theorem 4.3, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is graded regular. ∎
5.4. Deaconu-Renault groupoids
A monoid $P$ is called left reversible (or right Ore) if $aP\cap bP\neq\emptyset$ for all $a,b\in P$. If $P$ is cancellative and left reversible, then it embeds in its group of fractions $G$. One can construct $G$ as $(P\times P)/{\sim}$ where $(p,q)\sim(p^{\prime},q^{\prime})$ if the exist $x,y\in P$ with $(px,qx)=(p^{\prime}y,q^{\prime}y)$. One writes $pq^{-1}$ for the class of $(p,q)$ and the product is given by $pq^{-1}rs^{-1}=px(sy)^{-1}$ where $x,y\in P$ with $qx=ry$. The embedding of $P$ into $G$ is $p\mapsto pe^{-1}$, which we write as $p$ from now on. Note that $G=PP^{-1}$.
Following [30], we consider right actions of $P$ on a space $X$ by partial local homeomorphisms. So we have a partially defined mapping $X\times P\to X$ (written $(x,p)\mapsto xp$) such that if $U_{p}$ is the set of $x\in X$ with $xp$ defined and $V_{p}=\{xp\mid x\in U_{p}\}$, then $U_{p}$ and $V_{p}$ are open, and the following hold:
  1. (1)
    $x\mapsto xp$ is a local homeomorphism $U_{p}\to V_{p}$ for all $p\in P$.
  2. (2)
    $U_{e}=X$ and $xe=x$ for all $x\in X$.
  3. (3)
    $x\in U_{pq}$ if and only if $x\in U_{p}$ and $xp\in U_{q}$, in which case $(xp)q=x(pq)$.
The action is said to be directed if, for all $p,q\in P$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Notice that left reversibility is equivalent to this condition if $U_{p}=X$ for all $p\in P$. More generally, if $F\subseteq P$, then $F$ is said to be action directed if given $p,q\in F$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in F\cap pP\cap qP$ with $U_{p}\cap U_{q}=U_{r}$. It is shown in [30, Claim 3] that every finite subset of $P$ is contained in an action-directed finite subset of $P$.
From now on we assume that $X$ is locally compact, Hausdorff and totally disconnected.
For example, if $E$ is a directed graph, then there is an action of $\mathbb{N}$ on $\partial E$ by local homeomorphisms by putting $xn=\sigma^{n}(x)$ where $\sigma$ is the shift map. Here $U_{n}=\partial E^{\geq n}$. The action-directed condition is trivial since $r$ can be taken to be the maximum of $p,q$.
The Deaconu-Renault groupoid associated to the action of $P$ on $X$ is
$\mathscr{G}(X,P)=\{(x,g,y)\in X\times G\times Y\mid\exists p,q\in P,xp=yq,g=pq^{-1}\}.$
A pair $(x,g,y),(x^{\prime},g^{\prime},y^{\prime})$ is composable if and only if $y=x^{\prime}$, in which case $(x,g,y)(x^{\prime},g^{\prime},y^{\prime})=(x,gg^{\prime},y^{\prime})$. The inversion is given by $(x,g,y)^{-1}=(y,g^{-1},x)$. The unit space is $\mathscr{G}(X,P)^{(0)}=\{(x,e,x)\mid x\in X\}$, which we can identify with $X$.
A basis for the topology on $\mathscr{G}(X,P)$ consists of all sets of the form $D(U,p,q,V)=\{(x,pq^{-1},y)\mid x\in U,y\in V,xp=yq\}$ with $U$ and $V$ open subsets of $X$. Such a set is a compact open bisection if $U\subseteq U_{p}$, $V\subseteq V_{q}$ are compact open, $p$ is injective on $U$, $q$ is injective on $V$ and $Up=Vq$. These compact open bisections form a basis for the topology, and so the groupoid is ample. The topology on $\mathscr{G}(X,P)$ is finer than the product topology on $X\times G\times X$ but coincides with the product topology on sets of the form $D(U,p,q,V)$ (see [30, Lemma 5.11]). The groupoid $\mathscr{G}(X,P)$ is Hausdorff. There is a continuous cocycle $c\colon\mathscr{G}(X,P)\to G$ given by $c(x,g,y)=g$ (see [30, Proposition 5.12]).
Notice that the boundary path groupoid $\mathscr{G}_{E}$ for a graph $E$ is a Deaconu-Renault groupoid for the action of $\mathbb{N}$ via the shift map. Similarly, the boundary path groupoid of a higher rank graph (or $k$-graph) is a Deanconu-Renault groupoid for an action of $\mathbb{N}^{k}$ on the boundary path space [30].
Our goal is to show that $\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle provided that each set $U_{g}$ is clopen. This occurs in particular if each $U_{g}=X$. It also happens in the case of row-finite graphs and row-finite higher rank graphs. In this setting, the finite paths in the boundary path space are isolated points. Since every infinite path is in the domain of any shift map, it follows that the domains are clopen in the row-finite case. The following result recovers graded regularity for Leavitt path algebras of row-finite graphs and also establishes graded regularity for row-finite higher ranks graphs.
Theorem 5.15.
Let $K$ be a field and let $\mathscr{G}(X,P)$ be the Deaconu-Renault groupoid of the directed action of a cancellative left reversible monoid $P$ on a locally compact, Hausdorff and totally disconnected space $X$ by partial local homeomorphisms. Assume that the domain $U_{p}$ of the action of $p$ is clopen for all $p\in P$. Then $K\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle, where $G$ is the group of fractions of $P$.
Proof.
Note that $c^{-1}(e)=\{(x,e,y)\in\mathscr{G}(X,P)\}$ and hence is a principal groupoid. So it suffices by Theorem 4.3 to show that $c^{-1}(e)$ is approximately quasi-compact. Let $W$ be a compact open subset of $X$, and let $F$ be an action-directed finite subset of $P$. Put $\mathscr{G}(W,F)=\{(x,e,y)\mid x,y\in W,\exists p\in F,xp=yp\}$. We claim that $\mathscr{G}(W,F)$ is a compact open subgroupoid of $c^{-1}(e)$. First of all,
$\mathscr{G}(W,F)=\bigcup_{p\in F}D(W\cap U_{p},p,p,W\cap U_{p})$
and hence is compact open, since $W\cap U_{p}$ is compact open because $U_{p}$ is clopen, and the topology on $D(W\cap U_{p},p,p,W\cap U_{p})$ is the product topology on $(W\cap U_{p})\times\{e\}\times(W\cap U_{p})$ by [30, Lemma 5.11]. Indeed, since $X$ is Hausdorff, the set of elements of $U_{p}\times U_{p}$ with $xp=yp$ is closed in $U_{p}\times U_{p}$ and hence a compact subset of $(W\cap U_{p})\times(W\cap U_{p})$ as $U_{p}$ is clopen. It remains to check that $\mathscr{G}(W,F)$ is a subgroupoid. It is clearly closed under inversion. But if $(x,e,y),(y,e,z)\in\mathscr{G}(W,F)$, then we can find $p,q\in F$ with $xp=yp$ and $yq=zq$. Since $F$ is action directed, we can find $r\in F\cap pP\cap qP$ with $U_{r}=U_{p}\cap U_{q}$. Then $y\in U_{r}$. Let $r=pa=qb$. Then $yr$ defined implies that $a$ is defined on $yp=xp$ and $b$ is defined on $yq=zq$. Therefore, $xr=xpa$ and $zr=zqb$ are defined and $xr=xpa=ypa=yr=yqb=zqb=zr$. Thus $(x,e,z)\in\mathscr{G}(W,F)$. We conclude that $\mathscr{G}(W,F)$ is a compact open subgroupoid. If $F,F^{\prime}$ are action-directed finite subsets of $P$ and $W,W^{\prime}$ are compact open subsets of $X$, then there is an action-directed finite subset $F^{\prime\prime}$ with $F\cup F^{\prime}\subseteq F^{\prime\prime}$ by [30, Claim 3]. Then $\mathscr{G}(W,F),\mathscr{G}(W^{\prime},F^{\prime})\subseteq\mathscr{G}(W\cup W^{\prime},F^{\prime\prime})$ by construction. If $(x,e,y)\in c^{-1}(e)$, then we can find a compact open set $W$ containing $x,y$ and we can find $p\in P$ with $xp=yp$. Clearly, $\{p\}$ is action directed. It follows that $c^{-1}(e)$ is the directed union of the compact open subgroupoids $\mathscr{G}(W,F)$ and hence is approximately quasi-compact. The result follows. ∎
5.5. Self-similar group actions on graphs and Nekrashevych-Exel-Pardo algebras
Nekrashevych introduced algebras associated to self-similar groups [24, 25] and these were generalized by Exel-Pardo [17] to self-similar actions on graphs. To keep things simple, we will work with finite directed graphs with no sinks.
Let $E$ be a finite directed graph with no sinks. Then $\partial E=E^{\infty}$ is compact Hausdorff and totally disconnected. If $G$ is a discrete group, then a self-similar action of $G$ on $E$ consists of an action of $G$ on $E$ by digraph automorphisms and a mapping $G\times E^{1}\to G$, written $(g,e)\mapsto g|_{e}$ and called the section of $g$ at $e$, satisfying:
  1. (1)
    $g(v)=g|_{e}(v)$ for all $v\in E^{0}$ and $e\in E^{1}$;
  2. (2)
    $(gh)|_{e}=g|_{h(e)}h|_{e}$.
One easily checks that $G$ acts on $E^{*}$ via the recursive formula $g(ex)=g(e)g|_{e}(x)$ for $ex\in E^{*}$ and with $e\in E^{1}$. This product makes sense because of the first axiom and it is an action by the second. It follows that the action of $G$ on $E^{*}$ induces an action of $G$ on $\partial E=E^{\infty}$. The action of $g\in G$ on an infinite path can be computed recursively via the formula $g(ex)=g(e)g|_{e}(x)$ for $ex\in E^{\infty}$ and $e\in E^{1}$. Note that this action preserves the length of the longest common prefix of two infinite paths and hence $G$ acts on $\partial E$ by homeomorphisms. We say that the self-similar group action is faithful if the action of $G$ on $\partial E$ is faithful.
The classical notion of self-similar groups [23] concerns the special case when $E$ has a single vertex and at least two edges. The following notion is standard for self-similar group actions for single-vertex graphs and we define it now in the general case. The mapping $G\times E^{1}\to G$ given by $(g,e)\mapsto g|_{e}$ extends to an action of the free monoid on $E^{1}$ on the set $G$ by mappings. We say that $g\in G$ is finite state if the forward orbit of $g$ under this action is finite. We say that $G$ is finite state if each element of $G$ is finite state. It is easy to check (and well known for the case when $E$ is a single-vertex graph [23]) that if $X\subseteq G$ is any subset closed under taking sections (i.e., under the action of the free monoid), then the subgroup $H=\langle X\rangle$ acting on $E$ is self-similar, that is, $h|_{e}\in H$ for all $h\in H$ and $e\in E^{1}$. Moreover, if $|X|<\infty$, then each element of $H$ is finite state. Finitely generated, finite state self-similar groups are also known as automaton groups.
There is an algebra associated to any self-similar group action, generalizing the Leavitt path algebra [24, 17]. It can be given by generators and relations but we give the groupoid description. Exel and Pardo [17] construct an inverse semigroup $S_{G,E}$ and then use a groupoid of germs construction to build the groupoid. The inverse semigroup $S_{G_{,}E}$ can be defined as the inverse semigroup with zero whose nonzero elements consists of expressions of the form $\alpha g\beta^{*}$ with $\alpha,\beta\in E^{*}$ and $r(\alpha)=gr(\beta)$. The product is given by
$\alpha g\beta^{*}\cdot\lambda h\nu^{*}=\begin{cases}\alpha g(\gamma)g|_{\gamma}\nu^{*},&\text{if}\ \lambda=\beta\gamma,\\ \alpha(h^{-1}|_{\gamma})^{-1}(\nu h^{-1}(\gamma))^{*},&\text{if}\ \beta=\lambda\gamma,\\ 0,&\text{else.}\end{cases}$
The involution is given by $(pgq^{*})^{*}=qg^{-1}p^{*}$. If $E$ has a single vertex $v$, we can identify the group of units of $S_{G,E}$ with $G$, but otherwise $S_{G,E}$ is not monoid.
There is an action of $\theta\colon S_{G,E)}\to I_{\partial E}$ defined as follows. The zero acts as the empty map. If $\alpha g\beta^{*}\in S_{G,E}$ with $gr(\beta)=r(\alpha)$, the corresponding partial homeomorphism has domain $Z(\beta)$, range $Z(\alpha)$ and $\theta_{\alpha g\beta^{*}}(\beta\eta)=\alpha g(\eta)$ for $\eta\in Z(r(\beta))$.
The groupoid $\mathscr{G}_{G,E}$ is the groupoid of germs of the action of $S_{G,E}$ on $\partial E$. If $G$ is trivial, this recovers the boundary path groupoid $\mathscr{G}_{E}$. The groupoid $\mathscr{G}_{G,E}$ is not usually Hausdorff. The Exel-Pardo algebra of $(G,E)$ over $K$ is $K\mathscr{G}_{G,E}$. When $G$ is trivial, this is just $L_{K}(E)$.
There is a well-defined continuous cocycle $c\colon\mathscr{G}_{G,E}\to\mathbb{Z}$ given by putting $c([\alpha g\beta^{-1},x])=|\alpha|-|\beta|$ and this induces a $\mathbb{Z}$-grading on $K\mathscr{G}_{G,E}$. The following is an analogue of Theorem 5.4 for Exel-Pardo algebras.
Theorem 5.16.
Let $(G,E)$ be a faithful finite state self-similar group action of a group $G$ on a finite directed graph $E$ with no sinks and let $K$ be a field of characteristic 00. Then the Exel-Pardo algebra $K\mathscr{G}_{G,E}$ is graded regular with respect to its usual $\mathbb{Z}$-grading if and only if $G$ is locally finite.
Proof.
Suppose first that $K\mathscr{G}_{G,E}$ is graded regular. There is a homomorphism $\psi\colon G\to\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ given by $\psi(g)=\{[g(v)gv^{*},\nu]\mid v\in Z(v),v\in E^{0}\}$. Moreover, $\psi$ is injective because the action is faithful, and so if $g\neq 1$, then $g(x)\neq x$ for some $x\in\partial E$, whence $\mathop{\boldsymbol{r}}([g(s(x))gs(x)^{*},x])=gx\neq x=\mathop{\boldsymbol{s}}([g(s(x))gs(x)^{*},x])$. But $\psi(1)=G^{(0)}$. So $\psi(g)\neq\psi(1)$. By Theorem 3.7, $\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ is locally finite and hence $G\cong\psi(G)$ is locally finite.
For the converse, let $(S_{G,E})_{0}$ be the inverse subsemigroup of $S_{G,E}$ consisting of 00 and all elements of the form $\alpha g\beta^{*}$ with $r(\alpha)=r(\beta)$ and $|\alpha|=|\beta|$. Then $(S_{G,E})_{0}$ contains all the idempotents of $S_{G,E}$, from which it follows that $(\mathscr{G}_{G,E})_{0}$ is the groupoid of germs of the action of $(S_{G,E})_{0}$ on $\partial E$. Therefore, by Corollary 3.5 and Theorem 4.3 it suffices to show that $(S_{G,E})_{0}$ is locally finite. Since $G$ is finite state, if $X\subseteq G$ is finite then replacing $X$ by its forward orbit under the action of the free monoid on $E$, we see that $G$ is the direct limit of subgroups generated by finite subsets $Y$ of $G$ closed under sections. Since $G$ is locally finite, such a finite subset $Y$ generates a finite self-similar subgroup of $G$. Thus we are reduced to showing that $(S_{G,E})_{0}$ is locally finite when $G$ is finite. If $F$ is any finite subgraph of $E$, then since $G$ is finite, $G\cdot F$ is a finite $G$-invariant subgraph of $E$ containing $F$. Thus $(S_{G,E})_{0}$ is the direct limit of the $(S_{G,F})_{0}$ with $F$ running over finite $G$-invariant subgraphs of $E$. This reduces us to the case that $E$ is finite. But if $G$ and $E$ are finite, then we can write $(S_{G,E})_{0}$ as the direct limit of the finite inverse subsemigroups $S_{n}$ with $n\geq 0$, where $S_{n}$ consists of 00 and all elements $\alpha g\beta^{*}$ with $r(\alpha)=r(\beta)$, $|\alpha|=|\beta|\leq n$ and $g\in G$. Thus $(S_{G,E})_{0}$ is locally finite when $G$ and $E$ are finite. This completes the proof. ∎
Most famous examples of self-similar groups are finite state (like the Grigorchuk group) but few of the commonly studied ones are locally finite. Of course, when $G$ is trivial we recover Hazrat’s theorem on graded regularity of Leavitt path algebras in characteristic 00 [20] for the special case of finite graphs with no sinks.
Acknowledgments
We thank Gilles de Castro for showing Example 5.12 to us.

### v3

@@
 Abstract.
-We study the question of when the algebra of an ample groupoid over a field is von Neumann regular and, more generally, when the algebra of a graded ample groupoid is graded von Neumann regular.
-For fields of characteristic 00 we are able to provide a complete characterization.
-Over fields of positive characteristic, we merely have some sufficient and some necessary conditions for regularity.
-As applications we recover known results on regularity and graded regularity of Leavitt path algebras and inverse semigroups.
-We also prove a number of new results, in particular concerning graded regularity of algebras of Deaconu-Renault groupoids and Nekrashevych-Exel-Pardo algebras of self-similar groups.
+We completely characterize when the algebra of an ample groupoid with coefficients in an arbitrary unital ring is von Neumann regular and, more generally, when the algebra of a graded ample groupoid is graded von Neumann regular.
+Our main application is to resolve the question, open since 1970, of when the algebra of an inverse semigroup is von Neumann regular.
+As applications we recover known results on regularity and graded regularity of Leavitt path algebras, and prove a number of new results, in particular concerning graded regularity of algebras of Deaconu-Renault groupoids and Nekrashevych-Exel-Pardo algebras of self-similar groups.
 Key words and phrases:
@@
 It was later observed that regular rings are precisely the rings for which every module is flat, and so from the point of view of homological algebra they are a natural generalization of semisimple rings.
-A famous result of Connell [12] says that a group ring $KG$, over a field $K$, is regular if and only if $G$ is locally finite and the characteristic of $K$ does not divide the order of any element of $G$.
+A famous result of Connell [12] says that a group ring $RG$, over a unital ring $R$, is regular if and only if $R$ is regular, $G$ is locally finite and the order of every finite subgroup of $G$ is invertible in $R$.
 In particular, this says that regularity is a rare feature for groups rings.
 Leavitt path algebras [1] are also rarely regular: they are regular if and only if the (directed) graph defining the Leavitt path algebra is acyclic [3].
-However, every Leavitt path algebra is graded von Neumann regular by a result of Hazrat [20].
+However, every Leavitt path algebra over a field is graded von Neumann regular by a result of Hazrat [20].
 A graded ring $R$ is graded regular if the semigroup of homogeneous elements is regular.
-This has an important impact on the graded $K$-theory of Leavitt path algebras.
+This has an important impact on the graded $K$-theory of Leavitt path algebras [21].
 Okniński tried to generalize Connell’s results to inverse semigroup algebras (see his book [29]).
@@
 In positive characteristic he was able to give necessary conditions: $S$ is periodic, each maximal subgroup of $S$ is locally finite and the characteristic of $K$ cannot divide the order of any element of a maximal subgroup.
-It had been proven earlier that if $S$ is locally finite and the characteristic of $K$ divides the order of no element of any maximal subgroup, then $KS$ is regular.
+It had been proven earlier by Weissglass [34] that if $R$ is regular, $S$ is locally finite and the order of every finite subgroup of $S$ is invertible in $R$, then $RS$ is regular.
 But being periodic with locally finite maximal subgroups is a strictly weaker condition than being locally finite.
@@
 The proof is a reduction to Connell’s result.
-Our main result is a complete characterization of regularity of $K\mathscr{G}$ when $K$ is a field of characteristic 00.
+Our main result is a complete characterization of regularity of $R\mathscr{G}$ over any base unital ring $R$.
 Following the Bourbaki tradition, we say a space is quasi-compact if every open covering has a finite subcover and reserve “compact” to mean quasi-compact and Hausdorff.
-Then $K\mathscr{G}$ is regular if and only if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids.
-For principal, second countable Hausdorff groupoids this is equivalent to approximate finiteness in the sense of [18], hence we call such groupoids approximately quasi-compact.
-Like Okniński, we use analytic methods, but this time we embed $\mathbb{C}\mathscr{G}$ into the reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$.
-In positive characteristic, we are able to show that if $\mathscr{G}$ is approximately quasi-compact and the characteristic of $K$ does not divide the order of any isotropy element of $\mathscr{G}$, then $K\mathscr{G}$ is regular.
-But we are unable to prove in general the necessity of being approximately quasi-compact in positive characteristic.
+Theorem A.
+Let $\mathscr{G}$ be an ample groupoid and $R$ a unital ring. Then $R\mathscr{G}$ is von Neumann regular if and only if:
+  1. (1)
+    $R$ is von Neumann regular;
+  2. (2)
+    $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids;
+  3. (3)
+    the order of each finite subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
+For principal, second countable Hausdorff groupoids the second condition is equivalent to approximate finiteness in the sense of [18], hence we call such groupoids approximately quasi-compact.
+As a consequence, we resolve the problem of characterizing regular inverse semigroup algebras, dating back to the 1970 paper of Weissglass [34].
+Theorem B.
+Let $S$ be an inverse semigroup and $R$ a unital ring. Then $RS$ is von Neumann regular if and only if:
+  1. (1)
+    $R$ is von Neumann regular;
+  2. (2)
+    $S$ is locally finite;
+  3. (3)
+    the order of each finite subgroup of $S$ is invertible in $R$.
+The sufficiency of these conditions goes back to Weissglass [34].
 We also consider graded regularity of groupoid algebras for ample groupoids equipped with a locally constant cocycle to a group.
 We prove the algebra is graded regular if and only if the homogeneous component of the identity is regular.
-Since the homogeneous component of the identity is the algebra of the groupoid of elements mapping to the identity under the cocycle, our results for regularity of groupoid algebras apply.
+Since the homogeneous component of the identity is the algebra of the clopen subgroupoid of elements mapping to the identity under the cocycle, our results for regularity of groupoid algebras apply.
 From this we easily recover Hazrat’s result on graded regularity of Leavitt path algebras [20].
@@
 The final section, applies the main results to study regularity and graded regularity for the algebras of a number of well-studied families of ample groupoids.
+We added an appendix giving a groupoid-free proof of Theorem B for specialists in semigroup theory, although the proof is merely a retranslation of the groupoid proof.
+Acknowledgments
+We thank Gilles de Castro for showing Example 5.11 to us.
 2. Preliminaries
@@
 The reader is referred to [23] for details on the theory of inverse semigroups.
-A semigroup is a set together with an associative binary operation.
+A semigroup is a nonempty set together with an associative binary operation.
 An inverse semigroup is a semigroup $S$ such that for every $s\in S$ there is a unique $s^{*}\in S$ such that $s^{*}ss^{*}=s^{*}$ and $ss^{*}s=s$.
 Notice that $ss^{*},s^{*}s$ are idempotents for all $s\in S$.
-The set $E(S)$ of idempotents of $S$ is a commutative subsemigroup and hence a meet semilattice, with the partial order given by $e\leq f$ if and only if $ef=e$, and the meet by $e\wedge f=ef$.
+The set $E(S)$ of idempotents of $S$ is a commutative subsemigroup and hence a meet semilattice with the partial order given by $e\leq f$ if and only if $ef=e$, and the meet by $e\wedge f=ef$.
 The order extends to $S$ by defining $s\leq t$ if $s=te$ for some $e\in E(S)$.
@@
 A topological groupoid is a groupoid $\mathscr{G}$ endowed with a locally compact topology such that composition and inversion are continuous.
+All groupoids are assumed nonempty.
 We do not assume that $\mathscr{G}$ is Hausdorff in general, but we do assume that the unit space $\mathscr{G}^{(0)}=\{\gamma\gamma^{-1}\mid\gamma\in\mathscr{G}\}$ is Hausdorff with the relative topology inherited from $\mathscr{G}$.
@@
 Note that a finite union of quasi-compact sets is quasi-compact and a closed subset of a quasi-compact set is quasi-compact.
-An étale groupoid is a topological groupoid $\mathscr{G}$ such that the range map $\mathop{\boldsymbol{r}}\colon\mathscr{G}\to\mathscr{G}^{(0)}$ (or equivalently, the source map $\mathop{\boldsymbol{s}}$) is a local homeomorphism.
-A bisection of an étale groupoid $\mathscr{G}$ is an open subset $U\subseteq\mathscr{G}$ such that the restrictions $\mathop{\boldsymbol{r}}|_{U}$ and $\mathop{\boldsymbol{s}}|_{U}$ are homeomorphisms onto an open set in $\mathscr{G}^{(0)}$.
+An étale groupoid is a topological groupoid $\mathscr{G}$ such that the range map $\mathop{\boldsymbol{r}}\nolimits\colon\mathscr{G}\to\mathscr{G}^{(0)}$ (or equivalently, the source map $\mathop{\boldsymbol{s}}$) is a local homeomorphism.
+A bisection of an étale groupoid $\mathscr{G}$ is an open subset $U\subseteq\mathscr{G}$ such that the restrictions $\mathop{\boldsymbol{r}}\nolimits|_{U}$ and $\mathop{\boldsymbol{s}}|_{U}$ are homeomorphisms onto an open set in $\mathscr{G}^{(0)}$.
 In particular, bisections are always Hausdorff subspaces.
@@
 If $x\in\mathscr{G}^{(0)}$, then the isotropy group at $x$ is
-$\mathscr{G}^{x}_{x}=\{\gamma\in\mathscr{G}\mid\mathop{\boldsymbol{s}}(\gamma)=x=\mathop{\boldsymbol{r}}(\gamma)\}.$
+$\mathscr{G}^{x}_{x}=\{\gamma\in\mathscr{G}\mid\mathop{\boldsymbol{s}}(\gamma)=x=\mathop{\boldsymbol{r}}\nolimits(\gamma)\}.$
 Then $\mathop{\mathrm{Iso}}(\mathscr{G})=\bigcup_{x\in\mathscr{G}^{(0)}}\mathscr{G}^{x}_{x}$ is a closed subgroupoid called the isotropy subgroupoid (or bundle).
@@
 A groupoid with trivial isotropy groups is called principal.
-If $X\subseteq\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is the subgroupoid of all arrows $\gamma$ with $\mathop{\boldsymbol{s}}(\gamma),\mathop{\boldsymbol{r}}(\gamma)\in X$, i.e., $\mathscr{G}|_{X}=\mathop{\boldsymbol{s}}^{-1}(X)\cap\mathop{\boldsymbol{r}}^{-1}(X)$.
+If $X\subseteq\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is the subgroupoid of all arrows $\gamma$ with $\mathop{\boldsymbol{s}}(\gamma),\mathop{\boldsymbol{r}}\nolimits(\gamma)\in X$, i.e., $\mathscr{G}|_{X}=\mathop{\boldsymbol{s}}^{-1}(X)\cap\mathop{\boldsymbol{r}}\nolimits^{-1}(X)$.
 In particular, if $X$ is open (respectively, closed) in $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{X}$ is open (respectively, closed) in $\mathscr{G}$.
 The orbit of $x\in\mathscr{G}^{(0)}$ is the set $\mathcal{O}_{x}$ consisting of all $y\in\mathscr{G}^{(0)}$ such that there is an arrow $\gamma\colon x\to y$.
-Note that $\mathcal{O}_{x}=\mathop{\boldsymbol{r}}(\mathop{\boldsymbol{s}}^{-1}(x))=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(x))$.
+Note that $\mathcal{O}_{x}=\mathop{\boldsymbol{r}}\nolimits(\mathop{\boldsymbol{s}}^{-1}(x))=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}\nolimits^{-1}(x))$.
 The set of all orbits of $\mathscr{G}$ is denoted $\mathscr{G}^{(0)}/\mathscr{G}$.
@@
 $UV=\{\gamma\eta\in\mathscr{G}\mid\gamma\in U,\eta\in V\},\text{ and }U^{-1}=\{\gamma^{-1}\mid\gamma\in U\},$
-respectively. The semilattice of idempotents of $\Gamma_{c}(\mathscr{G})$ is given by the set of compact open subsets of $\mathscr{G}^{(0)}$.
-2.2. Algebras of ample groupoids
-Fix an ample groupoid $\mathscr{G}$ and a field $K$. The Steinberg algebra of $\mathscr{G}$ over $K$ is
-$K\mathscr{G}=\mathrm{span}_{K}\{1_{U}\mid U\in\Gamma_{c}(\mathscr{G})\},$
+respectively.
+The semilattice of idempotents of $\Gamma_{c}(\mathscr{G})$ is given by the set of compact open subsets of $\mathscr{G}^{(0)}$.
+Notice that if $\mathscr{G}$ is an ample group bundle, then $\Gamma_{c}(\mathscr{G})$ is a Clifford semigroup, as $\mathop{\boldsymbol{s}}(U)=\mathop{\boldsymbol{r}}\nolimits(U)$ for all $U\in\Gamma_{c}(\mathscr{G})$.
+2.2. Algebras of inverse semigroups and ample groupoids
+Fix a unital ring $R$ (not necessarily commutative). If $S$ is an inverse semigroup, then the inverse semigroup ring $RS$ is the free left $R$-module with basis $S$ and with multiplication
+$\sum_{s\in S}c_{s}s\cdot\sum_{t\in S}d_{t}t=\sum_{s,t\in S}c_{s}d_{t}st.$
+Fix an ample groupoid $\mathscr{G}$. Then the Steinberg algebra of $\mathscr{G}$ over $R$ is the left $R$-module
+$R\mathscr{G}=\mathrm{span}_{R}\{1_{U}\mid U\in\Gamma_{c}(\mathscr{G})\}\subseteq R^{\mathscr{G}},$
 with multiplication given by convolution:
-$f\ast g(\gamma)=\sum_{\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{s}}\alpha)}f(\gamma\alpha^{-1})g(\alpha)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
+$f\ast g(\gamma)=\sum_{\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{s}}(\alpha)}f(\gamma\alpha^{-1})g(\alpha)=\sum_{\alpha\beta=\gamma}f(\alpha)g(\beta).$
 Notice that $1_{U}*1_{V}=1_{UV}$ for all $U,V\in\Gamma_{c}(\mathscr{G})$.
-Thus $K\mathscr{G}$ is a quotient of the semigroup algebra $K\Gamma_{c}(\mathscr{G})$.
-If $K$ is a field with involution $\sigma$ (possibly trivial), then $K\mathscr{G}$ is a $\ast$-algebra with involution given by $f^{\ast}(\gamma)=\sigma(f(\gamma^{-1}))$.
-We shall only use this in the case $K=\mathbb{C}$ and $\sigma$ is complex conjugation.
-Note that $K\mathscr{G}$ is unital if and only if $\mathscr{G}^{(0)}$ is compact.
-See [31] for details.
-If $U$ is an open invariant subset of $\mathscr{G}^{(0)}$ and $X=\mathscr{G}^{(0)}\setminus U$ is the complementary closed invariant subspace, then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$ via restriction of functions by [10, Pages 1596–1597], a fact we shall use in the sequel without comment.
+Thus $R\mathscr{G}$ is a quotient of the semigroup algebra $R\Gamma_{c}(\mathscr{G})$.
+Note that $R\mathscr{G}$ is unital if and only if $\mathscr{G}^{(0)}$ is compact.
+See [33] for details (where the assumption that $R$ is commutative is not necessary).
+If $U$ is an open invariant subset of $\mathscr{G}^{(0)}$ and $X=\mathscr{G}^{(0)}\setminus U$ is the complementary closed invariant subspace, then $I=R\mathscr{G}|_{U}$ is an ideal of $R\mathscr{G}$ and $R\mathscr{G}/I\cong R\mathscr{G}|_{X}$ via restriction of functions by [10, Pages 1596–1597] (note an omission in the argument is resolved in [24, Corollar 4.4]), a fact we shall use in the sequel without comment.
 2.3. Groupoid of germs and the universal groupoid of an inverse semigroup
@@
 An action of $S$ on $X$ is a homomorphism $\varphi\colon S\to I_{X}$ such that if $X_{e}$ denotes the domain of an idempotent $e$, then $\bigcup_{e\in E}X_{e}=X$; this last condition says that the action is non-degenerate.
-Notice that $\psi_{s}\colon X_{s^{*}s}\to X_{ss^{*}}$ for each $s\in S$.
-We put $sx=\psi_{s}(x)$ for the action when no confusion can arise.
+Notice that $\varphi_{s}\colon X_{s^{*}s}\to X_{ss^{*}}$ for each $s\in S$.
+We put $sx=\varphi_{s}(x)$ for the action when no confusion can arise.
 If $S$ acts on a locally compact totally disconnected space $X$, the groupoid of germs $\mathscr{G}$ of the action is described as follows.
@@
 The unit space $\mathscr{G}_{S}^{(0)}=\{[e,x]\mid x\in X_{e}\}$ is identified with $X$ via the homeomorphism $[e,x]\mapsto x$.
-From this point of view, the domain and range maps are given by $\mathop{\boldsymbol{s}}([s,x])=x$ and $\mathop{\boldsymbol{r}}([s,x])=sx$.
+From this point of view, the domain and range maps are given by $\mathop{\boldsymbol{s}}([s,x])=x$ and $\mathop{\boldsymbol{r}}\nolimits([s,x])=sx$.
 There is a homomorphism $\psi\colon S\to\Gamma_{c}(\mathscr{G})$ given by $\psi(s)=(s,X_{s^{*}s})$ and $\mathscr{G}=\bigcup\psi(S)$.
@@
 $\alpha_{s}(x)(e)=x(s^{*}es)$
-is a homeomorphism, and the map $s\mapsto\alpha_{s}$ defines an action $\alpha\colon S\to I_{\widehat{E}}$ of $S$ on $\widehat{E}$ (see [29, Proposition 4.3.2] [^1] Note that Paterson works with right actions in [29].).
+is a homeomorphism, and the map $s\mapsto\alpha_{s}$ defines an action $\alpha\colon S\to I_{\widehat{E}}$ of $S$ on $\widehat{E}$ (see [30, Proposition 4.3.2] [^1] Note that Paterson works with right actions in [30].).
 The universal groupoid $\mathscr{G}_{S}$ associated with $S$ is the groupoid of germs $\mathscr{G}_{S}$ of the action of $S$ on $\widehat{E}$.
 For $s\in S$, define $\psi(s)=(s,D_{s^{*}s})$.
-Then $\psi$ is a homomorphism of $S$ into $\Gamma_{c}(\mathscr{G}_{S})$[29, Theorem 3.3.2] [^2] In fact, $\psi$ is a restriction of Paterson’s $\psi_{X}$ in [29, Theorem 3.3.2].
-In fact, $\psi$ is injective, and extends to an isomorphism $\widehat{\psi}\colon KS\to K\mathscr{G}_{S}$; see [31, Theorem 6.3].
+Then $\psi$ is a homomorphism of $S$ into $\Gamma_{c}(\mathscr{G}_{S})$[30, Theorem 3.3.2] [^2] In fact, $\psi$ is a restriction of Paterson’s $\psi_{X}$ in [30, Theorem 3.3.2].
+In fact, $\psi$ is injective, and extends to an isomorphism $\widehat{\psi}\colon RS\to R\mathscr{G}_{S}$; see [33, Theorem 6.3] where the assumption that $R$ is commutative is unnecessary.
 If $e\in E$ and $x_{e}$ is the principal character given by
@@
 2.4. Regular rings
-A ring $R$ is von Neumann regular, or regular for short, if for all $a\in R$, there exists $b\in R$ with $aba=a$. We do note require rings to be unital. The following facts are quite standard [19].
+A ring $A$ is von Neumann regular, or regular for short, if for all $a\in A$, there exists $b\in A$ with $aba=a$.
+We do not require rings to be unital, however the coefficient rings of our group, semigroup and groupoid algebras shall always be assumed unital.
+The following facts are quite standard; see [19, Chapter 1].
 Proposition 2.1.
@@
   1. (1)
-    The direct product of regular rings is regular.
+    An arbitrary direct product of regular rings is regular.
   2. (2)
-    If $R$ is a ring and $I$ is an ideal of $R$, then $R$ is regular if and only if $I$ and $R/I$ are regular.
+    If $A$ is a ring and $I$ is an ideal of $A$, then $A$ is regular if and only if $I$ and $A/I$ are regular.
   3. (3)
@@
   8. (8)
-    A ring $R$ is regular if and only if it has local units and each unitary $R$-module is flat.
-Recall that a semigroup is locally finite if all its finitely generated subsemigroups are finite, or, equivalently, it is a direct limit of finite semigroups.
+    A ring $A$ is regular if and only if it has local units and each unitary $A$-module is flat.
+A semigroup is locally finite if all its finitely generated subsemigroups are finite, or, equivalently, it is a direct limit of finite semigroups.
 Note that any subsemigroup of a locally finite semigroup is locally finite.
@@
 Theorem 2.3 (Connell).
-Let $G$ be a group and $R$ a ring. Then the group ring $RG$ is regular if and only if $R$ is regular, $G$ is locally finite and the order of each element of $G$ is a unit in $R$.
-Attempts have been made to extend this to inverse semigroups. The best result to date is in Okniński [27, Chapter 15].
-Theorem 2.4.
-Let $K$ be a field and $S$ an inverse semigroup.
-  1. (1)
-    If $K$ has characteristic 00, then $KS$ is regular if and only if $S$ is locally finite.
-  2. (2)
-    If $K$ has characteristic $p>0$, $S$ is locally finite and $p$ does not divide the order of any element of any maximal subgroup of $S$, then $KS$ is regular.
-  3. (3)
-    If $KS$ is regular and $K$ has characteristic $p>0$, then $S$ is periodic and each maximal subgroup of $S$ is locally finite with no element of order divisible by $p$.
-It is not known in positive characteristic if $KS$ regular implies that $S$ is locally finite for inverse semigroups in general, although this is known for Clifford semigroups, cf.
-[28, Chapter 15, Proposition 11].
-Since we wish to prove the difficult implications of Theorem 2.4 as consequences of our main results, we prove the straightforward implications of Theorem 2.4, which we shall need in our arguments.
-Proposition 2.5.
-Let $K$ be a field and $S$ a locally finite inverse semigroup such that the characteristic of $K$ does not divide the order of any element of any maximal subgroup of $S$. Then $KS$ is regular.
+Let $G$ be a group and $R$ a unital ring. Then the group ring $RG$ is regular if and only if $R$ is regular, $G$ is locally finite and the order of each finite subgroup of $G$ is invertible in $R$.
+Our main result extends Connell’s theorem to ample groupoids.
+We shall need the following result going back to Weissglass [34]. We sketch a proof for completeness.
+Proposition 2.4 (Weissglass).
+Let $R$ be a regular unital ring and let $S$ a locally finite inverse semigroup such that if a prime $p$ is the order of a cyclic subgroup of $S$, then $p$ is invertible in $R$. Then $RS$ is regular.
 Proof.
-Since the class of regular rings is closed under direct limits by Proposition 2.1, we may assume that $S$ is a finite inverse semigroup and $K$ is a field whose characteristic does not divide the order of any maximal subgroup of $S$.
-But then $KS$ is semisimple (in fact, isomorphic to a direct product of matrix algebras over group algebras of maximal subgroups) and hence regular, cf.
-[32, Corollary 9.4].
+Since the class of regular rings is closed under direct limits by Proposition 2.1, we may assume that $S$ is a finite inverse semigroup.
+Then it is well known that $RS$ is isomorphic to a finite direct product of matrix algebras over the group rings of its maximal subgroups (cf.
+[32, Theorem 4.6] where the assumption that $R$ is commutative is superfluous).
+Our assumptions imply that the order of any maximal subgroup $G$ of $S$ is invertible in $R$ by Cauchy’s theorem, and so $RG$ is regular by Theorem 2.3.
+Thus $RS$ is regular by Proposition 2.1.
 ∎
 3. von Neumann regularity of Steinberg algebras
-The following result is proved in [4]. We provide a proof for the reader’s convenience.
-Theorem 3.1.
-Let $K$ be a field and $\mathscr{G}$ an ample groupoid.
-If $K\mathscr{G}$ is regular, then each isotropy group of $\mathscr{G}$ is locally finite and the characteristic of $K$ divides the order of no isotropy group element.
+We begin by establishing the necessity of the conditions in Theorem A.
+Proposition 3.1.
+Suppose that $R\mathscr{G}$ is regular with $\mathscr{G}$ ample and $R$ a unital ring. Then $R$ is regular.
 Proof.
-By Connell’s theorem, it is enough to show that $K\mathscr{G}_{x}^{x}$ is regular for each $x\in\mathscr{G}^{(0)}$.
-Let $a\in K\mathscr{G}_{x}^{x}$, which we view as a finitely supported function $\mathscr{G}_{x}^{x}\to K$.
-Choose a compact open bisection $U_{\gamma}$ containing $\gamma$ for each $\gamma\in\mathscr{G}_{x}^{x}$ with $a(\gamma)\neq 0$.
-Consider $f=\sum_{\gamma\in\mathscr{G}^{x}_{x}}a(\gamma)U_{\gamma}$.
-By construction, $f|_{\mathscr{G}_{x}^{x}}=a$.
-Since $K\mathscr{G}$ is regular, we can find $g\in K\mathscr{G}$ with $fgf=f$.
-Notice that if $\alpha\in\mathscr{G}_{x}^{x}$, then
-$f(\alpha)=\sum_{\beta\gamma\nu=\alpha}f(\beta)g(\gamma)f(\nu).$ (3.1)
-But if $\alpha=\beta\gamma\nu$, then by construction of $f$ and the definition of a bisection if $f(\beta)\neq 0\neq f(\nu)$, we must have $\beta,\nu\in\mathscr{G}^{x}_{x}$ and hence $\gamma\in\mathscr{G}^{x}_{x}$.
-Therefore, (3.1) becomes
-$a(\alpha)=f(\alpha)=\sum_{\beta,\gamma,\nu\in\mathscr{G}_{x}^{x},\beta\gamma\nu=\alpha}a(\beta)g(\gamma)a(\nu).$
-Thus if we put $b=g|_{\mathscr{G}_{x}^{x}}$, then $b$ has finite support and $aba=a$. ∎
-We aim to improve on this, particularly in characteristic zero, where we can use some analysis. We shall also see that the necessary condition in Theorem 3.1 is by no means sufficient.
-Let us say that an ample groupoid $\mathscr{G}$ is uniformly bounded if there is an integer $M\geq 1$ with $|\mathop{\boldsymbol{r}}^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$.
-Note that any open or closed subgroupoid of a uniformly bounded ample groupoid is uniformly bounded.
+Let $r\in R$ and let $U$ be any nonempty compact open subset of $\mathscr{G}^{(0)}$.
+Let $f=r1_{U}$.
+Then there exists $g\in R\mathscr{G}$ with $fgf=f$.
+Let $x\in U$.
+Then $r=f(x)=fgf(x)=rg(x)r$, as $x=\alpha\beta\gamma$ with $\alpha,\gamma\in U$ implies $\alpha=x=\gamma$, and hence $\beta=x$.
+We conclude that $R$ is regular.
+∎
+Let us say that an ample groupoid $\mathscr{G}$ is uniformly bounded if there is an integer $M\geq 1$ with $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$ or, equivalently, $|\mathop{\boldsymbol{s}}^{-1}(x)|\leq M$ for all $x\in\mathscr{G}^{(0)}$.
+Note that any open or closed subgroupoid of a uniformly bounded ample groupoid is uniformly bounded, and so, in particular, a uniformly bounded groupoid has finite isotropy groups.
 Let us say that $\mathscr{G}$ is approximately uniformly bounded if it is a directed union of uniformly bounded open subgroupoids.
-An open or closed subgroupoid of an approximately uniformly bounded ample groupoid is again approximately uniformly bounded.
+An open or closed subgroupoid of an approximately uniformly bounded ample groupoid is again approximately uniformly bounded, and hence the isotropy groups of such a groupoid are locally finite.
 Let us say that $\mathscr{G}$ is approximately quasi-compact if $\mathscr{G}$ is a directed union of quasi-compact open subgroupoids.
-It is not difficult to verify that a principal second countable Hausdorff ample groupoid is approximately quasi-compact if and only if it is an approximately finite groupoid in the sense of [26, Section 11.5] (i.e., an approximately finite étale equivalence relation in the sense of [18]).
-Note that if $\mathscr{G}$ is a quasi-compact ample groupoid, then $\mathscr{G}^{(0)}$ is compact being the image of $\mathscr{G}$ under $\mathop{\boldsymbol{r}}$.
+It is not difficult to verify that a principal second countable Hausdorff ample groupoid is approximately quasi-compact if and only if it is an approximately finite groupoid in the sense of [28, Section 11.5] (i.e., an approximately finite étale equivalence relation in the sense of [18]).
+Note that if $\mathscr{G}$ is a quasi-compact ample groupoid, then $\mathscr{G}^{(0)}$ is compact being the image of $\mathscr{G}$ under $\mathop{\boldsymbol{r}}\nolimits$.
 The following proposition will be used without comment.
@@
 We may cover $C$ by finitely many compact open bisections $U_{1}\cup\cdots\cup U_{n}$ of $\mathscr{G}$.
-Then we have $|\mathop{\boldsymbol{r}}^{-1}(x)\cap\mathscr{H}|\leq n$ for all $x\in\mathscr{H}^{(0)}$ as $|\mathop{\boldsymbol{r}}^{-1}(x)\cap U_{i}|\leq 1$ for $i=1,\ldots,n$ and $\mathscr{H}\subseteq U_{1}\cup\cdots\cup U_{n}$.
+Then we have $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)\cap\mathscr{H}|\leq n$ for all $x\in\mathscr{H}^{(0)}$ as $|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)\cap U_{i}|\leq 1$ for $i=1,\ldots,n$ and $\mathscr{H}\subseteq U_{1}\cup\cdots\cup U_{n}$.
 ∎
@@
 Thus we may assume without loss of generality that $\mathscr{G}$ is uniformly bounded.
-Suppose that $|\mathop{\boldsymbol{r}}(x)|\leq n$ for all $x\in\mathscr{G}^{(0)}$.
-Then, for each $x\in\mathscr{G}^{(0)}$, we have that $|\mathscr{G}^{x}_{x}|\cdot|\mathcal{O}_{x}|=|\mathop{\boldsymbol{r}}^{-1}(x)|\leq n$.
-It follows that $|\mathscr{G}|_{\mathcal{O}_{x}}|=|\mathcal{O}_{x}|^{2}\cdot|\mathscr{G}_{x}^{x}|\leq n^{2}$, and hence $\mathscr{G}|_{\mathcal{O}_{x}}$ is a finite closed subgroupoid and $\Gamma_{c}(\mathscr{G}|_{\mathcal{O}_{x}})$ is a finite inverse semigroup of cardinality at most $2^{n^{2}}$.
+Suppose that $|\mathop{\boldsymbol{r}}\nolimits(x)|\leq n$ for all $x\in\mathscr{G}^{(0)}$.
+Then, for each $x\in\mathscr{G}^{(0)}$, we have that $|\mathscr{G}^{x}_{x}|\cdot|\mathcal{O}_{x}|=|\mathop{\boldsymbol{r}}\nolimits^{-1}(x)|\leq n$.
+It follows that $|\mathscr{G}|_{\mathcal{O}_{x}}|=|\mathcal{O}_{x}|^{2}\cdot|\mathscr{G}_{x}^{x}|\leq n^{2}$, and hence $\mathscr{G}|_{\mathcal{O}_{x}}$ is a finite closed subgroupoid, and $\Gamma_{c}(\mathscr{G}|_{\mathcal{O}_{x}})$ is a finite inverse semigroup of cardinality at most $2^{n^{2}}$.
 We have an injective homomorphism $\psi\colon\Gamma_{c}(\mathscr{G})\to\prod_{\mathcal{O}\in\mathscr{G}^{(0)}/\mathscr{G}}\Gamma_{c}(\mathscr{G}|_{\mathcal{O}})$ given by $\psi(U)_{\mathcal{O}}=U\cap\mathscr{G}|_{\mathcal{O}}$.
@@
 ∎
-To prove our main result, let us first recall the construction of the reduced $C^{*}$-algebra of an ample groupoid as described in [11].
-Let $\mathscr{G}$ be an ample groupoid.
-Then $\mathbb{C}\mathscr{G}$ is a $\ast$-algebra with involution $f^{\ast}(\gamma)=\overline{f(\gamma^{-1})}$.
-For each $x\in\mathscr{G}^{(0)}$, we consider a $\ast$-representation $\pi_{x}\colon\mathbb{C}\mathscr{G}\to\mathcal{B}(\ell_{2}(\mathscr{G}_{x}))$ given on the orthonormal basis $\delta_{\gamma}$, $\gamma\in\mathscr{G}_{x}$, by
-$\pi_{x}(f)\delta_{\gamma}=\sum_{\alpha\in\mathop{\boldsymbol{s}}^{-1}(\mathop{\boldsymbol{r}}(\gamma))}f(\alpha)\delta_{\alpha\gamma}.$
-Notice that if $U$ is a compact open bisection, then $\|\pi_{x}(1_{U})\|\leq 1$.
-Indeed, $U$ induces a partial permutation $\psi_{U}$ of $\mathscr{G}_{x}$ by observing that if $\gamma\in\mathscr{G}_{x}$, then either $U\{\gamma\}$ is empty, in which case $\psi_{U}(\gamma)$ is undefined, or $U\{\gamma\}=\{\alpha\gamma\}$ where $\alpha$ is the unique element of $U$ with $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\gamma)$, in which case $\psi_{U}(\gamma)=\alpha\gamma$.
-Therefore if $a=\sum_{\gamma\in\mathscr{G}_{x}}c_{\gamma}\delta_{\gamma}\in\ell_{2}(\mathscr{G}_{x})$, then writing $\mathrm{dom}(\psi_{U})$ for the domain of $\psi_{U}$, we have that
-$\pi_{x}(1_{U})a=\sum_{\gamma\in\mathscr{G}_{x}}\sum_{\alpha\in\mathop{\boldsymbol{r}}^{-1}(\gamma)}1_{U}(\alpha)c_{\gamma}\delta_{\alpha\gamma}=\sum_{\gamma\in\mathrm{dom}(\psi_{U})}c_{\gamma}\delta_{\psi_{U}(\gamma)}$
-and so $\|\pi_{x}(1_{U})a\|\leq\|a\|$.
-Thus $\|\pi_{x}(1_{U})\|\leq 1$.
-It follows that if $f=\sum_{i=1}^{n}c_{i}1_{U_{i}}\in\mathbb{C}\mathscr{G}$ with the $U_{i}$ compact open bisections, then $\|\pi_{x}(f)\|\leq\sum_{i=1}^{n}|c_{i}|$, and so $\|\pi_{x}(f)\|\leq|f|$ where ||$\cdot$|| ⋅ | is the norm defined in [33] by
-$|f|=\inf\left\{\sum_{i=1}^{n}|d_{i}|:f=\sum_{i=1}^{n}d_{i}1_{U_{i}},U_{i}\in\Gamma_{c}(\mathscr{G})\right\}.$
-(Note that this latter norm also bounds the $I$-norm.) The reduced $C^{*}$-norm on $\mathbb{C}\mathscr{G}$ is defined by $\|f\|_{r}=\sup_{x\in\mathscr{G}^{(0)}}\|\pi_{x}(f)\|$.
-The reduced $C^{*}$-algebra $C^{*}_{r}(\mathscr{G})$ of $\mathscr{G}$ is then the completion of $\mathbb{C}\mathscr{G}$ with respect to $\|\cdot\|_{r}$; see [11] for details.
-The fact that the reduced norm is truly a norm, and hence $\mathbb{C}\mathscr{G}$ embeds in $C^{*}_{r}(\mathscr{G})$, follows from the following estimate that we shall need in any event.
+Now we prove conditions (2) and (3) of Theorem A are necessary.
 Proposition 3.6.
-If $f\in\mathbb{C}\mathscr{G}$, then $\|f\|_{\infty}\leq\|f\|_{r}\leq|f|$.
+Let $R$ be a unital ring and $\mathscr{G}$ an ample groupoid such that $R\mathscr{G}$ is von Neumann regular.
+Then $\mathscr{G}$ is approximately quasicompact and the order of any finite subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
 Proof.
-We already observed that $\|f\|_{r}\leq|f|$.
-If $f\in\mathbb{C}\mathscr{G}$ and $\gamma\in\mathscr{G}$, then for $x=\mathop{\boldsymbol{s}}(\gamma)$, we have $\langle\pi_{x}(f)\delta_{x},\delta_{\gamma}\rangle=f(\gamma)$, and so by the Cauchy-Schwarz inequality $|f(\gamma)|\leq\|\pi_{x}(f)\|\leq\|f\|_{r}$.
-∎
-We now aim to characterize ample groupoids with regular algebras over fields of characteristic zero, generalizing Theorem 2.4.
-Our proof is inspired by Okniński [28, Chapter 13, Lemma 2], who traces his approach to an idea of Formanek.
-We shall need the well-known fact that any countable field $L$ of characteristic 00 can be embedded into $\mathbb{C}$.
-Indeed, choose a transcendence basis $S$ for $L/\mathbb{Q}$.
-Then $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{Q}(S)/\mathbb{Q}$ is purely transcendental.
-Since $L$ is countable, $S$ is also countable.
-On the other hand, $\mathbb{C}$ is uncountable and hence has an uncountable transcendence basis $T$ over $\mathbb{Q}$.
-Thus we can find an injective map $S\to T$ which extends to a field embedding $\mathbb{Q}(S)\hookrightarrow\mathbb{C}$.
-Since $L/\mathbb{Q}(S)$ is algebraic and $\mathbb{C}$ is algebraically closed, we can then extend this embedding to $L$.
-Theorem 3.7.
-Let $K$ be a field of characteristic zero and $\mathscr{G}$ an ample groupoid. Then the following are equivalent.
-  1. (1)
-    $K\mathscr{G}$ is regular.
-  2. (2)
-    $\mathscr{G}$ is approximately uniformly bounded.
-  3. (3)
-    $\mathscr{G}$ is approximately quasi-compact.
-  4. (4)
-    $\Gamma_{c}(\mathscr{G})$ is locally finite.
-  5. (5)
-    There is a locally finite inverse semigroup $S\subseteq\Gamma_{c}(\mathscr{G})$ with $\bigcup S=\mathscr{G}$.
-Proof.
-We already know that (2)–(5) are equivalent by Proposition 3.4.
-Let $S=\Gamma_{c}(\mathscr{G})$ be the inverse semigroup of compact open bisections of $\mathscr{G}$.
-Note that (4) implies that $KS$ is regular by Proposition 2.5, and hence $K\mathscr{G}$ is regular by Proposition 2.1, as it is quotient of $KS$.
-Thus (4) implies (1).
-It remains to show that (1) implies (2).
-Note that if $U$ is any compact open subset of $\mathscr{G}^{(0)}$, then $U\mathscr{G}U=\mathscr{G}|_{U}$ is an open subgroupoid with compact unit space $U$ and $K\mathscr{G}|_{U}=1_{U}K\mathscr{G}1_{U}$, and hence is regular by Proposition 2.1.
-Since $\mathscr{G}$ is the directed union of the $\mathscr{G}|_{U}$ as $U$ runs over all compact open subsets of $\mathscr{G}^{(0)}$, we may assume without loss of generality that $\mathscr{G}^{(0)}$ is compact, and hence $K\mathscr{G}$ is unital, by Propositions 3.2 and 3.4.
-Note that if $T$ is any inverse subsemigroup of $S$, then $\bigcup T$ is an open subgroupoid of $\mathscr{G}$.
-Also, we have that $\mathscr{G}=\bigcup S$.
-Since $S$ is a directed union of its finitely generated inverse subsemigroups, it suffices by Proposition 3.4 to prove that $\bigcup T$ is uniformly bounded for every finitely generated inverse subsemigroup $T$ of $S$.
-So let $X\subseteq S$ be a finite subset closed under inversion.
-We will show that $\mathscr{H}=\bigcup\langle X\rangle$ is uniformly bounded.
-Say $X=\{U_{1},\ldots,U_{n}\}$ and let $a=\frac{1}{2n}(1_{U_{1}}+\cdots+1_{U_{n}})\in K\mathscr{G}$.
-Then since $K\mathscr{G}$ is regular, we can find $b\in K\mathscr{G}$ such that $(1-a)b(1-a)=(1-a)$.
-We can write $b=c_{1}1_{V_{1}}+\cdots+c_{k}1_{V_{k}}$ with $V_{1},\ldots,V_{k}$ compact open bisections.
-Then $L=\mathbb{Q}(c_{1},\ldots,c_{k})$ is a finitely generated extension of $\mathbb{Q}$ such that $a,b\in L\mathscr{G}$.
-Since $L$ is countable, we can embed it over $\mathbb{Q}$ into $\mathbb{C}$ and hence $a,b\in\mathbb{C}\mathscr{G}$.
-Note that by Proposition 3.6, we have that $\|a\|_{r}\leq|a|\leq\frac{1}{2}$, and so $\sum_{m=0}^{\infty}a^{m}$ converges to $(1-a)^{-1}$ in $C^{*}_{r}(\mathscr{G})$.
-But in $\mathbb{C}\mathscr{G}\subseteq C^{*}_{r}(\mathscr{G})$, we have $(1-a)b(1-a)=(1-a)$, and so we deduce that $b=(1-a)^{-1}$ by cancelling $1-a$ on both sides in $C^{*}_{r}(\mathscr{G})$.
-Thus $b=\sum_{m=0}^{\infty}a^{m}$ in $C^{*}_{r}(\mathscr{G})$, that is, $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{r}=0$.
-But then since $\|f\|_{\infty}\leq\|f\|_{r}$, for $f\in\mathbb{C}\mathscr{G}$ by Proposition 3.6, we have $\lim_{N\to\infty}\||b-\sum_{m=0}^{N}a^{m}\|_{\infty}=0$.
-From the definition of $a$, we see that
-$a^{m}=\frac{1}{(2n)^{m}}\sum_{U_{i_{1}},\ldots,U_{i_{m}}\in X}1_{U_{i_{1}}\cdots U_{i_{m}}}.$
-Now if $\gamma\in\mathscr{H}=\bigcup\langle X\rangle$, then $\gamma\in U_{i_{1}}\cdots U_{i_{r}}$ for some $U_{i_{1}},\ldots,U_{i_{r}}\in X$ (not necessarily distinct).
-But then $\sum_{m=0}^{N}a^{m}(\gamma)\geq a^{r}(\gamma)\geq\frac{1}{(2n)^{r}}$ for any $N\geq r$.
-It follows that $b(\gamma)\neq 0$ for all $\gamma\in\mathscr{H}$.
-Thus $\mathscr{H}$ is contained in the quasi-compact subspace $V_{1}\cup\cdots\cup V_{k}$ of $\mathscr{G}$.
-It follows from Lemma 3.3 that $\mathscr{H}$ is uniformly bounded, completing the proof.
-∎
-Remark 3.8.
-Note that Theorem 3.7 gives us some additional information about the structure of $K\mathscr{G}$ when it is regular in characteristic 00.
-Namely, $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple subalgebras.
-In particular, if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial (i.e., a direct limit of finite direct products of matrix algebras over $K$).
-Indeed, if $\Gamma_{c}(\mathscr{G})$ is locally finite, then since $KT$ is a finite dimensional semisimple $K$-algebra for any finite inverse subsemigroup of $T$ [32, Chapter 9], we see that $K\Gamma_{c}(\mathscr{G})$ is of the desired form.
-Since a quotient algebra of such a directed union is again such a directed union, it follows that $K\mathscr{G}$ is of this form.
-In an attempt to understand the positive characteristic case, we will give sufficient conditions for the algebra of an approximately quasi-compact ample groupoid to be regular.
+First note that if $U$ is a compact open subset of $\mathscr{G}^{(0)}$, then $\mathscr{G}|_{U}$ is an open subgroupoid with compact unit space and $R\mathscr{G}|_{U}=1_{U}R\mathscr{G}1_{U}$.
+Hence if $R\mathscr{G}$ is regular, then so is $R\mathscr{G}|_{U}$ by Proposition 2.1.
+Since $\mathscr{G}$ is the directed union of open subgroupoids of the form $\mathscr{G}|_{U}$, we may assume without loss of generality that $\mathscr{G}^{(0)}$ is compact, that is, $R\mathscr{G}$ has the identity element $1_{\mathscr{G}^{(0)}}$.
+By Proposition 3.4, it suffices to show that $\mathscr{G}$ is approximately uniformly bounded, and that the order of any finite subgroup of an isotropy group is invertible in $R$.
+If $S\leq\Gamma_{c}(\mathscr{G})$ is an inverse submonoid, then $\bigcup S$ is an open subgroupoid of $\mathscr{G}$ containing $\mathscr{G}^{(0)}$.
+Moreover, since $\Gamma_{c}(\mathscr{G})$ is the directed union of its finitely generated inverse submonoids and $\mathscr{G}=\bigcup\Gamma_{c}(\mathscr{G})$, it suffices to show that if $S\leq\Gamma_{c}(\mathscr{G})$ is a finitely generated inverse submonoid, then $\mathscr{H}=\bigcup S$ is uniformly bounded and the order of any isotropy group of $\mathscr{H}$ is invertible in $R$.
+We proceed by adapting the idea of Connell for the case of groups [12].
+Let $I_{S}$ be the left ideal of $R\mathscr{G}$ generated by the elements $1_{U}-1_{UU^{-1}}$ with $U\in S$.
+Let $X$ be a finite generating set of $S$ as a semigroup which is closed under taking inverses.
+We prove that $I_{S}$ is generated as a left ideal by the finite set $Y$ of elements of the form $1_{V}-1_{VV^{-1}}$ with $V\in X$.
+We proceed by induction on word length, i.e., the minimum length of an expression of an element $U$ of $S$ as a product of elements of $X$, to show that $1_{U}-1_{UU^{-1}}\in K\mathscr{G}\cdot Y$, with the case of length $1$ being trivial.
+If $U=WV$ with $V\in X$ and $W\in S$ of length one shorter than $U$, then $W^{-1}$ also has length one shorter than $U$ (since $X$ is closed under inversion), and so
+$1_{U}-1_{UU^{-1}}=1_{W}(1_{V}-1_{VV^{-1}})-1_{WVV^{-1}}(1_{W^{-1}}-1_{W^{-1}W})\in K\mathscr{G}\cdot Y$
+by induction, as required.
+Since $R\mathscr{G}$ is a regular ring with identity and $I_{S}$ is a finitely generated left ideal, there is an idempotent $e\in R\mathscr{G}$ with $I_{S}=R\mathscr{G}e$ by [19, Theorem 1.1].
+Let $f=1_{\mathscr{G}^{(0)}}-e$.
+Then $f$ is an idempotent and $I_{S}f=R\mathscr{G}e(1-e)=0$, that is, $1_{U}f=1_{UU^{-1}}f$ for all $U\in S$.
+If $f=\sum_{i=1}^{r}c_{i}1_{U_{i}}$ with $U_{i}$ compact open bisections, then $f$ can be nonzero on at most $r$ elements from the same $\mathop{\boldsymbol{s}}$-fiber, a fact that we shall soon use.
+Let $\gamma\in\mathscr{H}$ and $\beta\in\mathscr{G}$ with $\mathop{\boldsymbol{s}}(\gamma)=\mathop{\boldsymbol{r}}\nolimits(\beta)$.
+We claim that $f(\gamma\beta)=f(\beta)$.
+Indeed, let $U\in S$ with $\gamma\in U$.
+Then $(1_{U}f)(\gamma\beta)=f(\beta)$ and $(1_{UU^{-1}})f(\gamma\beta)=f(\gamma\beta)$.
+We conclude that $f(\gamma\beta)=f(\beta)$ for all $\gamma\in\mathscr{H}\cap\mathop{\boldsymbol{s}}^{-1}(\mathop{\boldsymbol{r}}\nolimits(\beta))$.
+Let $\mathop{\boldsymbol{r}}\nolimits_{*}\colon R\mathscr{G}\to R\mathscr{G}^{(0)}$ be the left $R$-module homomorphism induced by the local homeomorphism $\mathop{\boldsymbol{r}}\nolimits$. So
+$\mathop{\boldsymbol{r}}\nolimits_{*}(f)(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\gamma)=x}f(\gamma).$
+Note that if $U\in\Gamma_{c}(\mathscr{G})$, then $\mathop{\boldsymbol{r}}\nolimits_{*}(1_{U})=1_{UU^{-1}}$.
+It is well known and easy to check that $\mathop{\boldsymbol{r}}\nolimits_{*}$ is a left $R\mathscr{G}$-module homomorphism where $R\mathscr{G}$ acts on $R\mathscr{G}^{(0)}$ via the rule
+$f_{1}\cdot g_{1}(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\gamma)=x}f_{1}(\gamma)g_{1}(\mathop{\boldsymbol{s}}(\gamma))$
+for $f_{1}\in R\mathscr{G}$ and $g_{1}\in R\mathscr{G}^{(0)}$. Notice that if $U\in\Gamma_{c}(\mathscr{G})$ and $V\subseteq\mathscr{G}^{(0)}$ is compact open, then $1_{U}\cdot 1_{V}=1_{UVU^{-1}}$.
+If $U\in S$, then $\mathop{\boldsymbol{r}}\nolimits_{*}(1_{U}-1_{UU^{-1}})=0$, and so $I_{S}\subseteq\ker\mathop{\boldsymbol{r}}\nolimits_{*}$.
+Therefore, $\mathop{\boldsymbol{r}}\nolimits_{*}(f)=\mathop{\boldsymbol{r}}\nolimits_{*}(1_{\mathscr{G}^{(0)}}-e)=\mathop{\boldsymbol{r}}\nolimits_{*}(1_{\mathscr{G}^{(0)}})=1_{\mathscr{G}^{(0)}}$, as $e\in I_{S}$.
+Let $x\in\mathscr{G}^{(0)}$.
+Then
+$1=\mathop{\boldsymbol{r}}\nolimits_{*}(f)(x)=\sum_{\mathop{\boldsymbol{r}}\nolimits(\beta)=x}f(\beta)$ (3.1)
+and so we can find $\beta\in\mathscr{G}$ with $\mathop{\boldsymbol{r}}\nolimits(\beta)=x$ and $f(\beta)\neq 0$.
+Then, for every $\gamma\in\mathscr{H}\cap\mathop{\boldsymbol{s}}^{-1}(x)$, we have $f(\gamma\beta)=f(\beta)\neq 0$.
+Since $f$ is nonzero on at most $r$ elements with source $\mathop{\boldsymbol{s}}(\beta)$, we deduce that $|\mathop{\boldsymbol{s}}^{-1}(x)\cap\mathscr{H}|<r+1$.
+Finally, the group $\mathscr{H}_{x}^{x}$ acts freely on the left of $\mathop{\boldsymbol{r}}\nolimits^{-1}(x)$, and so if we choose a transversal $T$ for this action and use $f(\gamma\beta)=f(\beta)$ for $\gamma\in\mathscr{H}_{x}^{x}$, we obtain from (3.1)
+$1=|\mathscr{H}_{x}^{x}|\cdot\sum_{t\in T}f(t),$
+and so $|\mathscr{H}_{x}^{x}|$ is invertible in $R$. This completes the proof. ∎
+We now prove the sufficiency of the conditions in Theorem A.
 Proposition 3.7.
-Let $\mathscr{G}$ be a quasi-compact ample group bundle and $K$ a field whose characteristic does not divide the order of any isotropy group element. Then $K\mathscr{G}$ is regular.
+Let $\mathscr{G}$ be a quasi-compact ample group bundle and $R$ a regular unital ring such that the order of any cyclic subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
+Then $R\mathscr{G}$ is regular.
 Proof.
 We know that $S=\Gamma_{c}(\mathscr{G})$ is locally finite by Proposition 3.4.
-We show that no maximal subgroup of $S$ contains an element of order divisible by the characteristic of $K$.
-It will then follow that $KS$ is regular by Proposition 2.5, and hence its quotient $K\mathscr{G}$ is regular by Proposition 2.1.
-Let $U\subseteq\mathscr{G}^{(0)}$ be the identity of a maximal subgroup $H$, and note that $H$ consists of those $V\in S$ with $\mathop{\boldsymbol{s}}(V)=U=\mathop{\boldsymbol{r}}(V)$.
-If $V\in H$ and $x\in U$, then denote by $v_{x}$ the unique element of $V$ with $\mathop{\boldsymbol{r}}(v_{x})=x$.
+We show that $RS$ is regular.
+It will then follow that its quotient $R\mathscr{G}$ is regular by Proposition 2.1.
+By Proposition 2.4 it suffices to show that if a prime $p$ is the order of a cyclic subgroup of $S$, then $p$ is invertible in $R$.
+Let $U\subseteq\mathscr{G}^{(0)}$ be the identity of a maximal subgroup $H$, and note that $H$ consists of those $V\in S$ with $\mathop{\boldsymbol{s}}(V)=U=\mathop{\boldsymbol{r}}\nolimits(V)$.
+If $V\in H$ and $x\in U$, then denote by $v_{x}$ the unique element of $V$ with $\mathop{\boldsymbol{r}}\nolimits(v_{x})=x$.
 Since $\mathscr{G}$ consists of isotropy, $\mathop{\boldsymbol{s}}(v_{x})=x$ as well.
 Then we have an injective group homomorphism $\psi\colon H\to\prod_{x\in U}\mathscr{G}_{x}^{x}$ given by $\psi(V)_{x}=v_{x}$.
-Suppose that $K$ has characteristic $p>0$.
-Note that if $V\in H$ satisfies $V^{p}=U$, then $\psi(V)^{p}=\psi(U)$ and hence $v_{x}^{p}=x$ for all $x\in U$.
-Since no isotropy group has an element of order $p$, we deduce that $v_{x}=x$ for all $x\in U$, and hence $V=U$.
-Thus no element of $H$ has order divisible by $p$.
-This concludes the proof.
+Note that if $V\in H$ has order $p$, then $\psi(V)$ has order $p$, and hence $v_{x}^{p}=x$ for all $x\in U$, and some $v_{x}$ has order $p$.
+Therefore, $p$ is invertible in $R$ by assumption.
 ∎
@@
 Lemma 3.8.
-Let $\mathscr{G}$ be a quasi-compact ample groupoid and $K$ a field whose characteristic does not divide the order of any isotropy group element.
+Let $\mathscr{G}$ be a quasi-compact ample groupoid and $R$ a regular unital ring such that the order of any cyclic subgroup of an isotropy group of $\mathscr{G}$ is invertible in $R$.
 Suppose that all orbits of $\mathscr{G}$ have the same cardinality.
-Then $K\mathscr{G}$ is regular.
+Then $R\mathscr{G}$ is regular.
 Proof.
 It follows from Lemma 3.3 that the size of each orbit is finite, say that this common size is $n$.
-Choose for each $x\in\mathscr{G}^{(0)}$, elements $\gamma_{x,1},\ldots,\gamma_{x,n}$ with $\mathop{\boldsymbol{s}}(\gamma_{x,i})=x$ and $\mathcal{O}_{x}=\{\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})\}$.
+Choose, for each $x\in\mathscr{G}^{(0)}$, elements $\gamma_{x,1},\ldots,\gamma_{x,n}$ with $\mathop{\boldsymbol{s}}(\gamma_{x,i})=x$ and $\mathcal{O}_{x}=\{\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,n})\}$.
 We may assume without loss of generality that $\gamma_{x,1}=x$ (the identity at $x$).
@@
 We may assume that $U^{x}_{1,1}\subseteq\mathscr{G}^{(0)}$.
-Since $\mathscr{G}^{(0)}$ is Hausdorff with a basis of compact open sets and $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x_{n}})$ are pairwise distinct, we can find pairwise disjoint compact open neighborhoods $V_{1},\ldots,V_{n}$ of $\mathop{\boldsymbol{r}}(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}(\gamma_{x,n})$, respectively, in $\mathscr{G}^{(0)}$.
-Replacing $U^{x}_{i,1}$ by $V_{i}U^{x}_{i,1}$, we may assume that the sets $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,1})$ are pairwise disjoint.
+Since $\mathscr{G}^{(0)}$ is Hausdorff with a basis of compact open sets and $\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x_{n}})$ are pairwise distinct, we can find pairwise disjoint compact open neighborhoods $V_{1},\ldots,V_{n}$ of $\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(\gamma_{x,n})$, respectively, in $\mathscr{G}^{(0)}$.
+Replacing $U^{x}_{i,1}$ by $V_{i}U^{x}_{i,1}$, we may assume that the sets $U^{x}_{i}=\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,1})$ are pairwise disjoint.
 Also, putting $U=\bigcap_{i=1}^{n}\mathop{\boldsymbol{s}}(U^{x}_{i,1})$ and replacing $U^{x}_{i,1}$ by $U^{x}_{i,1}U$, we may assume that $\mathop{\boldsymbol{s}}(U^{x}_{i,1})=\mathop{\boldsymbol{s}}(U^{x}_{1,1})=U^{x}_{1}$ for all $i$.
 Put $U^{x}_{i,j}=U^{x}_{i,1}(U^{x}_{j,1})^{-1}$.
-Note that $\mathop{\boldsymbol{s}}(U^{x}_{i,j})=U^{x}_{j}$ and $\mathop{\boldsymbol{r}}(U^{x}_{i,j})=U^{x}_{i}$ for all $i,j$.
+Note that $\mathop{\boldsymbol{s}}(U^{x}_{i,j})=U^{x}_{j}$ and $\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,j})=U^{x}_{i}$ for all $i,j$.
 Also note that $U^{x}_{i,i}=U^{x}_{i}$ for $i=1,\ldots,n$.
@@
 For suppose that $y\in U^{x}_{j}$.
-Then since $U^{x}_{j}=\mathop{\boldsymbol{s}}(U^{x}_{i,j})$ and $U^{x}_{i}=\mathop{\boldsymbol{r}}(U^{x}_{i,j})$ for all $i=1,\ldots,n$, we can find elements of the orbit of $y$ in each of $U^{x}_{1},\ldots,U^{x}_{n}$.
+Then since $U^{x}_{j}=\mathop{\boldsymbol{s}}(U^{x}_{i,j})$ and $U^{x}_{i}=\mathop{\boldsymbol{r}}\nolimits(U^{x}_{i,j})$ for all $i=1,\ldots,n$, we can find elements of the orbit of $y$ in each of $U^{x}_{1},\ldots,U^{x}_{n}$.
 Since these sets are pairwise disjoint, this yields $n$ elements in the orbit of $y$ and hence the whole orbit by assumption.
@@
 Another observation is that the isotropy subgroupoid $\mathop{\mathrm{Iso}}(\mathscr{G})$ of $\mathscr{G}$ is clopen, and hence quasi-compact.
-It is closed since $\mathscr{G}^{(0)}$ is Hausdorff and it is the equalizer of $\mathop{\boldsymbol{s}},\mathop{\boldsymbol{r}}$.
-It is open because if $\gamma\colon x\to x$ is isotropy and $V$ is a compact open bisection containing $\gamma$, then $U^{x}_{1}VU^{x}_{1}$ is a compact open bisection containing $\gamma$ and consisting of isotropy as $x\in U^{x}_{1}$, and $U^{x}_{1}$ contains exactly only one element from each orbit of the invariant compact open set $U^{x}$, whence $\mathop{\boldsymbol{s}}(\alpha),\mathop{\boldsymbol{r}}(\alpha)\in U^{x}_{1}$ implies $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}(\alpha)$.
+It is closed since $\mathscr{G}^{(0)}$ is Hausdorff and it is the equalizer of $\mathop{\boldsymbol{s}},\mathop{\boldsymbol{r}}\nolimits$.
+It is open because if $\gamma\colon x\to x$ is isotropy and $V$ is a compact open bisection containing $\gamma$, then $U^{x}_{1}VU^{x}_{1}$ is a compact open bisection containing $\gamma$ and consisting of isotropy as $x\in U^{x}_{1}$, and $U^{x}_{1}$ contains exactly only one element from each orbit of the invariant compact open set $U^{x}$, whence $\mathop{\boldsymbol{s}}(\alpha),\mathop{\boldsymbol{r}}\nolimits(\alpha)\in U^{x}_{1}$ implies $\mathop{\boldsymbol{s}}(\alpha)=\mathop{\boldsymbol{r}}\nolimits(\alpha)$.
 Since $\mathscr{G}^{(0)}$ is compact and the $U^{x}$ cover $\mathscr{G}^{(0)}$, we can find finitely may points $x_{1},\ldots,x_{r}$ such that $\mathscr{G}^{(0)}=U^{x_{1}}\cup\cdots\cup U^{x_{r}}$.
@@
 In particular note that $V_{i}U=UV_{i}$ for any compact open bisection $U$.
-It follows that if $\mathscr{G}_{i}=\mathscr{G}|_{V_{i}}$, then $K\mathscr{G}\cong\prod_{k=1}^{r}K\mathscr{G}_{k}$ and $K\mathscr{G}_{k}=1_{V_{k}}K\mathscr{G}1_{V_{k}}$.
-Note that since $\mathscr{G}_{k}=\mathop{\boldsymbol{r}}^{-1}(V_{k})$ is clopen in $\mathscr{G}$ it is quasi-compact.
+It follows that if $\mathscr{G}_{i}=\mathscr{G}|_{V_{i}}$, then $R\mathscr{G}\cong\prod_{k=1}^{r}R\mathscr{G}_{k}$ and $R\mathscr{G}_{k}=1_{V_{k}}R\mathscr{G}1_{V_{k}}$.
+Note that since $\mathscr{G}_{k}=\mathop{\boldsymbol{r}}\nolimits^{-1}(V_{k})$ is clopen in $\mathscr{G}$ it is quasi-compact.
 Also $\mathop{\mathrm{Iso}}(\mathscr{G}_{k})=\mathop{\mathrm{Iso}}(\mathscr{G})\cap\mathscr{G}_{k}$ is clopen and hence quasi-compact.
@@
 Also note that $U^{k}_{1,1}\subseteq U^{x_{k}}_{1}$ does not contain any two points in the same orbit by the previous discussion.
-Hence $U^{k}_{1,1}\mathscr{G}_{k}U^{k}_{1,1}=\mathscr{G}_{k}|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})\cap\mathop{\boldsymbol{r}}^{-1}(U^{k}_{1,1})$ is clopen (as $U^{k}_{1,1}$ is a clopen subspace of $\mathscr{G}^{(0)}$) and thus a quasi-compact group bundle.
-Note that $V_{k}=V_{k}U^{x_{k}}=\bigcup_{i=1}^{n}U^{k}_{i,i}$ and hence it follows from (3.2) that the $1_{U^{k}_{i,j}}$ form a set of matrix units for $K\mathscr{G}_{k}$.
-Thus $K\mathscr{G}_{k}\cong M_{n}(1_{U^{k}_{1,1}}K\mathscr{G}_{k}1_{U^{k}_{1,1}})=M_{n}(K\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}})$.
-We deduce that $K\mathscr{G}$ is regular from Propositions 2.1 and 3.9.
+Hence $U^{k}_{1,1}\mathscr{G}_{k}U^{k}_{1,1}=\mathscr{G}_{k}|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}}=\mathop{\mathrm{Iso}}(\mathscr{G}_{k})\cap\mathop{\boldsymbol{r}}\nolimits^{-1}(U^{k}_{1,1})$ is clopen (as $U^{k}_{1,1}$ is a clopen subspace of $\mathscr{G}^{(0)}$) and thus a quasi-compact group bundle.
+Note that $V_{k}=V_{k}U^{x_{k}}=\bigcup_{i=1}^{n}U^{k}_{i,i}$ and hence it follows from (3.2) that the $1_{U^{k}_{i,j}}$ form a set of matrix units for $R\mathscr{G}_{k}$.
+Thus $R\mathscr{G}_{k}\cong M_{n}(1_{U^{k}_{1,1}}R\mathscr{G}_{k}1_{U^{k}_{1,1}})=M_{n}(R\mathop{\mathrm{Iso}}(\mathscr{G}_{k})|_{U^{k}_{1,1}})$.
+We deduce that $R\mathscr{G}$ is regular from Propositions 2.1 and 3.7.
 ∎
@@
 Let $x=y_{1},y_{2}\ldots,y_{k}$ be $k$ distinct elements in the orbit of $x$.
-As in the proof of Lemma 3.10, we can find open bisections $U_{1},\ldots,U_{k}$ with $x\in\mathop{\boldsymbol{s}}(U_{i})$, $y_{i}\in\mathop{\boldsymbol{r}}(U_{i})$, for $i=1,\ldots,k$, and $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ pairwise disjoint (using that $\mathscr{G}^{(0)}$ is Hausdorff).
-Then $U=\bigcap_{i=1}^{k}\mathop{\boldsymbol{s}}(U_{i})$ is a neighborhood of $x$ and if $y\in U$, then each of $\mathop{\boldsymbol{r}}(U_{1}),\ldots,\mathop{\boldsymbol{r}}(U_{k})$ contains an element of the orbit of $y$.
+As in the proof of Lemma 3.8, we can find open bisections $U_{1},\ldots,U_{k}$ with $x\in\mathop{\boldsymbol{s}}(U_{i})$, $y_{i}\in\mathop{\boldsymbol{r}}\nolimits(U_{i})$, for $i=1,\ldots,k$, and $\mathop{\boldsymbol{r}}\nolimits(U_{1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(U_{k})$ pairwise disjoint (using that $\mathscr{G}^{(0)}$ is Hausdorff).
+Then $U=\bigcap_{i=1}^{k}\mathop{\boldsymbol{s}}(U_{i})$ is a neighborhood of $x$ and if $y\in U$, then each of $\mathop{\boldsymbol{r}}\nolimits(U_{1}),\ldots,\mathop{\boldsymbol{r}}\nolimits(U_{k})$ contains an element of the orbit of $y$.
 Since these sets are pairwise disjoint, we conclude that $y\in O_{k}$.
 ∎
-We can now prove our second main theorem.
-Theorem 3.12.
-Suppose that $K$ is a field and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that the order of no isotropy element is divisible by the characteristic of $K$.
-Then $K\mathscr{G}$ is regular.
+We can now prove the sufficiency of the conditions in Theorem A.
+Proposition 3.10.
+Suppose that $R$ is a regular unital ring and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that the order any finite subgroup of an isotropy group is invertible in $R$.
+Then $R\mathscr{G}$ is regular.
 Proof.
@@
 Let $X=\mathscr{G}^{(0)}\setminus U$.
-Then $I=K\mathscr{G}|_{U}$ is an ideal of $K\mathscr{G}$ and $K\mathscr{G}/I\cong K\mathscr{G}|_{X}$.
-Note that $\mathscr{G}|_{U}=\mathop{\boldsymbol{r}}^{-1}(U)$ is open and $\mathscr{G}|_{X}=\mathop{\boldsymbol{r}}^{-1}(X)$ is closed and hence quasi-compact.
-By construction, the orbits sizes in $\mathscr{G}|_{X}$ are $d_{1}<d_{2}<\cdots<d_{m-1}$, and so by induction $K\mathscr{G}|_{X}$ is regular.
-If we can show that $K\mathscr{G}|_{U}$ is regular, then Proposition 2.1 will imply that $K\mathscr{G}$ is regular.
+Then $I=R\mathscr{G}|_{U}$ is an ideal of $R\mathscr{G}$ and $R\mathscr{G}/I\cong R\mathscr{G}|_{X}$.
+Note that $\mathscr{G}|_{U}=\mathop{\boldsymbol{r}}\nolimits^{-1}(U)$ is open and $\mathscr{G}|_{X}=\mathop{\boldsymbol{r}}\nolimits^{-1}(X)$ is closed and hence quasi-compact.
+By construction, the orbits sizes in $\mathscr{G}|_{X}$ are $d_{1}<d_{2}<\cdots<d_{m-1}$, and so by induction $R\mathscr{G}|_{X}$ is regular.
+If we can show that $R\mathscr{G}|_{U}$ is regular, then Proposition 2.1 will imply that $R\mathscr{G}$ is regular.
 Let $V\subseteq U$ be a compact open set.
-Then $V$ is clopen in $\mathscr{G}^{(0)}$ and hence $\mathop{\boldsymbol{r}}^{-1}(V)\subseteq\mathscr{G}$ is clopen, hence quasi-compact and open.
-Thus $W=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}^{-1}(V))$ is a compact open invariant subset of $U$ containing $V$.
+Then $V$ is clopen in $\mathscr{G}^{(0)}$, and hence $\mathop{\boldsymbol{r}}\nolimits^{-1}(V)\subseteq\mathscr{G}$ is clopen, whence quasi-compact and open.
+Thus $W=\mathop{\boldsymbol{s}}(\mathop{\boldsymbol{r}}\nolimits^{-1}(V))$ is a compact open invariant subset of $U$ containing $V$.
 Hence the compact open invariant subsets of $U$ are cofinal among compact open subsets of $U$.
-Moreover, if $W$ is a compact open invariant subspace of $U$, then $W$ is clopen in $\mathscr{G}^{(0)}$ (as the latter is Hausdorff), and so $\mathscr{G}|_{W}=\mathop{\boldsymbol{r}}^{-1}(W)$ is a clopen subgroupoid of $\mathscr{G}$, hence quasi-compact, with all orbits of size $d_{m}$ (as $W\subseteq U$ is invariant).
-Hence $K\mathscr{G}|_{W}$ is regular by Lemma 3.10.
-Since $K\mathscr{G}|_{U}$ is the direct limit of the $K\mathscr{G}|_{W}$ where $W$ runs over the compact open invariant subsets of $U$, we conclude that $K\mathscr{G}|_{U}$ is regular from Proposition 2.1.
+Moreover, if $W$ is a compact open invariant subspace of $U$, then $W$ is clopen in $\mathscr{G}^{(0)}$ (as the latter is Hausdorff), and so $\mathscr{G}|_{W}=\mathop{\boldsymbol{r}}\nolimits^{-1}(W)$ is a clopen subgroupoid of $\mathscr{G}$, hence quasi-compact, with all orbits of size $d_{m}$ (as $W\subseteq U$ is invariant).
+Hence $R\mathscr{G}|_{W}$ is regular by Lemma 3.8.
+Since $R\mathscr{G}|_{U}$ is the direct limit of the $R\mathscr{G}|_{W}$ where $W$ runs over the compact open invariant subsets of $U$, we conclude that $R\mathscr{G}|_{U}$ is regular from Proposition 2.1.
 This completes the proof.
@@
 Remark 3.11.
-If one follows the proofs carefully, it is not difficult to see that if $K$ is a field of characteristic $p>0$ and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that no isotropy element has order divisible by $p$, then $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple algebras, and hence if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial.
+If one follows the proofs carefully, it is not difficult to see that if $K$ is a field and $\mathscr{G}$ is an approximately quasi-compact ample groupoid such that no isotropy element has order divisible by the characteristic of $K$, then $K\mathscr{G}$ is a directed union of finite dimensional unital semisimple algebras, and hence if $K$ is algebraically closed, then $K\mathscr{G}$ is ultramatricial.
 This just uses that the class under consideration is closed under matrix amplification and if an algebra has a finite ideal filtration with factors of this form, then the algebra has this form.
+We may now prove Theorem A.
+Proof of Theorem A.
+Necessity of (1)–(3) follows from Propositions 3.1 and 3.6. Sufficiency of these conditions follows from Proposition 3.10. ∎
 4. Graded von Neumann regularity of Steinberg algebras
@@
 In this section we characterize when ample groupoid algebras are graded von Neumann regular.
-Let $K$ be a field, $R$ a $K$-algebra and $H$ a group with identity $e$.
-Then $R$ is $H$-graded if $R$ has a direct sum decomposition $R=\bigoplus_{h\in H}R_{h}$, where each $R_{h}$ is a $K$-vector subspace of $R$ such that $R_{h}R_{k}\subseteq R_{hk}$.
-The summand $R_{h}$ is called the homogeneous component of degree $h$.
-The homogeneous component $R_{e}$ is a $K$-subalgebra of $R$.
+If $H$ is a group with identity $e$, then a ring $\Lambda$ is $H$-graded if $\Lambda$ has a direct sum decomposition $\Lambda=\bigoplus_{h\in H}\Lambda_{h}$, where each $\Lambda_{h}$ is an additive subgroup of $\Lambda$ such that $\Lambda_{h}\Lambda_{k}\subseteq\Lambda_{hk}$ for all $h,k\in H$.
+The summand $\Lambda_{h}$ is called the homogeneous component of degree $h$.
+The homogeneous component $\Lambda_{e}$ is a subring of $\Lambda$.
 An element is homogenous if it belongs to a homogeneous component.
-A graded ring $R$ is graded (von Neumann) regular if for every homogeneous element $x\in R$ there exists $y\in R$ such that $x=xyx$.
+A graded ring $\Lambda$ is graded (von Neumann) regular if for every homogeneous element $x\in\Lambda$ there exists $y\in\Lambda$ such that $x=xyx$.
 Let $\mathscr{G}$ be an ample groupoid and let $H$ be a discrete group.
-A cocycle from $\mathscr{G}$ to $H$ is a map $c\colon\mathscr{G}\to H$ such that $c(\gamma^{-1})=c(\gamma)^{-1}$, for all $\gamma\in\mathscr{G}$, and $c(\gamma\eta)=c(\gamma)c(\eta)$ for all $(\gamma,\eta)\in\mathscr{G}^{(2)}$.
-A continuous cocycle $c\colon\mathscr{G}\to H$ induces an an $H$-grading $\{(K\mathscr{G})_{h}\}_{h}$ of $K\mathscr{G}$.
+A cocycle from $\mathscr{G}$ to $H$ is a map $c\colon\mathscr{G}\to H$ such that $c(\gamma\eta)=c(\gamma)c(\eta)$ for all $(\gamma,\eta)\in\mathscr{G}^{(2)}$.
+A continuous cocycle $c\colon\mathscr{G}\to H$ induces an an $H$-grading $\{(R\mathscr{G})_{h}\}_{h}$ of $R\mathscr{G}$.
 This is done in detail in [9] for Hausdorff groupoids.
@@
 Call a compact open bisection $U\in\Gamma_{c}(\mathscr{G})$ homogenous of degree $h\in H$ if $U\subseteq c^{-1}(h)$ for some $h\in H$.
-Then $(K\mathscr{G})_{h}$ is the $K$-span of the characteristic functions of homogeneous compact open bisections of degree $h$.
-Note that each $c^{-1}(h)$ with $h\in H$ is clopen, and so if $U$ is any compact open bisection, then $U=\bigcup_{h\in H}(U\cap c^{-1}(h))$ and the $U\cap c^{-1}(h)$ are pairwise disjoint homogeneuous compact open bisections.
-By compactness of $U$, only finitely many of these intersections, say corresponding to $h_{1},\ldots,h_{n}$, are nonempty, and so $1_{U}=\sum_{i=1}^{n}1_{(c^{-1}(h_{i})\cap U)}$, showing that $\sum_{h\in H}(K\mathscr{G})_{h}=K\mathscr{G}$.
-The sum is direct since the $c^{-1}(h)$ are pairwise disjoint and elements of $(K\mathscr{G})_{h}$ are supported on $c^{-1}(h)$.
-One can, in fact, show that $(K\mathscr{G})_{h}$ consists of those elements of $K\mathscr{G}$ supported on $c^{-1}(h)$, but we shall never need this description.
+Then $(R\mathscr{G})_{h}$ is the $R$-span of the characteristic functions of homogeneous compact open bisections of degree $h$.
+Note that each $c^{-1}(h)$ with $h\in H$ is clopen, and so if $U$ is any compact open bisection, then $U=\bigcup_{h\in H}(U\cap c^{-1}(h))$ and the $U\cap c^{-1}(h)$ are pairwise disjoint homogeneous compact open bisections.
+By compactness of $U$, only finitely many of these intersections, say corresponding to $h_{1},\ldots,h_{n}$, are nonempty, and so $1_{U}=\sum_{i=1}^{n}1_{(c^{-1}(h_{i})\cap U)}$, showing that $\sum_{h\in H}(R\mathscr{G})_{h}=R\mathscr{G}$.
+The sum is direct since the $c^{-1}(h)$ are pairwise disjoint and elements of $(R\mathscr{G})_{h}$ are supported on $c^{-1}(h)$.
+One can, in fact, show that $(R\mathscr{G})_{h}$ consists of those elements of $R\mathscr{G}$ supported on $c^{-1}(h)$, but we shall never need this description.
 This argument also shows that the homogeneous compact open bisections form a basis for the topology of $\mathscr{G}$.
 Notice that $\mathscr{G}^{(0)}\subseteq c^{-1}(e)=\mathscr{G}_{e}$ and that $\mathscr{G}_{e}$ is a clopen subgroupoid and hence is ample.
-Moreover, $(K\mathscr{G})_{e}=K\mathscr{G}_{e}$ by definition.
-From now on, we write $K\mathscr{G}_{h}$ instead of the more cumbersome $(K\mathscr{G})_{h}$.
-The following is [21, Definition 2.9].
-Let $R=\bigoplus_{h\in H}R_{h}$ be an $H$-graded $K$-algebra.
-If, for every $h\in H$ and $r\in R_{h}$, there exist $\epsilon_{h}(r)\in R_{h}R_{h^{-1}}$ and $\epsilon_{h}^{\prime}(r)\in R_{h^{-1}}R_{h}$ such that $\epsilon_{g}(r)r=r=r\epsilon_{g}^{\prime}(r)$, then $R$ is called nearly epsilon-strongly $H$-graded.
+Moreover, $(R\mathscr{G})_{e}=R\mathscr{G}_{e}$ by definition.
+From now on, we write $R\mathscr{G}_{h}$ instead of the more cumbersome $(R\mathscr{G})_{h}$.
+The following is [22, Definition 2.9].
+Let $\Lambda=\bigoplus_{h\in H}\Lambda_{h}$ be an $H$-graded $K$-algebra.
+If, for every $h\in H$ and $r\in\Lambda_{h}$, there exist $\epsilon_{h}(r)\in\Lambda_{h}\Lambda_{h^{-1}}$ and $\epsilon_{h}^{\prime}(r)\in\Lambda_{h^{-1}}\Lambda_{h}$ such that $\epsilon_{g}(r)r=r=r\epsilon_{g}^{\prime}(r)$, then $\Lambda$ is called nearly epsilon-strongly $H$-graded.
 It is shown in [22] that being nearly epsilon-strongly $H$-graded is a necessary condition for graded regularity.
 Proposition 4.1.
-Let $\mathscr{G}$ be an ample groupoid with a continuous cocycle $c\colon\mathscr{G}\to H$ and let $K$ be a field. Then $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
+Let $\mathscr{G}$ be an ample groupoid with a continuous cocycle $c\colon\mathscr{G}\to H$ and let $R$ be a ring. Then $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
 Proof.
-Let $\mathscr{G}=\bigoplus_{h\in H}K\mathscr{G}_{h}$ be the $H$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to H$ as above.
+Let $\mathscr{G}=\bigoplus_{h\in H}R\mathscr{G}_{h}$ be the $H$-grading induced by the continuous cocycle $c\colon\mathscr{G}\to H$ as above.
 Fix any $h\in H$ and let $U\subseteq c^{-1}(h)$ be a compact open bisection. Define $\epsilon_{h}(1_{U})=1_{UU^{-1}}$. Then
-$\epsilon_{h}(1_{U})=1_{UU^{-1}}=1_{U}1_{U^{-1}}\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}},$
+$\epsilon_{h}(1_{U})=1_{UU^{-1}}=1_{U}1_{U^{-1}}\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}},$
 and
 $\epsilon_{h}(1_{U})1_{U}=1_{UU^{-1}}1_{U}=1_{UU^{-1}U}=1_{U}.$
-Similarly, defining $\epsilon^{\prime}_{h}(1_{U})=1_{U^{-1}U}$ yields that $\epsilon^{\prime}_{h}(1_{U})\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$ and $1_{U}\epsilon^{\prime}_{h}(1_{U})=1_{U}$.
-Now consider an arbitrary $f\in K\mathscr{G}_{h}$.
-Then there are compact open bisections $U_{1},U_{2},\ldots,U_{n}$, with each $U_{i}\subseteq c^{-1}(h)$, and $k_{1},k_{2},\cdots,k_{n}\in K$ such that $f=\sum_{i=1}^{n}k_{i}1_{U_{i}}$.
-In particular, $\epsilon_{h}(1_{U_{i}})\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$ for each $i$.
+Similarly, defining $\epsilon^{\prime}_{h}(1_{U})=1_{U^{-1}U}$ yields that $\epsilon^{\prime}_{h}(1_{U})\in R\mathscr{G}_{h^{-1}}\,R\mathscr{G}_{h}$ and $1_{U}\epsilon^{\prime}_{h}(1_{U})=1_{U}$.
+Now consider an arbitrary $f\in R\mathscr{G}_{h}$.
+Then there are compact open bisections $U_{1},U_{2},\ldots,U_{n}$, with each $U_{i}\subseteq c^{-1}(h)$, and $r_{1},r_{2},\cdots,r_{n}\in R$ such that $f=\sum_{i=1}^{n}r_{i}1_{U_{i}}$.
+In particular, $\epsilon_{h}(1_{U_{i}})\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}}$ for each $i$.
 Put $U=\bigcup_{i=1}^{n}U_{i}U_{i}^{-1}$.
@@
 Put $\epsilon_{h}(f)=1_{U}$.
-Since $U\subseteq\mathscr{G}^{(0)}$ and $\mathscr{G}^{(0)}$ is Hausdorff, it follows from the inclusion-exclusion principle that $1_{U}=\epsilon_{h}(f)$ is an integer combination of the $1_{U_{i}U_{i}^{-1}}$, and hence $\epsilon_{h}(f)\in K\mathscr{G}_{h}\,K\mathscr{G}_{h^{-1}}$.
-Similarly, putting $U^{\prime}=\bigcup_{i=1}^{n}U_{i}^{-1}U_{i}$ and defining $\epsilon^{\prime}_{h}(f)=1_{U^{\prime}}$, we have that $\epsilon^{\prime}_{h}(f)\in K\mathscr{G}_{h^{-1}}\,K\mathscr{G}_{h}$.
+Since $U\subseteq\mathscr{G}^{(0)}$ and $\mathscr{G}^{(0)}$ is Hausdorff, it follows from the inclusion-exclusion principle that $1_{U}=\epsilon_{h}(f)$ is an integer combination of the $1_{U_{i}U_{i}^{-1}}$, and hence $\epsilon_{h}(f)\in R\mathscr{G}_{h}\,R\mathscr{G}_{h^{-1}}$.
+Similarly, putting $U^{\prime}=\bigcup_{i=1}^{n}U_{i}^{-1}U_{i}$ and defining $\epsilon^{\prime}_{h}(f)=1_{U^{\prime}}$, we have that $\epsilon^{\prime}_{h}(f)\in R\mathscr{G}_{h^{-1}}\,R\mathscr{G}_{h}$.
 It is straightforward from the construction that $\epsilon_{h}(f)f=f=f\epsilon^{\prime}_{h}(f)$.
-Hence, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
+Hence, $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded.
 ∎
 Theorem 4.2.
-Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle.
-Then $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular.
+Let $\mathscr{G}$ be an ample groupoid, $R$ a unital ring, $H$ a discrete group and $c\colon\mathscr{G}\to H$ a continuous cocycle.
+Then $R\mathscr{G}$ is graded regular if and only if $R\mathscr{G}_{e}$ is regular.
 Proof.
-By Proposition 4.1, $K\mathscr{G}$ is nearly epsilon-strongly $H$-graded. Thus, by [21, Theorem 1.2], $K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}$ is regular. ∎
-Theorem 4.2 allows us to upgrade Theorems 3.7 and 3.12 to a theorem about graded regularity.
+By Proposition 4.1, $R\mathscr{G}$ is nearly epsilon-strongly $H$-graded. Thus, by [22, Theorem 1.2], $R\mathscr{G}$ is graded regular if and only if $R\mathscr{G}_{e}$ is regular. ∎
+Theorem 4.2 allows us to upgrade Theorem A to a theorem about graded regularity.
 Theorem 4.3.
-Let $\mathscr{G}$ be an ample groupoid, let $K$ be a field, let $H$ be a discrete group and let $c\colon\mathscr{G}\to H$ be a continuous cocycle.
+Let $\mathscr{G}$ be an ample groupoid, $R$ a unital ring, $H$ a discrete group and $c\colon\mathscr{G}\to H$ a continuous cocycle.
+Then $R\mathscr{G}$ is graded von Neumann regular (with respect to the $H$-grading induced by $c$) if and only if the following conditions hold:
   1. (1)
-    If $\mathscr{G}_{e}$ is approximately quasi-compact and the characteristic of $K$ divides the order of no isotropy element of $\mathscr{G}_{e}$, then $K\mathscr{G}$ is graded regular with respect to the $H$-grading induced by $c$.
+    $R$ is regular
   2. (2)
-    If $K$ is of characteristic 00, then $K\mathscr{G}$ is graded regular (with respect to the $H$-grading induced by $c$) if and only if $\mathscr{G}_{e}$ is approximately quasi-compact.
+    $\mathscr{G}_{e}$ is approximately quasi-compact
+  3. (3)
+    the order of each finite subgroup of an isotropy group of $\mathscr{G}_{e}$ is invertible in $R$.
 5. Applications
-Our first application is to recover the nontrivial part of Okniński’s result on regularity of inverse semigroups in characteristic zero from Theorem 3.7.
-Theorem 5.1 (Okniński).
-Let $S$ be an inverse semigroup and $K$ a field of characteristic zero. Then $KS$ is regular if and only if $S$ is locally finite.
-Proof.
-We already proved in Proposition 2.5 that if $S$ is locally finite, then $KS$ is regular.
-For the converse, assume that $KS$ is regular.
+Our first application is to prove Theorem B.
+Proof of Theorem B.
+The sufficiency of the conditions is proved in Proposition 2.4.
+For sufficiency, suppose that $RS$ is regular.
 Let $\mathscr{G}_{S}$ denote the universal groupoid of $S$.
-Since $KS\cong K\mathscr{G}_{S}$, we have that $K\mathscr{G}_{S}$ is regular.
-Then, by Theorem 3.7, we have that $\mathscr{G}_{S}$ is approximately quasi-compact, and hence $S$ is locally finite by Corollary 3.5 since the canonical mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G}_{S})$ given by $\psi(s)=(s,D_{s^{*}s})$ is injective.
+Then $R\mathscr{G}_{S}\cong RS$ is regular, and so $R$ is regular, $\mathscr{G}_{S}$ is approximately quasi-compact and the order of each finite subgroup of an isotropy group of $\mathscr{G}_{S}$ is invertible in $R$.
+Then $S$ is locally finite by Corollary 3.5 since the canonical mapping $\psi\colon S\to\Gamma_{c}(\mathscr{G}_{S})$ given by $\psi(s)=(s,D_{s^{*}s})$ is injective.
+A maximal subgroup $G_{e}$ of $S$ is isomorphic to the isotropy group at the principal character $x_{e}$, and hence the order of any finite subgroup of $G_{e}$ is invertible in $R$.
 ∎
-Next we give an example to show that the necessary conditions in Theorem 3.1 are not sufficient to guarantee regularity of the algebra of a groupoid by giving examples of principal groupoids with nonregular algebras.
-Recall that if a discrete group $G$ acts on a locally compact, Hausdorff and totally disconnected space $X$ by homeomorphisms, then the groupoid of germs of this action is just the so-called transformation groupoid $\mathscr{G}$ with underlying space $G\times X$ (with the product topology) and unit space $\{1\}\times X$ (identified with $X$) with $\mathop{\boldsymbol{s}}(g,x)=x$ and $\mathop{\boldsymbol{r}}(g,x)=gx$.
+Next we give an example to show that the necessary conditions for regularity in [4] are not sufficient to guarantee regularity of the algebra of a groupoid by giving examples of principal groupoids with nonregular algebras over any regular unital ring $R$.
+Recall that if a discrete group $G$ acts on a locally compact, Hausdorff and totally disconnected space $X$ by homeomorphisms, then the groupoid of germs of this action is just the so-called transformation groupoid $\mathscr{G}$ with underlying space $G\times X$ (with the product topology) and unit space $\{1\}\times X$ (identified with $X$) with $\mathop{\boldsymbol{s}}(g,x)=x$ and $\mathop{\boldsymbol{r}}\nolimits(g,x)=gx$.
 Products are given by $(h,gx)(g,x)=(hg,x)$ and $(g,x)^{-1}=(g^{-1},gx)$.
@@
 In particular, if the action of $G$ is free, then the transformation groupoid is principal.
-It is well known that $K\mathscr{G}$ is isomorphic to the skew group ring $C_{c}(X,K)\rtimes G$ where $C_{c}(X,K)$ is the ring of compactly supported locally constant functions $X\to K$ and $G$ acts on this ring via $(gf)(x)=f(g^{-1}x)$.
+It is well known that $K\mathscr{G}$ is isomorphic to the skew group ring $C_{c}(X,R)\rtimes G$ where $C_{c}(X,R)$ is the ring of compactly supported locally constant functions $X\to R$ and $G$ acts on this ring via $(gf)(x)=f(g^{-1}x)$.
 The first part of the following result extends a result in [18].
@@
   2. (2)
-    If $K$ is a field of characteristic 00, then $K\mathscr{G}$ is regular if and only if $G$ is locally finite.
+    If $R$ is a regular unital ring, then $R\mathscr{G}$ is regular if and only if $G$ is locally finite and the order of each finite subgroup of an isotropy group of the action is invertible in $R$.
 Proof.
-The first item follows from Corollary 3.5 since the map $\psi\colon G\to\Gamma_{c}(\mathscr{G})$ given by $\psi(g)=\{g\}\times X$ is injective.
-The second item follows from the first and Theorem 3.7.
-∎
-For example, if we consider a minimal (nonperiodic) action of $\mathbb{Z}$ on the Cantor set $X$ (e.g., a nonperiodic minimal shift space), then the associated transformation groupoid $\mathscr{G}$ is a principal groupoid (and hence trivially satisfies the necessary conditions in Theorem 3.1), but $\mathbb{C}\mathscr{G}$ is not regular.
+The first item follows from Corollary 3.5 since the map $\psi\colon G\to\Gamma_{c}(\mathscr{G})$ given by $\psi(g)=\{g\}\times X$ is injective. The second item follows from the first and Theorem A. ∎
+For example, if we consider a minimal (nonperiodic) action of $\mathbb{Z}$ on the Cantor set $X$ (e.g., a nonperiodic minimal shift space), then the associated transformation groupoid $\mathscr{G}$ is a principal groupoid (and hence trivially satisfies the necessary conditions in [4]), but $\mathbb{C}\mathscr{G}$ is not regular.
 5.1. Leavitt path algebras
@@
 $Z(\mu,F,\nu)=\{(x,|\mu|-|\nu|,y)\in\mathscr{G}_{E}\mid x\in Z(\mu\setminus F),y\in Z(\nu\setminus F)\}.$
-Then the collection $\{Z(\mu,F,\nu)\}$, ranging over $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, forms a basis of compact open bisections for a locally compact Hausdorff topology on $\mathscr{G}_{E}$ [26, Lemma 9.2].
+Then the collection $\{Z(\mu,F,\nu)\}$, ranging over $\mu,\nu\in E^{*}$ with $r(\mu)=r(\nu)$ and finite $F\subseteq\{e\in E^{1}\mid s(e)=r(\mu)\}$, forms a basis of compact open bisections for a locally compact Hausdorff topology on $\mathscr{G}_{E}$ [28, Lemma 9.2].
 In particular, $\mathscr{G}_{E}$ is an ample groupoid.
@@
 Hence, $\mathscr{G}_{E}$ is principal if and only if $E$ is acyclic.
-Let $K$ be a field and let $L_{K}(E)$ denote the Leavitt path algebra of $E$ over $K$ (see [2] for the definition of $L_{K}(E)$).
+Let $K$ be a commutative ring with unit and let $L_{K}(E)$ denote the Leavitt path algebra of $E$ over $K$ (see [2] for the definition of $L_{K}(E)$).
 Then $L_{K}(E)\cong K\mathscr{G}_{E}$ with the isomorphism given by $\mu\nu^{*}\mapsto 1_{Z(\mu,\nu)}$ [9, Proposition 4.3].
@@
 Theorem 5.2.
-Let $E$ be a directed graph and $K$ a field. The following are equivalent.
+Let $E$ be a directed graph and $K$ a von Neumann regular commutative ring. The following are equivalent.
   1. (1)
@@
 We already observed that acyclic is equivalent to principal.
-If $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular, then $\mathscr{G}_{E}$ has locally finite isotropy by Theorem 3.1 and hence $E$ must be principal as $\mathscr{G}_{E}$ has no infinite cyclic isotropy group.
+If $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular, then $\mathscr{G}_{E}$ has locally finite isotropy by Theorem A and hence $E$ must be principal, as $\mathscr{G}_{E}$ has no infinite cyclic isotropy group.
 Suppose now that $E$ is acyclic.
@@
 If $\mathscr{G}_{E}$ is approximately quasi-compact (and hence approximately uniformly bounded), then all its isotropy groups must be locally finite.
-Since the isotropy groups of $\mathscr{G}_{E}$ are always subgroups of $\mathbb{Z}$, they must then be trivial and so $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular by Theorem 3.12.
+Since the isotropy groups of $\mathscr{G}_{E}$ are always subgroups of $\mathbb{Z}$, they must then be trivial, and so $L_{K}(E)\cong K\mathscr{G}_{E}$ is regular by Theorem A.
 ∎
-Next we show that we can recover [20, Theorem 10], which states that the Leavitt path algebra of an arbitrary graph is graded von Neumann regular.
-Every Leavitt path algebra $L_{K}(E)$ over a field $K$ is $\mathbb{Z}$-graded, with the grading given by homogeneous components
+Next we show that we can recover [20, Theorem 10], which states that the Leavitt path algebra of an arbitrary graph over a field is graded von Neumann regular.
+In fact, we can get away with having coefficients in a commutative regular ring.
+Every Leavitt path algebra $L_{K}(E)$ over a commutative ring $K$ is $\mathbb{Z}$-graded, with the grading given by homogeneous components
 $L_{K}(E)_{n}=\mathrm{span}_{K}\{\alpha\beta^{*}\mid\alpha,\beta\in E^{*},|\alpha|-|\beta|=n\}.$
@@
 Theorem 5.3 (Hazrat).
-Let $E$ be an arbitrary graph and $K$ a field. Then the Leavitt path algebra $L_{K}(E)$ is a graded von Neumann regular ring.
+Let $E$ be an arbitrary graph and $K$ a unital commutative von Neumann regular unital ring. Then the Leavitt path algebra $L_{K}(E)$ is a graded von Neumann regular ring.
 Proof.
@@
 For a finite subgraph $F$ of $E$ and $n\geq 0$, we let
-$\mathscr{G}_{0}(F,n)=\bigcup_{\alpha,\beta\in F^{*},\mathop{\boldsymbol{r}}(\alpha)=\mathop{\boldsymbol{r}}(\beta),|\alpha|=|\beta|\leq n}Z(\alpha,\emptyset,\beta).$
+$\mathscr{G}_{0}(F,n)=\bigcup_{\alpha,\beta\in F^{*},\mathop{\boldsymbol{r}}\nolimits(\alpha)=\mathop{\boldsymbol{r}}\nolimits(\beta),|\alpha|=|\beta|\leq n}Z(\alpha,\emptyset,\beta).$
 Then $\mathscr{G}_{0}(F,n)$ is compact open, as there are only finitely many choices of $\alpha,\beta$ given $F$ and $n$, and is a subgroupoid by (5.1).
@@
 We refer the reader to [16] for a detailed treatment on partial actions and partial skew group rings.
-Let $X$ be a locally compact Hausdorff totally disconnected space, let $G$ be a discrete group with identity element $e$, and let $K$ be a field.
+Let $X$ be a locally compact Hausdorff totally disconnected space, let $G$ be a discrete group with identity element $e$, and let $K$ be a commutative ring.
 Let $\phi=(\{X_{g}\}_{g\in G},\{\phi_{g}\}_{g\in G})$ be a topological partial action of $G$ on $X$.
@@
 By Theorem 4.2, $C_{c}(X,K)\rtimes_{\varphi}G\cong K\mathscr{G}$ is graded regular if and only if $K\mathscr{G}_{e}\cong C_{c}(X,K)$ is regular.
-But $C_{c}(X,K)$ is trivially regular since if $f\in C_{c}(X,K)$ and then $ff^{\prime}f=f$ where $f^{\prime}(x)=1/f(x)$ whenever $f(x)\neq 0$, and $f^{\prime}(x)=0$ otherwise.
+But $C_{c}(X,K)$ is trivially regular.
+Indeed, if $f\in C_{c}(X,K)$, then we can write $f=\sum_{i=1}^{n}k_{i}1_{U_{i}}$ with the $U_{i}$ disjoint compact open sets (as $X$ is Hausdorff).
+If we choose $k_{i}^{\prime}\in K$ with $k_{i}k^{\prime}_{i}k_{i}=k_{i}$ and put $f^{\prime}=\sum_{i=1}^{n}k_{i}^{\prime}1_{U_{i}}$, then $f^{\prime}\in C_{c}(X,K)$ and $ff^{\prime}f=f$.
 ∎
@@
 Suppose $L_{K}(\mathscr{E})\cong K\mathscr{G}$ is regular.
-Then the isotropy subgroups of $\mathscr{G}$ are locally finite by Theorem 3.1.
+Then the isotropy subgroups of $\mathscr{G}$ are locally finite by Theorem A.
 Since all isotropy subgroups of $\mathscr{G}$ are infinite cyclic, it follows that $\mathscr{G}$ is principal.
@@
 Let $\mathscr{E}=(E,\mathscr{L},\mathscr{B})$ be normal labeled space and $K$ a field.
-Suppose that if $\xi^{\alpha}\in\mathsf{T}$ and $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$.
+Suppose that if $\xi^{\alpha}\in\mathsf{T}$ and $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{\geq 1}$.
 Then $L_{K}(\mathscr{E})$ is regular.
@@
 Then, $\mathscr{G}_{\mathscr{E}}$ is a directed union of $\mathscr{G}(F)$, ranging over all finite subgraphs $F$ of $E$, and we conclude that $\mathscr{G}_{\mathscr{E}}$ is approximately quasi-compact.
-Notice that the assumption $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{*}$ implies that all isotropy in $\mathscr{G}_{\mathscr{E}}$ are trivial [14, Proposition 6.10].
+Notice that the assumption $\alpha\neq\beta\gamma^{\infty}$ for any $\beta,\gamma\in\mathscr{L}^{\geq 1}$ implies that all isotropy in $\mathscr{G}_{\mathscr{E}}$ are trivial [14, Proposition 6.10].
 Thus, the order of no isotropy element is divisible by the characteristic of $K$.
-Hence, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is regular by Theorem 3.12.
+Hence, $L_{K}(\mathscr{E})\cong K\mathscr{G}_{\mathscr{E}}$ is regular by Theorem A.
 ∎
@@
 More generally, if $F\subseteq P$, then $F$ is said to be action directed if given $p,q\in F$ with $U_{p}\cap U_{q}\neq\emptyset$, there is $r\in F\cap pP\cap qP$ with $U_{p}\cap U_{q}=U_{r}$.
-It is shown in [30, Claim 3] that every finite subset of $P$ is contained in an action-directed finite subset of $P$.
+It is shown in [31, Claim 3] that every finite subset of $P$ is contained in an action-directed finite subset of $P$.
 From now on we assume that $X$ is locally compact, Hausdorff and totally disconnected.
@@
 These compact open bisections form a basis for the topology, and so the groupoid is ample.
-The topology on $\mathscr{G}(X,P)$ is finer than the product topology on $X\times G\times X$ but coincides with the product topology on sets of the form $D(U,p,q,V)$ (see [30, Lemma 5.11]).
+The topology on $\mathscr{G}(X,P)$ is finer than the product topology on $X\times G\times X$ but coincides with the product topology on sets of the form $D(U,p,q,V)$ (see [31, Lemma 5.11]).
 The groupoid $\mathscr{G}(X,P)$ is Hausdorff.
-There is a continuous cocycle $c\colon\mathscr{G}(X,P)\to G$ given by $c(x,g,y)=g$ (see [30, Proposition 5.12]).
+There is a continuous cocycle $c\colon\mathscr{G}(X,P)\to G$ given by $c(x,g,y)=g$ (see [31, Proposition 5.12]).
 Notice that the boundary path groupoid $\mathscr{G}_{E}$ for a graph $E$ is a Deaconu-Renault groupoid for the action of $\mathbb{N}$ via the shift map.
@@
 Theorem 5.14.
-Let $K$ be a field and let $\mathscr{G}(X,P)$ be the Deaconu-Renault groupoid of the directed action of a cancellative left reversible monoid $P$ on a locally compact, Hausdorff and totally disconnected space $X$ by partial local homeomorphisms.
+Let $R$ be a regular unital ring and let $\mathscr{G}(X,P)$ be the Deaconu-Renault groupoid of the directed action of a cancellative left reversible monoid $P$ on a locally compact, Hausdorff and totally disconnected space $X$ by partial local homeomorphisms.
 Assume that the domain $U_{p}$ of the action of $p$ is clopen for all $p\in P$.
-Then $K\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle, where $G$ is the group of fractions of $P$.
+Then $R\mathscr{G}(X,P)$ is graded regular with respect to the $G$-grading coming from the canonical cocycle, where $G$ is the group of fractions of $P$.
 Proof.
-Note that $c^{-1}(e)=\{(x,e,y)\in\mathscr{G}(X,P)\}$ and hence is a principal groupoid.
+Note that $c^{-1}(e)=\{(x,e,y)\in\mathscr{G}(X,P)\}$, and hence it is a principal groupoid.
 So it suffices by Theorem 4.3 to show that $c^{-1}(e)$ is approximately quasi-compact.
@@
 $\mathscr{G}(W,F)=\bigcup_{p\in F}D(W\cap U_{p},p,p,W\cap U_{p})$
-and hence is compact open, since $W\cap U_{p}$ is compact open because $U_{p}$ is clopen, and the topology on $D(W\cap U_{p},p,p,W\cap U_{p})$ is the product topology on $(W\cap U_{p})\times\{e\}\times(W\cap U_{p})$ by [30, Lemma 5.11].
+and hence is compact open, since $W\cap U_{p}$ is compact open because $U_{p}$ is clopen, and the topology on $D(W\cap U_{p},p,p,W\cap U_{p})$ is the product topology on $(W\cap U_{p})\times\{e\}\times(W\cap U_{p})$ by [31, Lemma 5.11].
 Indeed, since $X$ is Hausdorff, the set of elements of $U_{p}\times U_{p}$ with $xp=yp$ is closed in $U_{p}\times U_{p}$ and hence a compact subset of $(W\cap U_{p})\times(W\cap U_{p})$ as $U_{p}$ is clopen.
@@
 We conclude that $\mathscr{G}(W,F)$ is a compact open subgroupoid.
-If $F,F^{\prime}$ are action-directed finite subsets of $P$ and $W,W^{\prime}$ are compact open subsets of $X$, then there is an action-directed finite subset $F^{\prime\prime}$ with $F\cup F^{\prime}\subseteq F^{\prime\prime}$ by [30, Claim 3].
+If $F,F^{\prime}$ are action-directed finite subsets of $P$ and $W,W^{\prime}$ are compact open subsets of $X$, then there is an action-directed finite subset $F^{\prime\prime}$ with $F\cup F^{\prime}\subseteq F^{\prime\prime}$ by [31, Claim 3].
 Then $\mathscr{G}(W,F),\mathscr{G}(W^{\prime},F^{\prime})\subseteq\mathscr{G}(W\cup W^{\prime},F^{\prime\prime})$ by construction.
@@
 Theorem 5.15.
-Let $(G,E)$ be a faithful finite state self-similar group action of a group $G$ on a finite directed graph $E$ with no sinks and let $K$ be a field of characteristic 00.
+Let $(G,E)$ be a faithful finite state self-similar group action of a group $G$ on a finite directed graph $E$ with no sinks and let $R$ be a unital ring.
 Then the Exel-Pardo algebra $K\mathscr{G}_{G,E}$ is graded regular with respect to its usual $\mathbb{Z}$-grading if and only if $G$ is locally finite.
@@
 There is a homomorphism $\psi\colon G\to\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ given by $\psi(g)=\{[g(v)gv^{*},\nu]\mid v\in Z(v),v\in E^{0}\}$.
-Moreover, $\psi$ is injective because the action is faithful, and so if $g\neq 1$, then $g(x)\neq x$ for some $x\in\partial E$, whence $\mathop{\boldsymbol{r}}([g(s(x))gs(x)^{*},x])=gx\neq x=\mathop{\boldsymbol{s}}([g(s(x))gs(x)^{*},x])$.
+Moreover, $\psi$ is injective because the action is faithful, and so if $g\neq 1$, then $g(x)\neq x$ for some $x\in\partial E$, whence $\mathop{\boldsymbol{r}}\nolimits([g(s(x))gs(x)^{*},x])=gx\neq x=\mathop{\boldsymbol{s}}([g(s(x))gs(x)^{*},x])$.
 But $\psi(1)=G^{(0)}$.
 So $\psi(g)\neq\psi(1)$.
-By Theorem 3.7, $\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ is locally finite and hence $G\cong\psi(G)$ is locally finite.
+By Theorem A, $\Gamma_{c}((\mathscr{G}_{G,E})_{0})$ is locally finite and hence $G\cong\psi(G)$ is locally finite.
 For the converse, let $(S_{G,E})_{0}$ be the inverse subsemigroup of $S_{G,E}$ consisting of 00 and all elements of the form $\alpha g\beta^{*}$ with $r(\alpha)=r(\beta)$ and $|\alpha|=|\beta|$.
@@
 Of course, when $G$ is trivial we recover Hazrat’s theorem on graded regularity of Leavitt path algebras in characteristic 00 [20] for the special case of finite graphs with no sinks.
-Acknowledgments
-We thank Gilles de Castro for showing Example 5.12 to us.
+Appendix: A groupoid-free proof of Theorem B
+In this appendix we give a groupoid-free proof of Theorem B.
+Let $S$ be an inverse semigroup, and let $R$ be a unital ring.
+Then $S$ acts on the left of $E(S)$ via the rule $s\mathrel{\diamond}e=ses^{*}$.
+The $\mathscr{R}$-class of $e\in E(S)$ is the set $R_{e}=\{s\in S\mid ss^{*}=e\}$.
+Dually the $\mathscr{L}$-class of $e$ is $L_{e}=\{s\in S\mid s^{*}s=e\}$.
+The maximal subgroup of $S$ at an idempotent $e$ is the group of units $G_{e}=R_{e}\cap L_{e}$ of the inverse monoid $eSe$ (with identity $e$).
+Note that $G_{e}$ acts freely on the left of $R_{e}$ by multiplication since if $g\in G_{e}$ and $s\in R_{e}$, then we have $gss^{*}g^{*}=geg^{*}=e$, and $s=gs$ implies $e=ss^{*}=gss^{*}=ge=g$.
+Lemma 5.16.
+Let $S$ be an semigroup and suppose that there is an integer $N>0$ with $|L_{e}|\leq N$ for all $e\in E(S)$. Then $S$ is locally finite.
+Proof.
+The semigroup $S$ acts on the left of $L_{e}$ by partial injective mappings via the left Schützenberger representation:
+$s\cdot t=\begin{cases}st,&\text{if}\ st\in L_{e}\\ \text{undefined},&\text{else}\end{cases}$
+for $s\in S$ and $t\in L_{e}$.
+Moreover, the Schützenberger representations separate points [11].
+That is, if $s_{1},s_{2}\in S$ act the same in each Schützenberger representation, then $s_{1}=s_{1}\cdot s_{1}^{*}s_{1}=s_{2}\cdot s_{1}^{*}s_{1}$, and so $s_{1}\leq s_{2}$.
+Dually, we have $s_{2}\leq s_{1}$, and so they are equal.
+It follows that the points of $S$ can be separated by homomorphisms into the symmetric inverse monoid $I_{N}$ of all partial bijections on an $N$-element set.
+Thus $S$ is locally finite by  2.2.
+∎
+Proof of Theorem B.
+The sufficiency of these conditions was established in [34] and was proved in Proposition 2.4.
+Suppose next that $RS$ is regular.
+The augmentation $\varepsilon\colon RS\to R$ given by on the $R$-basis $S$ by $s\mapsto 1$ is a surjective ring homomorphism, and hence $R$ is regular (cf.
+[19, Lemma 1.3]).
+Next observe that if $M=S^{1}$ is the inverse monoid resulting from adjoining an identity to $S$, then $RS$ is an ideal of $RM$ and we have that $R\cong RM/RS$.
+Therefore $RM$ is regular by [19, Lemma 1.3] as $R$ and $RS$ are regular.
+Thus we may assume without loss of generality that $S$ is a monoid.
+Suppose that $X$ is a finite subset of $S$ closed under inversion $x\mapsto x^{*}$.
+We show that $T=\langle X\rangle$ is finite and the order of each maximal subgroup of $T$ is invertible in $R$.
+Since every element of a maximal subgroup of $S$ belongs to some finitely generated inverse subsemigroup, it will follow that the order of each element of a maximal subgroup of $S$ is invertible in $R$.
+There is a left module homomorphism $\pi\colon RS\to RE(S)$ given by $\pi(s)=ss^{*}$ for $s\in S$.
+Indeed, $\pi(ts)=tss^{*}t^{*}=t\mathrel{\diamond}(ss^{*})=t\mathrel{\diamond}\pi(s)$ for $s,t\in S$.
+Evidently, $\ker\pi$ is spanned over $R$ by the elements of the form $s-ss^{*}$ with $s\in S$.
+Let $L\subseteq\ker\pi$ be the left ideal generated by the elements $t-tt^{*}$ with $t\in T$.
+We claim that $L$ is generated by the finite set $Y=\{x-xx^{*}\mid x\in X\}$.
+Indeed, we prove by induction on word length that if $t\in T$, then $t-tt^{*}\in RS\cdot Y$.
+There is nothing to prove if $t\in X$.
+Assume that $t=t_{0}x$ with $x\in X$ and $t_{0}$ having shorter word length.
+Since $t_{0}^{*}$ has the same word length as $t_{0}$, because $X$ is closed under inversion, it follows by induction that $x-xx^{*}$ and $t_{0}^{*}-t_{0}^{*}t_{0}$ are in $RS\cdot Y$.
+Therefore,
+$t-tt^{*}=t_{0}(x-xx^{*})-t_{0}xx^{*}(t_{0}^{*}-t_{0}^{*}t_{0})\in RS\cdot Y$
+as required.
+Since $RS$ is regular, it follows from [19, Theorem 1.1] that there is an idempotent $e\in RS$ with $L=RSe$.
+Let $f=1-e$ and note that $Lf=RSe(1-e)=0$.
+Therefore, $\pi(f)=\pi(1)-\pi(e)=1$ (as $e\in L\subseteq\ker\pi$), and also $tf=tt^{*}f$ for all $t\in T$ (as $(t-tt^{*})f=0$).
+Let $f=\sum_{i=1}^{N}r_{i}s_{i}$ with $r_{i}\neq 0$ for all $i$.
+We claim that $|L_{x}\cap T|\leq N$ for all $x\in E(T)$.
+It will then follow that $T$ is finite by Lemma 5.16.
+Define [^3] Note that $\widehat{f}$ is essentially the image of $f$ under the isomorphism of $RS$ with $R\mathscr{G}_{S}$, cf. [33] $\widehat{f}\colon S\to R$ by letting
+$\widehat{f}(s)=\sum_{s\leq s_{i}}r_{i}.$
+Notice that since $s\leq s_{i}$ if and only if $s=ss^{*}s_{i}$, if and only if $s=s_{i}s^{*}s$, we deduce that $\widehat{f}(s)$ is coefficient of $s$ in $ss^{*}f$ and also the coefficient of $s$ in $fs^{*}s$.
+If $x\in E(S)$, then there are at most $N$ elements in the support of $fx$, and so $\widehat{f}$ has at most $N$ elements of $L_{x}$ in its support.
+We claim that if $t\in T\cap L_{ss^{*}}$, then $\widehat{f}(s)=\widehat{f}(ts)$.
+Indeed, $\widehat{f}(s)$ is the coefficient of $s$ in $ss^{*}f$ and $\widehat{f}(ts)$ is the coefficients of $ts$ in $tss^{*}t^{*}f=tss^{*}t^{*}tf=tss^{*}f$ because $t^{*}\in T$ implies $t^{*}f=t^{*}tf$.
+But since $t^{*}t=ss^{*}$, left multiplication by $t$ is bijective mapping $ss^{*}S=t^{*}tS\to tt^{*}S=tS$ (with inverse left multiplication by $t^{*}$).
+Therefore, the coefficient of $ts$ in $tss^{*}f$ is the same as the coefficient of $s$ in $ss^{*}f$, and so $\widehat{f}(ts)=\widehat{f}(s)$.
+We may now complete the proof. Let $x\in E(T)$. From
+$\sum_{i=1}^{N}r_{i}xs_{i}s_{i}^{*}x^{*}=\pi(xf)=x\mathrel{\diamond}\pi(f)=x\pi(f)x^{*}=x1x^{*}=x$ (5.5)
+we deduce that there is $s$ in the support of $xf$ with $ss^{*}=x$.
+Therefore, $\widehat{f}(s)\neq 0$.
+Now if $t\in L_{x}\cap T$, then $\widehat{f}(ts)=\widehat{f}(s)\neq 0$ by the above.
+Since if $t\in L_{x}$, we have that $(ts)^{*}(ts)=s^{*}t^{*}ts=s^{*}ss^{*}s=s^{*}s$, all the elements $ts$ with $t\in L_{x}\cap T$ are in the support of $\widehat{f}$, and so there are at most $N$ such elements.
+On the other hand, if $t_{1},t_{2}\in L_{x}$ and $t_{1}s=t_{2}s$, then $t_{1}=t_{1}x=t_{1}ss^{*}=t_{2}ss^{*}=t_{2}x=t_{2}$.
+It follows that $L_{x}\cap T$ has at most $N$ elements.
+Finally, we show that if $H_{x}=G_{x}\cap T\subseteq L_{x}\cap T$ is the maximal subgroup of $T$ at $x$, then $|H_{x}|$ is invertible in $R$.
+Indeed, $H_{x}$ acts freely on the left of $R_{x}$.
+Let $A$ be a transversal for $H_{x}\backslash R_{x}$.
+Note that (5.5) implies that
+$\sum_{s\in R_{x}}\widehat{f}(s)=\sum_{xs_{i}s_{i}^{*}x^{*}=x}r_{i}=1.$
+On the other hand, if $s\in R_{x}$, we already observed that $\widehat{f}(ts)=\widehat{f}(s)$ for all $t\in L_{x}\cap T\supseteq H_{x}$.
+Since $H_{x}$ acts freely on the left of $R_{x}$, it follows that
+$1=\sum_{s\in R_{x}}\widehat{f}(s)=\sum_{a\in A}\sum_{h\in H_{x}}f(ha)=|H_{x}|\sum_{a\in A}f(a),$
+and so $|H_{x}|$ is invertible in $R$. This completes the proof. ∎